### v1

# EventFlow: Forecasting Continuous-Time Event Data with Flow Matching

Abstract
Continuous-time event sequences, in which events occur at irregular intervals, are ubiquitous across a wide range of industrial and scientific domains. The contemporary modeling paradigm is to treat such data as realizations of a temporal point process, and in machine learning it is common to model temporal point processes in an autoregressive fashion using a neural network. While autoregressive models are successful in predicting the time of a single subsequent event, their performance can be unsatisfactory in forecasting longer horizons due to cascading errors. We propose EventFlow, a non-autoregressive generative model for temporal point processes. Our model builds on the flow matching framework in order to directly learn joint distributions over event times, side-stepping the autoregressive process. EventFlow is likelihood-free, easy to implement and sample from, and either matches or surpasses the performance of state-of-the-art models in both unconditional and conditional generation tasks on a set of standard benchmarks.
1 Introduction
Many stochastic processes, ranging from consumer behavior (Hernandez et al., 2017) to the occurrence of earthquakes (Ogata, 1998), are best understood as a sequence of discrete events which occur at random times. Any observed event sequence, consisting of one or more event times, may be viewed as a draw from a temporal point process (TPP) (Daley & Vere-Jones, 2003) which characterizes the distribution over such sequences. Given a collection of observed event sequences, faithfully modeling the underlying TPP is critical in both understanding and forecasting the phenomenon of interest.
While multiple different parametric TPP models have been proposed (Hawkes, 1971; Isham & Westcott, 1979), their limited flexibility limits their application when modeling complex real-world sequences. This has motivated the use of neural networks (Du et al., 2016; Mei & Eisner, 2017) in modeling TPPs. To date, most neural network based TPP models are autoregressive in nature (Shchur et al., 2020a; Zhang et al., 2020), where a model is trained to predict the next event time given an observed history of events. However, in many tasks, we are interested not only in the next event, but in the entire sequence of events which is to follow. While these models can achieve high likelihoods, their performance in many-step forecasting tasks can be unsatisfactory due to compounding errors arising from the autoregressive sampling procedure (Xue et al., 2022; Lüdke et al., 2023).
Moreover, existing models are typically trained via a maximum likelihood procedure (see Section 3) which involves computing the CDF implied by the learned model. When using a neural model, computing this CDF necessitates techniques such as Monte Carlo estimation to properly compute the loss (Mei & Eisner, 2017). In addition, sampling from intensity-based models (Du et al., 2016; Mei & Eisner, 2017; Yang et al., 2022) is nontrivial, requiring an expensive and difficult to implement approach based on the thinning algorithm (Lewis & Shedler, 1979; Ogata, 1981; Xue et al., 2024).
Motivated by these limitations, we propose EventFlow, a generative model which directly learns the joint event time distributions, thus allowing us to avoid autoregressive sampling altogether. Our proposed model extends the flow matching framework (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) to the setting of TPPs, where we learn a continuous flow from a reference TPP to our data TPP. At an intuitive level, samples from our model are generated by drawing a collection of event times from a reference distribution and flowing these events along a learned vector field. The number of events is fixed throughout this process, decoupling the event counts and their times, so that the distribution over event counts can be learned or otherwise specified. See Figure 1 for an illustration. As our primary contribution regards the modeling of the event times themselves, we focus on unmarked point processes in this work. More specifically,
  • We propose EventFlow, a novel generative model for temporal point processes. Our model is suitable for both unconditional generation tasks (i.e., generating draws from the underlying data TPP) and conditional generation tasks (e.g., forecasting future events given a history), and is able to forecast multiple events simultaneously.
  • Our model provides a new perspective on modeling TPPs and sidesteps common pitfalls in existing approaches. In particular, EventFlow is likelihood-free, non-autoregressive, easy to sample from, and straightforward to implement.
  • On standard benchmark datasets, EventFlow obtains strong performance on a multi-step forecasting task, outperforming the second-best model by a factor of 2.2 to 4, and matches or exceeds the performance of state-of-the-art models for unconditional generation.
Figure 1: All illustration of forecasting with our EventFlow method. The horizontal axis indicates the flow time $s$, and the vertical axis indicates the support of the TPP $\mathcal{T}=[0,T]$. We first encode the observed history $\mathcal{H}$ into an embedding $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$. At $s=0$, we independently draw $n$ events in the forecasting window $[T_{0},T_{0}+\Delta T]$ from a fixed reference distribution, constituting a sample $\gamma_{0}$ from a mixed-binomial TPP. Each event can be thought of as a particle, which is assigned a velocity by a neural network $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$. Each particle flows along its corresponding velocity field until reaching its terminal point at $s=1$, whereby we obtain a forecasted sequence $\gamma_{1}$.
2 Related Work
Temporal Point Processes
The statistical modeling of temporal point processes (TPPs) is a classical subject with a long history (Daley & Vere-Jones, 2003; Hawkes, 1971; Isham & Westcott, 1979). The contemporary modeling paradigm, based on neural networks (Du et al., 2016), typically operates by learning a history encoder and an event decoder. The history encoder seeks to learn a fixed-dimensional vector representation of the history of a sequence up to some given time, and the decoder seeks to model a distribution over the subsequent event(s).
Numerous models have been proposed for both components. Popular choices for the history encoder include RNN-based models (Du et al., 2016; Shchur et al., 2020a; Mei et al., 2019) or attention-based models (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022). While attention-based encoders can provide longer-range contexts, this benefit typically comes at the cost of additional memory overhead. Similarly, a wide range of forms for the event decoder have also been proposed. The most common approach is to parametrize a conditional intensity function via a neural network.
For instance, several authors (Mei & Eisner, 2017; Zuo et al., 2020; Zhang et al., 2020) model the conditional intensity using a parametric form inspired by the Hawkes process (Hawkes, 1971), and Du et al. (2016) model the (log-)conditional intensity through an affine function of the history embedding. Similarly, Okawa et al. (2019) model the conditional intensity using a mixture of Gaussian kernels.
Most closely related to our work are approaches which use generative models as decoders. These models often do not assume a parametric form for the decoder, enhancing their flexibility. For instance, Xiao et al. (2017b) propose the use of W-GANs to generate new events. Similarly, Shchur et al. (2020a) learn the distribution over the next inter-arrival time via a normalizing flow. Lin et al. (2022) benchmark several choices of generative models, including diffusion, GANs, and VAEs. Despite the flexibility of these models, these approaches are all autoregressive in nature, making them ill-suited for multi-step forecasting tasks. In contrast, Lüdke et al. (2023) propose a diffusion-style model which is able to avoid autoregressive sampling via an iterative refinement procedure.
Our work can be viewed as a novel approach for building flexible decoders for TPPs, extending flow matching to the setting of continuous-time event sequences. In contrast to prior work using generative models, our model is likelihood-free and non-autoregressive, achieving strong performance on long-term forecasting tasks. The work of Lüdke et al. (2023) is perhaps most closely related to ours, but we emphasize that the method of Lüdke et al. (2023) requires an involved training and sampling procedure. In contrast, our method is straightforward to both implement and sample from, while simultaneously outperforming existing approaches.
Flow Matching
The recently introduced flow matching framework (or stochastic interpolants) (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) describes a class of generative models which are closely related to both normalizing flows (Papamakarios et al., 2021) and diffusion models (Ho et al., 2020; Song et al., 2021). Intuitively, these models learn a path of probability distributions which interpolates between a fixed reference distribution and the data distribution. These models are a popular alternative to diffusion, providing greater flexibility in model design, with recent applications in image generation (Ma et al., 2024; Dao et al., 2023), DNA and protein design (Stark et al., 2024; Campbell et al., 2024), and point cloud generation (Buhmann et al., 2023; Wu et al., 2023). To the best of our knowledge, our work is the first to explore flow matching for TPPs.
3 Autoregressive TPP Models
We first provide a brief review of autoregressive point process models and discuss their shortcomings. Informally, one may think of an event sequence as a set $\{t_{k}\}_{k=1}^{n}$ of increasing event times. We will use $\mathcal{H}_{t}$ to represent the history of a sample up to (and including) time $t$, i.e., $\mathcal{H}_{t}=\{t_{k}:t_{k}\leq t\}$. Similarly, we use $\mathcal{H}_{t^{-}}=\{t_{k}:t_{k}<t\}$ to represent the history of a sample prior to time $t$. In the autoregressive setting, the time of a single future event $t$ is modeled conditioned on the observed history of a sequence. This is typically achieved by either directly modeling a distribution over $t$ (Shchur et al., 2020a), or equivalently by modeling a conditional intensity function (Du et al., 2016).
In the first approach, a conditional probability density of the form $p(t\mid\mathcal{H}_{t_{n}})$ is learned, allowing us to specify a joint distribution over event times $p(t_{1},\dots,t_{n})$ autoregressively via ${p(t_{1},\dots,t_{n})=p(t_{1})\prod_{k=2}^{n}p(t_{k}\mid\mathcal{H}_{t_{k-1}})}$.
Alternatively, we may define the conditional intensity $\lambda^{*}(t):=\lambda(t\mid\mathcal{H}_{t^{-}})=p(t\mid\mathcal{H}_{t_{n}})/(1-F(t\mid\mathcal{H}_{t_{n}}))$, where $F(t\mid\mathcal{H}_{t_{n}})=\int_{t_{n}}^{t}p(s\mid\mathcal{H}_{t_{n}})\,\mathrm{d}s$ is the CDF associated with $p(t\mid\mathcal{H}_{t_{n}})$. Informally, the conditional intensity $\lambda^{*}(t)$ can be thought of (Rasmussen, 2011) as the instantaneous rate of occurrence of events at time $t$ given the previous $n$ events and given that no events have occurred since $t_{n}$. By integrating $\lambda^{*}(t)$, one can show that
$F(t\mid\mathcal{H}_{t_{n}})=1-\exp\left(-\int_{t_{n}}^{t}\lambda^{*}(s)\,\mathrm{d}s\right)\qquad p(t\mid\mathcal{H}_{t_{n}})=\lambda^{*}(t)\exp\left(-\int_{t_{n}}^{t}\lambda^{*}(s)\,\mathrm{d}s\right)$ (1)
and thus one may recover the conditional distribution from the conditional intensity under mild additional assumptions (Rasmussen, 2011, Prop 2.2).
The Likelihood Function
Suppose we observe an event sequence $\{t_{k}\}_{k=1}^{n}$ on the interval $[0,T]$. The likelihood of this sequence can be seen loosely as the probability of seeing precisely $n$ events at these times. The likelihood may be expressed in terms of either the density or intensity via
$L\left(\{t_{k}\}\right)=p(t_{1},\dots,t_{n})\left(1-F(T\mid\mathcal{H}_{t_{n}})\right)=\left(\prod_{k=1}^{n}\lambda^{*}(t_{k})\right)\exp\left(-\Lambda^{*}(T)\right)$ (2)
where the CDF term is included to indicate that no events beyond $t_{n}$ have occurred and $\Lambda^{*}(T)=\int_{0}^{T}\lambda^{*}(s)\,\mathrm{d}s$ is the total intensity. Autoregressive models are typically trained by maximizing this likelihood (Du et al., 2016; Mei & Eisner, 2017; Shchur et al., 2020a). We emphasize that this likelihood is not simply the joint event-time density $p_{n}(t_{1},\dots,t_{n})$, as the likelihood measures the fact that no events occur after $t_{n}$.
It is worth noting that evaluating $L(\{t_{k}\})$ can be non-trivial in practice. For models which parametrize $\lambda^{*}(t)$ via a neural network (Du et al., 2016; Mei & Eisner, 2017), computing the total intensity $\Lambda^{*}(T)$ is often done via a Monte Carlo integral, requiring several forward passes of the model to evaluate $\lambda^{*}(t)$ at different values of $t$. Models which directly parametrize the density $p(t\mid\mathcal{H}_{t})$ suffer from the same drawback when computing the corresponding CDF in Equation (2). Moreover, some approaches, such as the diffusion-based approach of Lin et al. (2022), are only trained to maximize an ELBO of $p(t\mid t_{1},\dots,t_{n})$, and are thus unable to compute the proper likelihood in Equation (2).
Sampling Autoregressive Models
In many tasks, we are interested not only in an accurate model of the intensity (or distribution), but also sampling new event sequences from the corresponding distribution. For instance, when forecasting an event sequence, we may want to generate several forecasts in order to provide uncertainty quantification over these predictions. Stochastic forecasts also have the potential to enable probabilistic queries (Boyd et al., 2023; 2022), e.g., estimating hitting times (the time it will take until a specified number of events occurs) or estimating the number of events in some subwindow. However, sampling from existing autoregressive models can be difficult.
For instance, the flow-based model of Shchur et al. (2020a) requires a numerical approximation to the inverse of the model to perform sampling. Similarly, the diffusion-based approach of Lin et al. (2022) can require several hundred forward passes of the model to generate a single event time, rendering it costly when generating long sequences. Moreover, the predictive performance of autoregressive models is often unsatisfactory on multi-step generation tasks due to the accumulation of errors over many steps (Lin et al., 2021; Lüdke et al., 2023). This difficulty is particularly pronounced for intensity-based models (Du et al., 2016; Mei & Eisner, 2017; Zhang et al., 2020), where naively computing the implied distribution in Equation (1) is prohibitively expensive. Instead, sampling from intensity-based models is typically achieved via the thinning algorithm (Ogata, 1981; Lewis & Shedler, 1979). However, this algorithm has several hyperparameters to tune, is challenging to parallelize, and can be difficult for practitioners to implement (Xue et al., 2024).
4 EventFlow
Motivated by the limitations of autoregressive models, we propose EventFlow, which has a number of distinct advantages over existing approaches. First, EventFlow directly models the joint distribution over event times, thereby avoiding autoregression entirely. Second, our model is likelihood-free, avoiding the Monte Carlo estimates needed to estimate the likelihood in Equation (2) during training. Third, sampling from our model amounts to solving an ordinary differential equation. This is straightforward to implement and parallelize, allowing us to avoid the difficulties of thinning-based approaches used in existing models. We build upon the flow matching (or stochastic interpolant) framework (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) to develop our model. We begin below by focusing on the unconditional setting, and later discuss how to extend our method for conditional generation as necessary.
4.1 Preliminaries
We first introduce some necessary background and notation. Let $\mathcal{T}=[0,T]$ be a finite-length interval. The set $\Gamma$ denotes the configuration space of $\mathcal{T}$ (Albeverio et al., 1998), i.e., the set of all finite counting measures on the set $[0,T]$. A point $\gamma\in\Gamma$ corresponds to a measure of the form $\gamma=\sum_{k=1}^{n}\delta[t^{k}]$, i.e., a finite collection of Dirac deltas located at event times $t^{k}\in\mathcal{T}$. A temporal point process (TPP) on $\mathcal{T}$ is a probability distribution $\mu$ over the configuration space $\Gamma$. Informally, $\mu$ represents a distribution over sequences $\gamma$ living in the configuration space $\Gamma$ which constitutes the set of valid sequences. We use $N:\Gamma\to\mathbb{Z}_{\geq 0}$ to denote the counting functional, i.e., $N(\gamma)$ is the number of events in the TPP realization $\gamma$ [^1] This can be thoughts of in terms of the counting process, i.e. $N(\gamma)$ corresponds to the value of the associated counting process at the ending time $T$, or the total number of events in $\gamma$ that have occurred in the interval $[0,T]$. While it is common to represent TPPs as distributions over random sets of event times, in our approach it will be more convenient to represent TPPs as random measures (Kallenberg et al., 2017).
We assume all TPP distributions are atomless (Kallenberg et al., 2017, Ch. 1), i.e., the probability of observing an event at any singleton is zero. In addition, we assume all TPPs are simple (Kallenberg et al., 2017, Ch. 2), i.e., no more than one event can occur simultaneously. Under these assumptions, a TPP $\mu$ can be fully characterized (Daley & Vere-Jones, 2003, Prop. 5.3.II) by a probability distribution which specifies the number of events and a collection of joint densities corresponding to the event times themselves. In a slight abuse of notation, we will write $\mu(n)$ for the corresponding distribution over event counts, and $\{\mu^{n}(t^{1},\dots,t^{n})\}_{n=1}^{\infty}$ for the collection of joint distributions. In other words, for any given $n\in\mathbb{Z}_{\geq 0}$, the probability of observing $n$ events in the interval $\mathcal{T}$ is $\mu(n)$, and $\mu^{n}(t^{1},\dots,t^{n})$ describes the corresponding joint distribution of event times. We further restrict each $\mu^{n}$ to be supported only on the ordered sets, so that we may assume $t^{1}<t^{2}<\dots<t^{n}$.
Let $\mu_{1}$ represent the data distribution and $\mu_{0}$ represent a reference distribution. That is, both $\mu_{0},\mu_{1}\in\mathbb{P}(\Gamma)$ are TPP distributions. To construct our model, we will define a path of TPPs $\eta_{s}\in\mathbb{P}(\Gamma)$ which approximately interpolates from our reference distribution to our data distribution. Throughout, we use $s\in[0,1]$ to denote a flow time and $t\in[0,T]$ to denote an event time. These two time axes are in a sense orthogonal to one another (see Figure 1).
4.2 Balanced Couplings
Our first step is to define a useful notion of couplings (Villani et al., 2009), allowing us to pair event sequences drawn from $\mu_{0}$ with those drawn from $\mu_{1}$. A coupling between two TPPs $\mu,\nu\in\mathbb{P}(\Gamma)$ is a joint probability measure $\rho\in\mathbb{P}(\Gamma\times\Gamma)$ over pairs of event sequences $(\gamma_{0},\gamma_{1})$ such that the marginal distributions of $\rho$ are $\mu$ and $\nu$. We say that the coupling $\rho$ is balanced if draws $(\gamma_{0},\gamma_{1})\sim\rho$ are such that $N(\gamma_{0})=N(\gamma_{1})$ almost surely. In other words, balanced couplings only pair event sequences with equal numbers of events. We will use $\Pi_{b}(\mu,\nu)$ to denote the set of balanced couplings of $\mu,\nu$. We can show that the set $\Pi_{b}(\mu,\nu)$ is nonempty if and only if the event count distributions of $\mu$ and $\nu$ are equal, placing a structural constraint on the suitable choices of a reference measure.
Proposition 1 (Existence of Balanced Couplings).
Let $\mu,\nu\in\mathbb{P}(\Gamma)$ be two TPPs. The set of balanced couplings $\Pi_{b}(\mu,\nu)$ is nonempty if and only if $\mu(n)=\nu(n)$ have the same distribution over event counts.
We provide a proof in Appendix B. To provide some intuition for the proof, in one direction, we may explicitly construct a balanced coupling (which we describe in the next paragraph). Conversely, if a balanced coupling exists, the event count distribution of $\mu$ (and hence $\nu$) may be obtained by sampling from the coupling and counting the number of events occurring in either coordinate.
In practice, we follow a simple strategy for choosing both the reference TPP $\mu_{0}$ and the coupling $\rho$. Suppose $q$ is any given density on our state space $\mathcal{T}$, e.g., a uniform distribution. We take $\mu_{0}$ to be a mixed binomial process (Kallenberg et al., 2017, Ch. 3) whose event count distribution is given by that of the data $\mu_{1}(n)$ and joint event distributions given by independent products of $q$ (up to sorting). That is, to sample from $\mu_{0}$, one can simply sample $n\sim\mu_{1}(n)$ from the empirical event count distribution followed by sampling and sorting $n$ i.i.d. points $t^{k}\sim q$.
To draw a sample from our coupling $\rho$, we first sample a data sequence $\gamma_{1}\sim\mu_{1}$, followed by sampling $N(\gamma_{1})$ events independently from $q$ and sorting to produce a draw $\gamma_{0}\sim\mu_{0}$. We call this coupling the independent balanced coupling of $\mu$ and $\nu$.
4.3 Interpolant Construction
We now proceed to construct our interpolant $\eta_{s}\in\mathbb{P}(\Gamma)$. We will construct this path of TPPs via a local procedure which we then marginalize over a given balanced coupling. To that end, let $\rho$ be any balanced coupling of the reference measure $\mu_{0}$ and the data measure $\mu_{1}$, and suppose $z:=(\gamma_{0},\gamma_{1})\sim\rho$ is a draw from this coupling. As $\rho$ is balanced, we have that
$\gamma_{0}=\sum_{k=1}^{n}\delta[t_{0}^{k}]\qquad\gamma_{1}=\sum_{k=1}^{n}\delta[t_{1}^{k}]$ (3)
are both a collection of $n$ events. As TPPs are fully characterized by their joint distributions over event times, we will henceforth describe our procedure for a fixed (but arbitrary) number of events $n$. First, we the define measure $\gamma_{s}^{z}\in\Gamma$ via
$\gamma_{s}^{z}=\sum_{k=1}^{n}\delta[t_{s}^{k}]\qquad t_{s}^{k}=(1-s)t_{0}^{k}+st_{1}^{k}\qquad 0\leq s\leq 1$ (4)
where we use the superscript $z$ to denote the dependence on the pair $z=(\gamma_{0},\gamma_{1})$. In other words, $\gamma_{s}^{z}$ linearly interpolates each corresponding event in $\gamma_{0}$ and $\gamma_{1}$. This defines a path $(\gamma_{s}^{z})_{s=0}^{1}$ in the configuration space $\Gamma$ which evolves the reference sample $\gamma_{0}$ into the data sample $\gamma_{1}$.
In order to perform the marginalization step, we now lift this deterministic path $(\gamma_{s}^{z})_{s=0}^{1}\subset\Gamma$ to a path of TPP distributions $(\eta_{s}^{z})_{s=0}^{1}\subset\mathbb{P}(\Gamma)$. We define the point process distribution $\eta_{s}^{z}\in\mathbb{P}(\Gamma)$ implicitly by adding independent Gaussian noise to each of the events in $\gamma_{s}^{z}$. That is, a draw $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ may be simulated via
$\hat{\gamma}_{s}^{z}=\sum_{k=1}^{n}\delta\left[t_{s}^{k}+\epsilon^{k}\right]\qquad\epsilon^{k}\sim\mathcal{N}(0,\sigma^{2}).$ (5)
In principle this means that the support of $\eta_{s}^{z}$ is larger than $\mathcal{T}$, but in practice we choose $\sigma^{2}$ sufficiently small such that this is not a concern. The addition of noise $\epsilon_{k}$ is instrumental in obtaining a well-specified model, but in practice the noise variance $\sigma^{2}$ is not a critical hyperparameter. We note that this noising step is typical in flow matching models (Lipman et al., 2023; Tong et al., 2024).
Finally, for any $s\in[0,1]$, we define the marginal TPP measure $\eta_{s}$ via $\eta_{s}=\int\eta_{s}^{z}\,\mathrm{d}\rho(z)$. Observe that, by construction, the event count distribution $\eta_{s}(n)$ is given by $\mu_{1}(n)$ for all $s\in[0,1]$. This path of TPP distributions $\eta_{s}$ approximately interpolates from the reference TPP $\mu_{0}$ at $s=0$ to the data TPP $\mu_{1}$ at $s=1$, in the sense that at the endpoints, the joint event time distributions $\eta_{0}^{n}(t^{1},\dots,t^{n})$ and $\eta_{1}^{n}(t^{1},\dots,t^{n})$ are given by a convolution of $\mu_{0}^{n}(t^{1},\dots,t^{n})$ and $\mu_{1}^{n}(t^{1},\dots,t^{n})$ with the Gaussian $\mathcal{N}(0,\sigma^{2}I_{n})$. As $\sigma^{2}\downarrow 0$, it is clear that we recover a genuine interpolant (Tong et al., 2024).
We now shift our attention to the transport of a single event $t_{s}^{k}$ for a fixed $k$. Through the addition of Gaussian noise, we have constructed a path of Gaussian distributions $\mathcal{N}(t_{s}^{k},\sigma^{2})$ whose mean is determined by the location of the $k$th event at the flow time $s$. This transport of a Gaussian can be achieved infinitesimally through the constant vector field ${v_{s}^{k}:[0,T]\to\mathbb{R}}$ given by $v_{s}^{k}(t)=t_{1}^{k}-t_{0}^{k}$ (Tong et al., 2024). Thus, the evolution in (5) is generated by the vector field $v_{s}^{z}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ given by
$v_{s}^{z}(\gamma)=\begin{bmatrix}v_{s}^{1},\dots,v_{s}^{n}\end{bmatrix}^{\sf T}=\begin{bmatrix}t_{1}^{1}-t_{0}^{1},&\dots,&t_{1}^{n}-t_{0}^{n}\end{bmatrix}^{\sf T}\qquad 0\leq s\leq 1.$ (6)
Informally, we view $v_{s}^{z}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ as prescribing a constant (but different) velocity to each of the $n$ events. For a fixed pair $z=(\gamma_{0},\gamma_{1})$ and a given sample $\gamma_{0}^{\prime}\sim\eta_{0}^{z}$, solving the system of ordinary differential equations $\,\mathrm{d}\gamma_{s}^{\prime}=v_{s}^{z}(\gamma_{s}^{\prime})\,\mathrm{d}s$ with initial condition $\gamma_{0}^{\prime}$ will result in a collection of events which is concentrated around the true event times $\gamma_{1}$. Note that here, we view this differential equation as an ODE in $\mathcal{T}^{n}$. If we draw many samples $\gamma_{0}\sim\eta_{0}^{z}$ and solve the corresponding ODE, the distribution over events at any intermediate time $s$ will be given by $\eta_{s}^{z}$.
In other words, the vector field $v_{s}^{z}$ generates the path of distributions $\eta_{s}^{z}$. However, this path is conditioned on $z$, and we would like to find the vector field $v_{s}$ which generates the unconditional path $\eta_{s}$. As is standard in flow matching (Lipman et al., 2023; Tong et al., 2024; Albergo & Vanden-Eijnden, 2023), the unconditional vector field $v_{s}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ may be obtained through a weighted marginalization procedure via
$v_{s}(\gamma)=\int v_{s}^{z}(\gamma)\frac{\,\mathrm{d}\eta_{s}^{z}}{\,\mathrm{d}\eta_{s}}(\gamma)\,\mathrm{d}\rho(z).$ (7)
We have thus far described a procedure for interpolating between a given reference distribution $\mu_{0}^{n}$ and the data distribution $\mu_{1}^{n}$ for a given, fixed number of events $n$. As $n$ was arbitrary, we have successfully constructed a family of interpolants which will enable us to sample from the joint event distribution for any $n$. However, to fully characterize the TPP distribution, we need to also specify the event count distribution. For unconditional generation tasks, this is straightforward – we simply follow the empirical event count distribution see in the training data. We describe our approach for modeling the event count distribution in conditional tasks in the following section.
4.4 Training, Parametrization, and Sampling
To train the model, we would like to perform regression on the vector fields $v_{s}$ in Equation (7). If we knew this vector field $v_{s}$, we could draw samples from the data TPP by drawing a sample event sequence $\gamma_{0}\sim\mu_{0}$ from the reference TPP, and flowing each event along the vector field $v_{s}$.
1 Sample $\gamma_{1}\sim\mu_{1}$, $s\sim\mathcal{U}[0,1]$, and $\epsilon\sim\mathcal{N}(0,1)$
2 $e_{\mathcal{H}}=\varnothing$
3  /* Null history */
4 if forecast then
5       Sample split time $T_{0}\in[\Delta T,T-\Delta T]$
6       Construct history $\mathcal{H}\leftarrow\{t\in\gamma_{1}:t\leq T_{0}\}$
7       Embed history $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H})$
8       Construct future $\gamma_{1}\leftarrow\{t\in\gamma_{1}:T_{0}<t\leq T_{0}+\Delta T\}$
9
10Set $n\leftarrow N(\gamma_{1})$
11 Sample $t_{0}^{1},\dots,t_{0}^{n}\sim q$ and sort to construct $\gamma_{0}$
12 Compute $\gamma_{s}^{z}$ via $t_{s}^{k}=(1-s)t_{0}^{k}+st_{s}^{k}$
Take a gradient step on $\left\lVert\gamma_{1}-\gamma_{0}-v_{\theta}\left(\gamma_{s}^{z}+\epsilon,s,e_{\mathcal{H}}\right)\right\rVert^{2}$
Algorithm 1 Training Step for EventFlow
Training
Foremost, although the marginal vector field in Equation (7) admits an analytical form, it is intractable to compute in practice as the marginal measure $\eta_{s}$ is not available. To overcome this, we may instead perform regression on the conditional vector fields $v_{s}^{z}$. Here, $v_{\theta}(\gamma_{s},s)$ will represent a neural network with parameters $\theta$ which takes in a sequence $\gamma_{s}$ of $N(\gamma_{s})=n$ event times, along with the flow time $s$. That is, we seek to minimize the loss
$J(\theta)=\mathbb{E}_{s,(\gamma_{0},\gamma_{1}),\hat{\gamma}_{s}^{z}}\left[\left\lVert\gamma_{1}-\gamma_{0}-v_{\theta}(\hat{\gamma}_{s}^{z},s)\right\rVert^{2}\right]$ (8)
which can be shown to be equal to MSE regression on the unconditional $v_{s}$ up to an additive constant not depending on $\theta$ (Lipman et al., 2023; Tong et al., 2024). We note here that, although the regression target $v_{s}^{z}$ is linear, the unconditional vector field $v_{s}$ will in general be nonlinear. In practice, this loss is estimated by uniformly sampling a flow time $s\in[0,1]$, a pair $z=(\gamma_{0},\gamma_{1})\sim\rho$ from our balanced independent coupling, and drawing a noisy interpolant $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ according to Equation (5).
To train the model on a forecasting task, where the goal is to predict a future sequence of events conditioned on a history $\mathcal{H}$, we embed $\mathcal{H}$ into a fixed-dimensional vector representation $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$ via a learned encoder $f_{\theta}$ before providing this to the model $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$ and minimizing Equation (8). Note that we jointly train the encoder $f_{\theta}$ and vector field $v_{\theta}$. See Algorithm 1.
Parametrization
The second challenge is that we must learn a vector field $v_{\theta}(\gamma,s)$ in $n$ dimensions for arbitrary values of $n$. In other words, $v_{\theta}$ is a neural network which takes in a flow time $s\in[0,1]$ and a sequence of events $\gamma$, and must produce $N(\gamma)$ scalar outputs. We achieve this through an attention-based architecture, which we detail in Appendix C. At a high level, the flow time $s$ is transformed via a learnable embedding into a fixed-dimensional vector. Similarly, each event in $\gamma$ is transformed into fixed-dimensional vector via a learned embedding (which is shared across the events, but not the flow time). The flow-time embedding is then added to each event embedding, and the resulting sequence is passed through a standard transformer architecture (Yang et al., 2022; Vaswani, 2017), resulting in a sequence of $N(\gamma)$ vectors. Finally, each of these vectors is projected into one dimension via a linear layer to produce the sequence of $N(\gamma)$ velocities.
For conditional tasks, we must also compute an encoding $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$ of the history $\mathcal{H}$. This is done by a separate transformer encoder, which operates in the same fashion as described in the previous paragraph, but without the use of the flow-time $s$ as an input and without the final linear projection layer. This embedding is fed into the intermediate layers of our velocity network via cross-attention.
Lastly, for forecasting tasks we must also learn a model of the event count distribution $p_{\phi}(n\mid\mathcal{H})$. We treat this as a classification problem, where the goal is to predict the number of events $n$ occurring in the forecast window given the history $\mathcal{H}$. We again use an attention-based model, analogous to our velocity model, but where we aggregate the final sequence embedding by averaging and passing this through a small MLP. The model $p_{\phi}(n\mid\mathcal{H})$ is trained by minimizing the cross-entropy loss.
1 Choose a flow time discretization $10=s_{0}<s_{1}<\dots<s_{K}=$ 1
2 $e_{\mathcal{H}}=\varnothing$
3  /* Null history */
4 if forecast then
5       Embed history $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H}$)
6       Sample $n\sim p_{\phi}(n\mid\mathcal{H})$
7
8else
9       Sample $n\sim\mu_{1}(n)$
10
11Sample $t_{0}^{1},\dots,t_{0}^{n}\sim q$ and sort to construct $\gamma_{0}$
12 for $k=1,2,\dots,K$ do
13       $h_{k}\leftarrow s_{k}-s_{k-1}$
14       $\gamma_{s_{k}}\leftarrow\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1},e_{\mathcal{H}})$
15
Algorithm 2 Sampling Step for EventFlow
Sampling
Once the network $v_{\theta}$ is learned, we may draw samples from the model by first drawing a reference sequence $\gamma_{0}\sim\mu_{0}$ and solving the corresponding system of ODEs parametrized by $v_{\theta}$. More concretely, we first fix a number of events $n$ to generate. When seeking to unconditionally generate new sequences from the underlying data TPP $\mu_{1}$, we simply sample $n$ from the empirical event count distribution $\mu_{1}(n)$. For conditional tasks, we draw $n\sim q(n\mid\mathcal{H})$ from the learned conditional distribution over event counts.
Next, we draw $n$ initial events, corresponding to $s=0$, by sampling and sorting $t_{0}^{1},\dots,t_{0}^{n}\sim q$. In practice, we use $q=\mathcal{N}(0,I_{n})$ as we normalize our sequences into the range $[-1,1]$ during training and sampling (followed by renormalization to the data scale). Since we have fixed $n$, we may view this initial draw as a vector $\gamma_{0}=[t_{0}^{1},\dots,t_{0}^{n}]\in\mathcal{T}^{n}$. This event sequence $\gamma_{0}$ then serves as the initial condition for the system of ODEs $d\gamma_{s}=v_{\theta}(\gamma_{s},s)\,\mathrm{d}s$ which can be solved numerically. In our experiments, we use the forward Euler scheme, i.e., we specify a discretization $\{0=s_{0}<s_{1}<\dots<s_{K}=1\}$ of the flow time (in practice, uniform) and recursively compute
$\gamma_{s_{k}}=\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1})\qquad k=1,2,\dots,K$ (9)
where $h_{k}=s_{k}-s_{k-1}$ represents a scalar step size. While other choices of numerical solvers are certainly possible, we found that this simple scheme was sufficient as the model sample paths are typically close to linear. See Algorithm 2 for the full procedure.
5 Experiments
We study our proposed EventFlow model under two settings. The first is a conditional generative modeling task, where we seek to forecast both the number and times of future events given a history. The second is an unconditional task, where we aim to learn a representation of the underlying TPP distribution from empirical observations and generate new sequences from this distribution. In a sense, this second task can be viewed as a special case of the first with no observed history. Our overall experimental procedure is inspired by that of Lüdke et al. (2023).
Datasets
We evaluate our model across a diverse set of datasets encompassing a wide range of possible point process behaviors. First, we use a collection of six synthetic datasets produced by Omi et al. (2019). We additionally evaluate our model on seven real-world datasets, which are a standard benchmark for modeling unmarked TPPs (Shchur et al., 2020b; Bosser & Taieb, 2023; Lüdke et al., 2023). See Appendix A for additional information regarding our datasets.
Baseline Models
As a point of comparison, we train and evaluate several baseline models. These models were selected as they constitute a set of diverse and highly performant models. For an intensity-based method, we compare against the Neural Hawkes Process (NHP) (Mei & Eisner, 2017). We additionally compare against two intensity-free methods, namely the flow-based IFTPP model (Shchur et al., 2020a) and the diffusion-based model of Lin et al. (2022). Lastly, our strongest baseline is the recently proposed Add-and-Thin model of Lüdke et al. (2023), which can be loosely viewed as a non-autoregressive diffusion model. These models use an RNN-based history encoder, with the exception of Add-and-Thin which uses a CNN-based encoder. See Appendix D for additional details.
Metrics
Evaluating generative TPP models is challenging, as one must take into account both the variable locations and numbers of events. This is particularly challenging for the unconditional setting, where unlike forecasting, we do not have a ground-truth sequence to compare against. Our starting point is a metric (Xiao et al., 2017a) on the space of sequences $\Gamma$, allowing us to measure the distance between two sequences $\gamma=\sum_{k=1}^{n}\delta[t_{k}^{\gamma}]$ and $\eta=\sum_{k=1}^{m}\delta[t_{k}^{\eta}]$ with possibly different numbers of events. Without loss of generality, we assume $n\leq m$, so the distance is given by
$d(\gamma,\eta)=\sum_{k=1}^{n}|t_{k}^{\gamma}-t_{k}^{\eta}|+\sum_{k=n+1}^{m}(T-t_{k}^{\eta})$ (10)
where we recall that we assume the sequences are supported on $\mathcal{T}=[0,T]$. This distance can be understood either as an $L^{1}$ distance between the corresponding counting processes of $\gamma,\eta$ or as a generalization of the $1$-Wasserstein distance to measures of unequal mass. This metric allows us to compare two sequences of possibly unequal lengths, and will be used in our forecasting experiment.
For our unconditional experiment, we require a metric which will capture the distance between the TPP distributions themselves. To do so, we follow the approach suggested by Shchur et al. (2020b) and Lüdke et al. (2023), and use the distance in Equation (10) to calculate an MMD (Gretton et al., 2012). Namely, the MMD between two point process distributions $\mu,\nu\in\mathbb{P}(\Gamma)$ is given by
$\mathsf{MMD}(\mu,\nu)=\mathbb{E}_{\gamma,\gamma^{\prime}\sim\mu}[k(\gamma,\gamma^{\prime})]-2\mathbb{E}_{\gamma\sim\mu,\eta\sim\nu}[k(\gamma,\eta)]+\mathbb{E}_{\eta,\eta^{\prime}\sim\nu}[k(\eta,\eta^{\prime})]$ (11)
where $k:\Gamma\to\Gamma\to\mathbb{R}_{\geq 0}$ is a specified kernel. In accordance with prior work (Shchur et al., 2020b; Lüdke et al., 2023), we use an exponential kernel $k(\gamma,\eta)=\exp\left(-d(\gamma,\eta)/(2\sigma^{2})\right)$ with $\sigma$ chosen to be the median distance between all sequences.
5.1 Forecasting Event Sequences
Table 1: Sequence distance (10) between the forecasted and ground-truth event sequences on a held-out test set. Lower is better. We report the mean $\pm$ one standard deviation over five random seeds. The best mean distance on each dataset is indicated in bold, and the second best by an underline.
PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
IFTPP $2.6{\scriptstyle\pm 0.38}$ $20.4{\scriptstyle\pm 1.74}$ $43.8{\scriptstyle\pm 5.70}$ $3.4{\scriptstyle\pm 0.17}$ $1.9{\scriptstyle\pm 0.15}$ $1.4{\scriptstyle\pm 0.12}$ $2.6{\scriptstyle\pm 0.21}$
NHP $2.2{\scriptstyle\pm 0.07}$ $29.9{\scriptstyle\pm 2.17}$ $67.6{\scriptstyle\pm 1.2}$ $3.4{\scriptstyle\pm 0.50}$ $2.6{\scriptstyle\pm 0.53}$ $1.1{\scriptstyle\pm 0.04}$ $2.1{\scriptstyle\pm 0.52}$
Diffusion $4.5{\scriptstyle\pm 1.26}$ $20.8{\scriptstyle\pm 2.4}$ $64.5{\scriptstyle\pm 19.8}$ $3.8{\scriptstyle\pm 0.51}$ $1.8{\scriptstyle\pm 0.18}$ $1.1{\scriptstyle\pm 0.02}$ $2.6{\scriptstyle\pm 0.12}$
Add-and-Thin $\underline{2.0}{\scriptstyle\pm 0.03}$ $\underline{13.7}{\scriptstyle\pm 0.46}$ $\underline{21.5}{\scriptstyle\pm 0.51}$ $\underline{2.2}{\scriptstyle\pm 0.06}$ $\underline{1.5}{\scriptstyle\pm 0.06}$ $\underline{0.9}{\scriptstyle\pm 0.02}$ $\underline{1.4}{\scriptstyle\pm 0.04}$
EventFlow (ours) $\mathbf{0.5}{\scriptstyle\pm 0.01}$ $\mathbf{4.8}{\scriptstyle\pm 0.25}$ $\mathbf{6.8}{\scriptstyle\pm 0.44}$ $\mathbf{1.0}{\scriptstyle\pm 0.10}$ $\mathbf{0.4}{\scriptstyle\pm 0.02}$ $\mathbf{0.3}{\scriptstyle\pm 0.02}$ $\mathbf{0.5}{\scriptstyle\pm 0.04}$
We first evaluate our model on a multi-step conditional forecasting task, focusing on the real-world datasets. To do so, we set a forecast horizon $\Delta T$ for each of our real-world datasets, and generate event sequences in the range $[T_{0},T_{0}+\Delta T]$ for some given $T_{0}$, conditioned on the history of events $\mathcal{H}_{T_{0}}$. Up to a shift, this means we are taking $\mathcal{T}=[0,\Delta T]$ as the support for our model TPP. The forecast horizon $\Delta T$ is chosen such that the window typically contains multiple events.
At training time, we uniformly sample $T_{0}\in[\Delta T,T-\Delta T]$ and split a given data sequence $\gamma_{1}$ into a history on $[0,T_{0}]$ and a future $[T_{0},T_{0}+\Delta T]$. We encode the history $\mathcal{H}_{T_{0}}$ before training the model to fit the events occurring in the future. At testing time, we perform the same splitting procedure, sampling 50 values of $T_{0}$ for each test set sequence. We then forecast the sequence in $[T_{0},T_{0}+\Delta T]$ via the model and compute the distance (10) between the ground-truth and generated sequences.
We report the results of this experiment in Table 1. Our proposed EventFlow method obtains the lowest average forecasting error across all datasets, obtaining an error which is 2.2 to 4 times lower than the second-best method, Add-and-Thin. Given that the non-autoregressive models (EventFlow, Add-and-Thin) consistently outperform the autoregressive baselines, this is clear evidence that autoregressive models can struggle on multi-step predictions. This is especially true on the Reddit-C and Reddit-S datasets, which exhibit long sequence lengths.
5.2 Unconditional Generation of Event Sequences
Next, we evaluate our model on an unconditional generation task, where we aim to generate new sequences from the underlying data distribution. This task serves as a benchmark to evaluate the methods in terms of how well they are able to fit the underlying TPP. Moreover, learning a general-purpose TPP prior could enable downstream tasks, such as data augmentation (Graikos et al., 2022).
In Tables 2 and 3, we report MMD values (11) for each of the synthetic and real-world datasets. We tune the hyperparameters for all models and perform model selection based on the validation set MMD. MMDs are calculated by sampling 1,000 sequences from each trained model, and estimating Equation (11) using the generated and test set samples. The first row (“data”) is the MMD calculated between samples in the training and validation sets, giving us a sense of the best-case performance.
Overall, we find that our EventFlow method (mean rank: $\mathbf{1.8}$) exhibits uniformly strong performance, obtaining either the best or second best MMD on 11 of the 13 datasets. This is particularly pronounced on the real-world datasets, where we obtain the lowest MMD on 5 of the 7 datasets. We see that IFTPP (mean rank: $\mathbf{2.1}$) is a strong baseline, obtaining results which are competitive with our method. The Add-and-Thin method (mean rank: $\mathbf{2.4}$) is often similarly strong, but struggles on the SC dataset. While the NHP (mean rank: $\mathbf{3.7}$) can obtain good fits, this appears to be dataset dependent, with weak results on the NSR, SC, and Reddit-S datasets. The diffusion baseline (mean rank: $\mathbf{4.8}$) is our weakest baseline, which is perhaps unsurprising as this model can only be trained to maximize the likelihood of a subsequent event and not the overall sequence likelihood.
Table 2: MMDs (1e-2) between the test set and $1,000$ generated sequences on our synthetic datasets. Lower is better. We report the mean $\pm$ one standard deviation over five random seeds. The lowest MMD distance on each dataset is indicated in bold, and the second lowest is indicated by an underline.
Hawkes1 Hawkes2 NSP NSR SC SR
Data $1.3$ $1.3$ $1.8$ $3.0$ $5.7$ $1.1$
IFTPP $\mathbf{1.5}{\scriptstyle\pm 0.4}$ $\mathbf{1.4}{\scriptstyle\pm 0.5}$ $\mathbf{2.3}{\scriptstyle\pm 0.7}$ $\underline{6.2}{\scriptstyle\pm 2.1}$ $\mathbf{5.8}{\scriptstyle\pm 0.5}$ $\mathbf{1.3}{\scriptstyle\pm 0.3}$
NHP $\underline{1.9}{\scriptstyle\pm 0.3}$ $5.2{\scriptstyle\pm 1.6}$ $3.6{\scriptstyle\pm 1.3}$ $12.6{\scriptstyle\pm 1.8}$ $25.4{\scriptstyle\pm 11.5}$ $5.0{\scriptstyle\pm 0.7}$
Diffusion $4.8{\scriptstyle\pm 2.7}$ $5.5{\scriptstyle\pm 3.3}$ $10.8{\scriptstyle\pm 7.5}$ $15.0{\scriptstyle\pm 3.6}$ $9.1{\scriptstyle\pm 1.8}$ $5.1{\scriptstyle\pm 2.8}$
Add-and-Thin $\underline{1.9}{\scriptstyle\pm 0.5}$ $2.5{\scriptstyle\pm 0.3}$ $\underline{2.6}{\scriptstyle\pm 0.5}$ $7.4{\scriptstyle\pm 1.2}$ $22.5{\scriptstyle\pm 0.5}$ $2.2{\scriptstyle\pm 0.8}$
EventFlow (ours) $\underline{1.9}{\scriptstyle\pm 0.2}$ $\underline{2.2}{\scriptstyle\pm 0.1}$ $3.8{\scriptstyle\pm 1.2}$ $\mathbf{4.2}{\scriptstyle\pm 0.5}$ $\underline{8.3}{\scriptstyle\pm 0.4}$ $\underline{1.7}{\scriptstyle\pm 0.3}$
Table 3: MMDs (1e-2) between the test set and $1,000$ generated sequences on our real-world datasets. Lower is better. We report the mean $\pm$ one standard deviation over five random seeds. The lowest MMD distance on each dataset is indicated in bold, and the second lowest is indicated by an underline.
PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
Data $1.3$ $0.6$ $0.4$ $3.1$ $2.6$ $3.6$ $3.1$
IFTPP $5.7{\scriptstyle\pm 1.8}$ $1.3{\scriptstyle\pm 1.2}$ $\underline{1.9}{\scriptstyle\pm 0.6}$ $5.8{\scriptstyle\pm 0.9}$ $\mathbf{2.9}{\scriptstyle\pm 0.6}$ $8.2{\scriptstyle\pm 4.7}$ $\underline{5.1}{\scriptstyle\pm 0.7}$
NHP $7.2{\scriptstyle\pm 0.2}$ $2.2{\scriptstyle\pm 1.6}$ $22.5{\scriptstyle\pm 0.3}$ $\underline{5.0}{\scriptstyle\pm 0.1}$ $7.3{\scriptstyle\pm 0.7}$ $6.7{\scriptstyle\pm 1.5}$ $6.1{\scriptstyle\pm 2.3}$
Diffusion $14.3{\scriptstyle\pm 6.5}$ $3.9{\scriptstyle\pm 1.2}$ $6.2{\scriptstyle\pm 3.3}$ $11.7{\scriptstyle\pm 1.8}$ $12.5{\scriptstyle\pm 1.9}$ $10.9{\scriptstyle\pm 3.8}$ $10.5{\scriptstyle\pm 5.2}$
Add-and-Thin $\underline{2.8}{\scriptstyle\pm 0.5}$ $\underline{1.2}{\scriptstyle\pm 0.27}$ $2.7{\scriptstyle\pm 0.3}$ $5.2{\scriptstyle\pm 0.6}$ $\underline{4.8}{\scriptstyle\pm 0.4}$ $\mathbf{4.5}{\scriptstyle\pm 0.2}$ $\mathbf{3.0}{\scriptstyle\pm 0.5}$
EventFlow (ours) $\mathbf{1.5}{\scriptstyle\pm 0.2}$ $\mathbf{0.7}{\scriptstyle\pm 0.1}$ $\mathbf{0.7}{\scriptstyle\pm 0.1}$ $\mathbf{3.5}{\scriptstyle\pm 0.1}$ $4.9{\scriptstyle\pm 0.7}$ $\underline{6.6}{\scriptstyle\pm 1.2}$ $\mathbf{3.0}{\scriptstyle\pm 0.5}$
6 Conclusion
In this work, we propose EventFlow, a non-autoregressive and likelihood-free generative model for temporal point processes. We demonstrate that EventFlow is able to achieve state-of-the-art results on a multi-step forecasting task and strong performance on unconditional generation.
There are several directions in which our work could be extended. First, we do not explicitly enforce that the support of our model TPP is $[0,T]$. This would necessitate moving beyond the Gaussian setup, which is non-trivial to carry out. Second, more sophisticated approaches to learning the event count distribution $p_{\phi}(n\mid\mathcal{H})$ could lead to improved performance. Overall, we believe that our work lays a foundation for future flow-based TPPs, and exploring their application in tasks such as imputation, marked TPPs, and querying (Boyd et al., 2023) are exciting directions.
Reproducibility Statement
Our experiments were conducted with fixed random seeds and tested to ensure reproducibility. After publication, we will open-source our code to facilitate reproducibility.
Acknowledgments
This research was supported by the National Institutes of Health under award R01-LM013344 and by the Hasso Plattner Institute (HPI) Research Center in Machine Learning and Data Science at the University of California, Irvine.
Appendix A Datasets
In this section, we provide some additional details regarding the datasets used in this work. In Table 4, we report the number of sequences in each dataset, some basic statistics regarding the number of events in each sequence, and their support $[0,T]$ and chosen forecast window $\Delta T$. In all datasets, we use 60% of the data for training, 20% for validation, and the remaining 20% for testing.
Synthetic Datasets
Our synthetic datasets are adopted from those proposed by Omi et al. (2019). Each of these datasets consists of $1,000$ sequences supported on $\mathcal{T}=[0,100]$. These synthetic datasets are chosen as they exhibit a wide range of behavior, ranging from i.i.d. inter-arrival times to self-correcting processes which discourage rapid bursts of events. We refer to Section 4 of Omi et al. (2019) for details.
Real-World Datasets
We use the set of real-world datasets proposed in Shchur et al. (2020b), which constitute a set of standard benchmark datasets for unmarked TPPs. We refer to Appendix D of Shchur et al. (2020b) for additional details. With the exception of PUBG, these datasets are supported on $\mathcal{T}=[0,24]$, i.e. each sequence corresponds to a single day. For the PUBG dataset, $\mathcal{T}=[0,38]$ corresponds to the maximum length (in minutes) of an online game of PUBG. We note that PUBG has the largest number of sequences (which can lead to slow training), and the Reddit-C and Reddit-S datasets have long sequences (which can lead to slow training and high memory costs).
Table 4: Some basic summary statistics of the datasets we consider in this work.
Sequences Mean length Std length Range length Support $\Delta T$
Hawkes1 1000 95.4 45.8 [14, 300] [0, 100] {-}-
Hawkes2 1000 97.2 49.1 [18, 355] [0, 100] {-}-
Nonstationary Poisson 1000 100.3 9.8 [71, 134] [0, 100] {-}-
Nonstationary Renewal 1000 98 2.9 [86, 100] [0, 100] {-}-
Stationary Renewal 1000 109.2 38.1 [1, 219] [0, 100] {-}-
Self-Correcting 1000 100.3 0.74 [98, 102] [0, 100] {-}-
PUBG 3001 76.5 8.8 [26, 97] [0, 38] $5$
Reddit-C 1356 295.7 317.9 [1, 2137] [0, 24] $4$
Reddit-S 1094 1129 359.5 [363, 2658] [0, 24] $4$
Taxi 182 98.4 20 [12, 140] [0, 24] $4$
Twitter 2019 14.9 14 [1, 169] [0, 24] $4$
Yelp-Airport 319 30.5 7.5 [9, 55] [0, 24] $4$
Yelp-Miss. 319 55.2 15.9 [3, 107] [0, 24] $4$
Appendix B Proofs
See 1
Proof.
Let $A_{1},A_{2}\subseteq\Gamma$ be Borel measurable (Daley & Vere-Jones, 2003, Prop. 5.3) subsets of the configuration space $\Gamma$, i.e. each of $A_{1},A_{2}$ is a measurable collection of event sequences. Observe that for $i=1,2$, each $A_{i}$ can be written as a disjoint union
$A_{i}^{n}=\bigcup_{n=0}^{\infty}\mathcal{T}^{n}\cap A_{i}$ (12)
i.e. $A_{i}^{n}\subseteq A_{i}$ is the subset of $A_{i}$ containing only sequences with $n$ events. Note each $A_{i}^{n}$ is a Borel measurable subset of $\mathcal{T}^{n}$.
Now, suppose that $\mu(n)=\nu(n)$ have equal event count distributions. We define the coupling $\rho\in\mathbb{P}(\Gamma\times\Gamma)$ by
$\rho(A_{1}\times A_{2})=\sum_{n=0}^{\infty}\mu(n)\mu^{n}(A_{1}^{n})\nu^{n}(A_{2}^{n}).$ (13)
Here, in a slight abuse of notation, we use $\mu^{n},\nu^{n}$ to denote the corresponding joint probability measures over $n$ events, i.e., both are Borel probability measures on $\mathcal{T}^{n}$. Since the $n$-dimensional projection of $\Gamma$ in Equation (12) is simply $\mathcal{T}^{n}$, it is immediate that $\rho(A_{1}\times\Gamma)=\mu(A_{1})$ and $\rho(\Gamma\times A_{2})=\nu(A_{2})$, so that $\rho$ is indeed a coupling. Moreover, it is clear that the coupling is balanced.
Conversely, suppose $\rho\in\Pi_{b}(\mu_{0},\mu_{1})$ is a balanced coupling. Let $N:\Gamma\to\mathbb{Z}_{\geq 0}$ be the event counting functional and let $\pi^{1},\pi^{2}:\Gamma\times\Gamma\to\Gamma$ denote the canonical projections of $\Gamma\times\Gamma$ onto its components. That is, $\pi^{1}:(\gamma_{0},\gamma_{1})\mapsto\gamma_{0}$ and $\pi^{2}:(\gamma_{0},\gamma_{1})\mapsto\gamma_{1}$. Furthermore, let $(N,N):\Gamma\times\Gamma\to\mathbb{Z}_{\geq 0}\times\mathbb{Z}_{\geq 0}$ denote the product of the counting functional, i.e. $(N,N)(\gamma_{0},\gamma_{1})=(N(\gamma_{0}),N(\gamma_{1}))$. Note that the pushforward $N_{\#}\mu$ yields the event count distribution $\mu(n)$ of $\mu$ (and analogously for $\nu$).
Now, observe that composing the projections and counting functionals yields
$\pi^{1}\circ(N,N)=N\circ\pi^{1}\qquad\pi^{2}\circ(N,N)=N\circ\pi^{2}.$ (14)
As $\rho$ is a coupling, we have that $\mu=\pi^{1}_{\#}\rho$ and $\nu=\pi^{2}_{\#}\rho$. From these observations, it follows that
$N_{\#}\mu$ $=N_{\#}\left(\pi^{1}_{\#}\rho\right)$ (15)
$=(N\circ\pi^{1})_{\#}\rho$ (16)
$=(\pi^{1}\circ(N,N))_{\#}\rho$ (17)
$=\pi^{1}_{\#}\left((N,N)_{\#}\rho\right)$ (18)
$=\pi^{2}_{\#}\left((N,N)_{\#}\rho\right)$ (19)
$=N_{\#}\nu$ (20)
where the equality in the penultimate line follows from the fact that $\rho$ is balanced. Thus, we have shown that the existence of a balanced coupling implies that $N_{\#}\mu=N_{\#}\nu$, i.e. the event count distributions are equal. ∎
Appendix C EventFlow Architecture and Training Details
Here, we provide additional details regarding the parametrization and training of our EventFlow model. In general, our model is based on the transformer architecture (Vaswani, 2017; Yang et al., 2022), due to its general ability to handle variable length inputs and outputs, high flexibility, and ability to incorporate long-range interactions. In all settings, our reference measure $\mu_{0}$ is specified with $q=\mathcal{N}(0,I)$.
Model Parametrization
For our unconditional model, we first embed the sequence times $\gamma_{s}$, the flow-time $s$, and the sequence position indices using sinusoidal embeddings followed by an additional linear layer. There are three linear layers in total – one for the flow time, one shared across the sequence times, and one for the position indices. These embeddings are added together to create a representation of the sequence, and we apply a standard transformer to this sequence to produce a sequence of vectors of length $N(\gamma_{s})$. Finally, each of these vectors is projected to one dimension via a final linear layer with shared weights to produce the vector field $v_{\theta}(\gamma_{s},s)$. See Figure 2.
For the conditional model, we use a standard transformer encoder-decoder architecture. We first embed the history sequence times $\mathcal{H}$ and the sequence position indices in a manner analogous to the above. In addition, the model was provided the start of the prediction window $T_{0}$ by concatenating it as the final event in $\mathcal{H}$. This yielded better results than encoding the start of the prediction window separately. We feed these embeddings through the transformer encoder produce an intermediate representation $e_{\mathcal{H}}$.
For the decoder, we provide the model with the current state $\gamma_{s}$ (corresponding to the generated event times at flow-time $s$), the flow-time $s$, and the corresponding positional indices. These are embedded as previously described, before being passed into the transformer decoder. The history encoding $e_{\mathcal{H}}$ is provided to the decoder via cross-attention in the intermediate layer. This produces a sequence of $N(\gamma_{s})$ vectors, which we again pass through a final linear layer to produce the final conditional vector field $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$. See Figure 3.
Our architecture for predicting the number of future events given a history, i.e. $p(n\mid\mathcal{H})$, is again based on the transformer decoder, sharing the same overall architecture as our unconditional model. However, the key difference is that we instead take a mean of the final sequence embeddings before passing this through a small MLP to produce the final logit. See Figure 4.
Training and Tuning
We normalize all sequences to the range $[-1,1]$, using the overall min/max event time seen in the training data. All sequences are generated on this normalized scale, prior to re-scaling the sequence back to the original data range before evaluation. Our model is trained with the Adam (Kingma & Ba, 2015) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for $30,000$ steps with a cosine scheduler (Loshchilov & Hutter, 2017), which cycled every $10,000$ steps. Final hyperparameters were selected by best performance on the validation dataset achieved at any point during the training, where models were evaluated 10 times throughout their training.
To tune our model, we performed a grid search over learning rates in $\{5\times 10^{-3},10^{-3},5\times 10^{-4}\}$ and dropout probabilities in $\{0,0.1,0.2\}$. Overall, we found that learning rates of $10^{-2}$ or larger often caused the model to diverge, and a dropout of $0.1$ yielded the best results across all settings. We use $6$ transformer layers, $8$ attention heads, and an embedding dimension of $512$ across all settings, except for the Reddit-C and Reddit-S datasets where we use $4$ heads and an embedding dimension of $128$ due to the increased memory cost of these datasets.
Figure 2: Overview of our model architecture for unconditional generation. The model takes as input the flow time $s$ and current sequence state $\gamma_{s}=\sum_{k=1}^{n}\delta[t_{s}^{k}]$. Each input is projected to a fixed-length vector via a learnable embedding. The resulting embeddings are added and passed to the transformer model, which produces a sequence of output velocities $v_{\theta}(\gamma_{s},s)$ with $N(\gamma_{s})$ components.
Figure 3: Overview of our model architecture for conditional generation. The encoder (left) takes as input the observed history $\mathcal{H}$, which is embedded in a fashion analogous to our unconditional model. The decoder (right) takes as input the flow time $s$ and current state $\gamma_{s}=\sum_{k=1}\delta[t_{s}^{k}]$. These are embedded and passed through the decoder, which applies cross attention to produce the conditional velocities $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$.
Figure 4: Overview of our architecture modeling the event count distribution $p_{\phi}(n\mid\mathcal{H})$. The model takes as input an observed history $\mathcal{H}$. As in our other architectures, the events are embedded and passed through a transformer. Here, the final sequence embedding output by the transformer is averaged and passed through an additional residual MLP with three layers to produce the logit corresponding to $p(n\mid\mathcal{H})$.
Table 5: The best hyperparameter settings found for the vector field $v_{\theta}$ in our EventFlow method on the unconditional generation task.
Learning Rate Emb. Dim. MLP Dim Heads Transformer Layers
Hawkes1 $10^{-3}$ $512$ $2048$ $8$ $6$
Hawkes2 $10^{-3}$ $512$ $2048$ $8$ $6$
Nonstationary Poisson $10^{-3}$ $512$ $2048$ $8$ $6$
Nonstationary Renewal $10^{-3}$ $512$ $2048$ $8$ $6$
Stationary Renewal $10^{-3}$ $512$ $2048$ $8$ $6$
Self-Correcting $10^{-3}$ $512$ $2048$ $8$ $6$
PUBG $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Reddit-C $10^{-3}$ $128$ $256$ $4$ $6$
Reddit-S $5\times 10^{-3}$ $128$ $256$ $4$ $6$
Taxi $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Twitter $10^{-3}$ $512$ $2048$ $8$ $6$
Yelp-Airport $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Yelp-Miss. $10^{-3}$ $512$ $2048$ $8$ $6$
Table 6: The best hyperparameter settings found for the vector field $v_{\theta}$ in our EventFlow method on the forecasting task.
Learning Rate Emb. Dim. MLP Dim. Heads Transformer Layers
PUBG $10^{-3}$ $512$ $2048$ $8$ $6$
Reddit-C $10^{-3}$ $128$ $256$ $4$ $6$
Reddit-S $10^{-3}$ $128$ $256$ $4$ $6$
Taxi $10^{-3}$ $512$ $2048$ $8$ $6$
Twitter $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Yelp-Airport $10^{-3}$ $512$ $2048$ $8$ $6$
Yelp-Miss. $10^{-3}$ $512$ $2048$ $8$ $6$
Table 7: The best hyperparameter settings found for the event count predictor $p(n\mid\mathcal{H})$ in our EventFlow method on the forecasting task.
Learning Rate Emb. Dim. MLP Dim. Heads Transformer Layers
PUBG $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Reddit-C $10^{-3}$ $128$ $256$ $4$ $6$
Reddit-S $10^{-3}$ $128$ $256$ $4$ $6$
Taxi $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Twitter $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Yelp-Airport $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Yelp-Miss. $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Appendix D Additional details on baselines
In this section, we provide additional details regarding our baseline methods. All methods are trained at a batch size of $64$ for $1,000$ epochs, using early stopping on the validation set loss. In early experiments, we also evaluated AttNHP (Zuo et al., 2020), a variant of the NHP which uses an attention-based encoder, but found it to be prohibitively expensive in terms of memory cost (requiring more than 24 GB of VRAM) and, as a result, do not include it in our results.
IFTPP
Our first baseline is the intensity-free TPP model of Shchur et al. (2020a). This model uses an RNN encoder and a mixture of log-normal distributions to parametrize the decoder. We directly use the implementation provided by the authors. [^2] URL: https://github.com/shchur/ifl-tpp. We train for $1,000$ epochs with early stopping based on the validation set loss. To tune this baseline, we performed a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$, weight decays in $\{0,10^{-6},10^{-5},10^{-4}\}$, history embedding dimensions $\{32,64,128\}$, and mixture component counts $\{8,16,32,64\}$. Our best hyperparameters can be found in Table 8 and Table 9.
Table 8: The best hyperparameter settings found for IFTPP on the unconditional generation task.
Learning Rate Weight Decay Embedding Dimension Mixture Components
Hawkes1 $10^{-3}$ $10^{-4}$ $32$ $8$
Hawkes2 $10^{-2}$ 00 $32$ $8$
Nonstationary Poisson $10^{-3}$ $10^{-6}$ $128$ $8$
Nonstationary Renewal $10^{-2}$ $10^{-6}$ $64$ $16$
Stationary Renewal $10^{-3}$ $10^{-4}$ $32$ $8$
Self-Correcting $10^{-3}$ $10^{-6}$ $32$ $64$
PUBG $10^{-2}$ 00 $128$ $32$
Reddit-C $10^{-3}$ $10^{-4}$ $64$ $16$
Reddit-S $10^{-2}$ $10^{-4}$ $64$ $16$
Taxi $10^{-2}$ $10^{-5}$ $128$ $64$
Twitter $10^{-3}$ $10^{-4}$ $64$ $6$
Yelp-Airport $10^{-2}$ $10^{-6}$ $64$ $64$
Yelp-Miss. $10^{-3}$ $10^{-4}$ $32$ $8$
Table 9: The best hyperparameter settings found for IFTPP on the forecasting task.
Learning Rate Weight Decay Embedding Dimension Mixture Components
PUBG $10^{-4}$ $10^{-6}$ $32$ $32$
Reddit-C $10^{-2}$ 00 $64$ $8$
Reddit-S $10^{-2}$ 00 $64$ $16$
Taxi $10^{-3}$ $10^{-6}$ $128$ $8$
Twitter $10^{-2}$ $10^{-5}$ $32$ $8$
Yelp-Airport $10^{-2}$ $10^{-6}$ $128$ $32$
Yelp-Miss. $10^{-2}$ $10^{-6}$ $32$ $8$
NHP
We additionally compare against the Neural Hawkes Process of Mei & Eisner (2017). This model uses an LSTM encoder and a parametric form, whose weights are modeled by a neural network, to model the conditional intensity function. In practice, we use the implementation proved by the EasyTPP benchmark (Xue et al., 2024), as this version implements the necessary thinning algorithm for sampling. [^3] URL: https://github.com/ant-research/EasyTemporalPointProcess We perform a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$ and embedding dimensions in $\{32,64,128\}$. These hyperparameters are chosen as the EasyTPP implementation allows these to be configured easily. Our best hyperparameters are reported in Table 10 and Table 11.
Diffusion
Our diffusion baseline is based on the implementation of Lin et al. (2022), and our decoder model architecture is taken directly from the code of Lin et al. (2022). [^4] URL: https://github.com/EDAPINENUT/GNTPP At a high level, this model is a discrete-time diffusion model (Ho et al., 2020) trained to generate a single inter-arrival time given a history embedding. Note that as the likelihood is not available in diffusion models, the CDF in the likelihood in Equation (2) is not tractable. Instead, the model is trained by maximizing an ELBO of only the subsequent inter-arrival time.
In preliminary experiments, we found that the codebase provided by Lin et al. (2022) often produced NaN values during sampling, prompting us to make several changes. First, we use the RNN encoder from Shchur et al. (2020a), i.e. the same encoder as the IFTPP baseline, to reduce the memory requirements of the model. Second, we do not log-scale the inter-arrival times as suggested by Lin et al. (2022), as we found that this often led to overflow and underflow issues at sampling time. Third, we do not normalize the data via standardization (i.e., subtracting off the mean inter-arrival time and dividing by the standard deviation), but rather, we scale the inter-arrival times so that they are in the bounded range $[-1,1]$. This is aligned with standard diffusion implementations (Ho et al., 2020), and allows us to perform clipping at sampling time to avoid the accumulation of errors. With these changes, our diffusion baseline is competitive, and able to obtain stronger results than previous work has reported (Lüdke et al., 2023).
We use $1000$ diffusion steps and the cosine beta schedule (Nichol & Dhariwal, 2021), and we train the model on the simplified $\epsilon$-prediction loss of Ho et al. (2020). We train for $1,000$ epochs with early stopping based on the validation set loss. To tune this baseline, we performed a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$, weight decays in $\{0,10^{-6},10^{-5},10^{-4}\}$, history embedding dimensions $\{32,64,128\}$, and layer numbers $\{2,4,6\}$. Our best hyperparameters can be found in Table 12 and Table 13.
Add-and-Thin
We compare to the Add-and-Thin model of Lüdke et al. (2023) as a recently proposed non-autoregressive baseline. We directly run the code provided by the authors without additional modifications. [^5] URL: https://github.com/davecasp/add-thin We do, however, perform a slightly larger hyperparameter sweep than Lüdke et al. (2023), in order to ensure a fair comparison between the methods considered. We train for $1,000$ epochs with early stopping on the validation loss. Tuning is performed via a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$ and number of mixture components in $\{8,16,32,64\}$. We choose to tune only these hyperparameters in order to follow the implementation provided by the authors. Our best hyperparameters can be found in Table 14 and Table 15.
Table 10: The best hyperparameter settings found for NHP on the unconditional generation task.
Learning Rate Embedding Dimension
Hawkes1 $10^{-3}$ $64$
Hawkes2 $10^{-3}$ $64$
Nonstationary Poisson $10^{-3}$ $64$
Nonstationary Renewal $10^{-4}$ $64$
Stationary Renewal $10^{-3}$ $64$
Self-Correcting $10^{-3}$ $64$
PUBG $10^{-4}$ $64$
Reddit-C $10^{-2}$ $64$
Reddit-S $10^{-2}$ $64$
Taxi $10^{-2}$ $64$
Twitter $10^{-4}$ $64$
Yelp-Airport $10^{-3}$ $128$
Yelp-Miss. $10^{-2}$ $64$
Table 11: The best hyperparameter settings found for NHP on the forecasting task.
Learning Rate Embedding Dimension
PUBG $10^{-3}$ $128$
Reddit-C $10^{-2}$ $64$
Reddit-S $10^{-2}$ $64$
Taxi $10^{-2}$ $128$
Twitter $10^{-2}$ $128$
Yelp-Airport $10^{-3}$ $64$
Yelp-Miss. $10^{-2}$ $64$
Table 12: The best hyperparameter settings found for diffusion on the unconditional generation task.
Learning Rate Weight Decay Embedding Dimension Layers
Hawkes1 $10^{-3}$ $10^{-6}$ $64$ $2$
Hawkes2 $10^{-2}$ $10^{-5}$ $64$ $4$
Nonstationary Poisson $10^{-3}$ $10^{-5}$ $128$ $2$
Nonstationary Renewal $10^{-3}$ $10^{-4}$ $64$ $2$
Stationary Renewal $10^{-2}$ 00 $32$ $6$
Self-Correcting $10^{-3}$ 00 $32$ $6$
PUBG $10^{-3}$ 00 $64$ $2$
Reddit-C $10^{-3}$ $10^{-6}$ $128$ $4$
Reddit-S $10^{-3}$ 00 $64$ $4$
Taxi $10^{-2}$ 00 $128$ $4$
Twitter $10^{-3}$ $10^{-4}$ $64$ $6$
Yelp-Airport $10^{-2}$ 00 $32$ $2$
Yelp-Miss. $10^{-2}$ $10^{-5}$ $128$ $2$
Table 13: The best hyperparameter settings found for diffusion on the forecasting task.
Learning Rate Weight Decay Embedding Dimension Layers
PUBG $10^{-4}$ $10^{-5}$ $32$ $6$
Reddit-C $10^{-2}$ $10^{-6}$ $64$ $6$
Reddit-S $10^{-3}$ 00 $64$ $4$
Taxi $10^{-3}$ $10^{-6}$ $32$ $2$
Twitter $10^{-4}$ $10^{-5}$ $64$ $6$
Yelp-Airport $10^{-4}$ $10^{-5}$ $64$ $6$
Yelp-Miss. $10^{-3}$ $10^{-5}$ $32$ $4$
Table 14: The best hyperparameter settings found for Add-and-Thin on the unconditional generation task.
Learning Rate Mixture Components
Hawkes1 $10^{-3}$ $32$
Hawkes2 $10^{-2}$ $32$
Nonstationary Poisson $10^{-2}$ $16$
Nonstationary Renewal $10^{-2}$ $8$
Stationary Renewal $10^{-2}$ $8$
Self-Correcting $10^{-4}$ $8$
PUBG $10^{-3}$ $8$
Reddit-C $10^{-2}$ $32$
Reddit-S $10^{-2}$ $16$
Taxi $10^{-2}$ $8$
Twitter $10^{-4}$ $32$
Yelp-Airport $10^{-4}$ $8$
Yelp-Miss. $10^{-2}$ $64$
Table 15: The best hyperparameter settings found for Add-and-Thin on the forecasting task.
Learning Rate Mixture Components
PUBG $10^{-2}$ $64$
Reddit-C $10^{-2}$ $16$
Reddit-S $10^{-2}$ $64$
Taxi $10^{-2}$ $8$
Twitter $10^{-3}$ $8$
Yelp-Airport $10^{-2}$ $32$
Yelp-Miss. $10^{-3}$ $16$

### v2

@@
-EventFlow: Forecasting Continuous-Time Event Data with Flow Matching
+EventFlow: Forecasting Temporal Point Processes with Flow Matching
 Abstract
@@
 The contemporary modeling paradigm is to treat such data as realizations of a temporal point process, and in machine learning it is common to model temporal point processes in an autoregressive fashion using a neural network.
-While autoregressive models are successful in predicting the time of a single subsequent event, their performance can be unsatisfactory in forecasting longer horizons due to cascading errors.
+While autoregressive models are successful in predicting the time of a single subsequent event, their performance can degrade when forecasting longer horizons due to cascading errors and myopic predictions.
 We propose EventFlow, a non-autoregressive generative model for temporal point processes.
-Our model builds on the flow matching framework in order to directly learn joint distributions over event times, side-stepping the autoregressive process.
-EventFlow is likelihood-free, easy to implement and sample from, and either matches or surpasses the performance of state-of-the-art models in both unconditional and conditional generation tasks on a set of standard benchmarks.
+The model builds on the flow matching framework in order to directly learn joint distributions over event times, side-stepping the autoregressive process.
+EventFlow is simple to implement and achieves a 20%-53% lower error than the nearest baseline on standard TPP benchmarks while simultaneously using fewer model calls at sampling time.
 1 Introduction
-Many stochastic processes, ranging from consumer behavior (Hernandez et al., 2017) to the occurrence of earthquakes (Ogata, 1998), are best understood as a sequence of discrete events which occur at random times.
+Many stochastic processes, ranging from the occurrence of earthquakes (Ogata, 1998) to consumer behavior (Xu et al., 2014), are best understood as a sequence of discrete events which occur at random times.
 Any observed event sequence, consisting of one or more event times, may be viewed as a draw from a temporal point process (TPP) (Daley & Vere-Jones, 2003) which characterizes the distribution over such sequences.
@@
 However, in many tasks, we are interested not only in the next event, but in the entire sequence of events which is to follow.
-While these models can achieve high likelihoods, their performance in many-step forecasting tasks can be unsatisfactory due to compounding errors arising from the autoregressive sampling procedure (Xue et al., 2022; Lüdke et al., 2023).
+Despite their popularity, their performance in many-step forecasting tasks can be unsatisfactory due to compounding errors arising from the autoregressive sampling procedure (Xue et al., 2022; Lüdke et al., 2023).
 Moreover, existing models are typically trained via a maximum likelihood procedure (see Section 3) which involves computing the CDF implied by the learned model.
@@
 See Figure 1 for an illustration.
-As our primary contribution regards the modeling of the event times themselves, we focus on unmarked point processes in this work.
 More specifically,
@@
 Our model is suitable for both unconditional generation tasks (i.e., generating draws from the underlying data TPP) and conditional generation tasks (e.g., forecasting future events given a history), and is able to forecast multiple events simultaneously.
-  • Our model provides a new perspective on modeling TPPs and sidesteps common pitfalls in existing approaches.
-In particular, EventFlow is likelihood-free, non-autoregressive, easy to sample from, and straightforward to implement.
-  • On standard benchmark datasets, EventFlow obtains strong performance on a multi-step forecasting task, outperforming the second-best model by a factor of 2.2 to 4, and matches or exceeds the performance of state-of-the-art models for unconditional generation.
+  • We provide new perspective on modeling TPPs which sidesteps common pitfalls.
+In particular, the key idea of EventFlow is to decompose the generative process into a learned event count distribution and a flow-based generative model for the joint distribution of event times.
+Our model is non-autoregressive, easy to sample from, and straightforward to implement.
+  • On standard benchmarks, we demonstrate that EventFlow obtains state-of-the-art performance on a multi-step forecasting task and matches or exceeds the performance of state-of-the-art models for unconditional generation.
+In particular, we achieve a 20%-53% reduction in forecasting error from the nearest baseline.
+Compared with existing methods, which can require hundreds of model evaluations at sampling time, we demonstrate that EventFlow can achieve competitive performance with only a single forward pass.
 Figure 1: All illustration of forecasting with our EventFlow method.
@@
 Similarly, a wide range of forms for the event decoder have also been proposed.
-The most common approach is to parametrize a conditional intensity function via a neural network.
-For instance, several authors (Mei & Eisner, 2017; Zuo et al., 2020; Zhang et al., 2020) model the conditional intensity using a parametric form inspired by the Hawkes process (Hawkes, 1971), and Du et al.
-(2016) model the (log-)conditional intensity through an affine function of the history embedding.
-Similarly, Okawa et al.
-(2019) model the conditional intensity using a mixture of Gaussian kernels.
-Most closely related to our work are approaches which use generative models as decoders.
+The most common approach is to parametrize a conditional intensity function via a neural network (Mei & Eisner, 2017; Zuo et al., 2020; Du et al., 2016).
+An alternative approach is to use generative models as decoders.
 These models often do not assume a parametric form for the decoder, enhancing their flexibility.
@@
 (2023) propose a diffusion-style model which is able to avoid autoregressive sampling via an iterative refinement procedure.
-Our work can be viewed as a novel approach for building flexible decoders for TPPs, extending flow matching to the setting of continuous-time event sequences.
-In contrast to prior work using generative models, our model is likelihood-free and non-autoregressive, achieving strong performance on long-term forecasting tasks.
+Our work can be viewed as a novel approach to building flexible decoders for TPPs, extending flow matching to the setting of continuous-time event sequences.
+In contrast previous generative approaches, our model is trained in a likelihood-free manner and is non-autoregressive.
 The work of Lüdke et al.
-(2023) is perhaps most closely related to ours, but we emphasize that the method of Lüdke et al.
-(2023) requires an involved training and sampling procedure.
+(2023) is most closely related to ours, but we emphasize that Lüdke et al.
+(2023) require an involved training and sampling procedure.
 In contrast, our method is straightforward to both implement and sample from, while simultaneously outperforming existing approaches.
@@
 This is typically achieved by either directly modeling a distribution over $t$ (Shchur et al., 2020a), or equivalently by modeling a conditional intensity function (Du et al., 2016).
-In the first approach, a conditional probability density of the form $p(t\mid\mathcal{H}_{t_{n}})$ is learned, allowing us to specify a joint distribution over event times $p(t_{1},\dots,t_{n})$ autoregressively via ${p(t_{1},\dots,t_{n})=p(t_{1})\prod_{k=2}^{n}p(t_{k}\mid\mathcal{H}_{t_{k-1}})}$.
+In the first approach, a conditional probability density of the form $p_{n}(t\mid\mathcal{H}_{t_{n}})$ is learned, allowing us to specify a joint distribution over event times $p(t_{1},\dots,t_{n})$ autoregressively via ${p_{n}(t_{1},\dots,t_{n})=p(t_{1})\prod_{k=2}^{n}p(t_{k}\mid\mathcal{H}_{t_{k-1}})}$.
 Alternatively, we may define the conditional intensity $\lambda^{*}(t):=\lambda(t\mid\mathcal{H}_{t^{-}})=p(t\mid\mathcal{H}_{t_{n}})/(1-F(t\mid\mathcal{H}_{t_{n}}))$, where $F(t\mid\mathcal{H}_{t_{n}})=\int_{t_{n}}^{t}p(s\mid\mathcal{H}_{t_{n}})\,\mathrm{d}s$ is the CDF associated with $p(t\mid\mathcal{H}_{t_{n}})$.
@@
 The likelihood may be expressed in terms of either the density or intensity via
-$L\left(\{t_{k}\}\right)=p(t_{1},\dots,t_{n})\left(1-F(T\mid\mathcal{H}_{t_{n}})\right)=\left(\prod_{k=1}^{n}\lambda^{*}(t_{k})\right)\exp\left(-\Lambda^{*}(T)\right)$ (2)
+$L\left(\{t_{k}\}_{k=1}^{n}\right)=p_{n}(t_{1},\dots,t_{n})\left(1-F(T\mid\mathcal{H}_{t_{n}})\right)=\left(\prod_{k=1}^{n}\lambda^{*}(t_{k})\right)\exp\left(-\Lambda^{*}(T)\right)$ (2)
 where the CDF term is included to indicate that no events beyond $t_{n}$ have occurred and $\Lambda^{*}(T)=\int_{0}^{T}\lambda^{*}(s)\,\mathrm{d}s$ is the total intensity.
 Autoregressive models are typically trained by maximizing this likelihood (Du et al., 2016; Mei & Eisner, 2017; Shchur et al., 2020a).
-We emphasize that this likelihood is not simply the joint event-time density $p_{n}(t_{1},\dots,t_{n})$, as the likelihood measures the fact that no events occur after $t_{n}$.
+We emphasize that this likelihood is not simply the joint event-time density $p_{n}(t_{1},\dots,t_{n})$, as the likelihood also measures the fact that no events occur between $t_{n}$ and $T$.
 It is worth noting that evaluating $L(\{t_{k}\})$ can be non-trivial in practice.
-For models which parametrize $\lambda^{*}(t)$ via a neural network (Du et al., 2016; Mei & Eisner, 2017), computing the total intensity $\Lambda^{*}(T)$ is often done via a Monte Carlo integral, requiring several forward passes of the model to evaluate $\lambda^{*}(t)$ at different values of $t$.
+For models that parametrize $\lambda^{*}(t)$ via a neural network (Du et al., 2016; Mei & Eisner, 2017), computing the total intensity $\Lambda^{*}(T)$ is often done via a Monte Carlo integral, requiring many forward passes of the model to evaluate $\lambda^{*}(t)$ at different values of $t$.
 Models which directly parametrize the density $p(t\mid\mathcal{H}_{t})$ suffer from the same drawback when computing the corresponding CDF in Equation (2).
 Moreover, some approaches, such as the diffusion-based approach of Lin et al.
-(2022), are only trained to maximize an ELBO of $p(t\mid t_{1},\dots,t_{n})$, and are thus unable to compute the proper likelihood in Equation (2).
+(2022), are only trained to maximize an ELBO of $p(t_{k}\mid t_{1},\dots,t_{k-1})$, and are thus unable to compute the proper likelihood in Equation (2).
 Sampling Autoregressive Models
-In many tasks, we are interested not only in an accurate model of the intensity (or distribution), but also sampling new event sequences from the corresponding distribution.
+In many tasks, we are interested not only in an accurate model of the intensity (or distribution), but also sampling new event sequences.
 For instance, when forecasting an event sequence, we may want to generate several forecasts in order to provide uncertainty quantification over these predictions.
-Stochastic forecasts also have the potential to enable probabilistic queries (Boyd et al., 2023; 2022), e.g., estimating hitting times (the time it will take until a specified number of events occurs) or estimating the number of events in some subwindow.
-However, sampling from existing autoregressive models can be difficult.
-For instance, the flow-based model of Shchur et al.
+However, sampling from existing models can be difficult.
+The flow-based model of Shchur et al.
 (2020a) requires a numerical approximation to the inverse of the model to perform sampling.
-Similarly, the diffusion-based approach of Lin et al.
+The diffusion-based approach of Lin et al.
 (2022) can require several hundred forward passes of the model to generate a single event time, rendering it costly when generating long sequences.
@@
 However, this algorithm has several hyperparameters to tune, is challenging to parallelize, and can be difficult for practitioners to implement (Xue et al., 2024).
+For instance, the thinning algorithm requires one to know an upper bound on the intensity, which is in practice selected heuristically.
 4 EventFlow
-Motivated by the limitations of autoregressive models, we propose EventFlow, which has a number of distinct advantages over existing approaches.
+Motivated by the limitations of autoregressive models, we propose EventFlow, which has a number of distinct advantages.
 First, EventFlow directly models the joint distribution over event times, thereby avoiding autoregression entirely.
-Second, our model is likelihood-free, avoiding the Monte Carlo estimates needed to estimate the likelihood in Equation (2) during training.
+Second, our model is likelihood-free at training time, avoiding the Monte Carlo estimates needed to estimate the likelihood in Equation (2).
 Third, sampling from our model amounts to solving an ordinary differential equation.
-This is straightforward to implement and parallelize, allowing us to avoid the difficulties of thinning-based approaches used in existing models.
+This is straightforward to implement and parallelize, allowing us to avoid the difficulties of thinning-based approaches.
+In fact, EventFlow can achieve competitive performance using only a single forward pass at sampling time.
 We build upon the flow matching (or stochastic interpolant) framework (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) to develop our model.
-We begin below by focusing on the unconditional setting, and later discuss how to extend our method for conditional generation as necessary.
-4.1 Preliminaries
+We begin below by focusing on the unconditional setting, followed by an extension for conditional generation.
+Preliminaries
 We first introduce some necessary background and notation.
-Let $\mathcal{T}=[0,T]$ be a finite-length interval.
+Let $\mathcal{T}=[0,T]$ be a finite length interval.
 The set $\Gamma$ denotes the configuration space of $\mathcal{T}$ (Albeverio et al., 1998), i.e., the set of all finite counting measures on the set $[0,T]$.
@@
 A temporal point process (TPP) on $\mathcal{T}$ is a probability distribution $\mu$ over the configuration space $\Gamma$.
-Informally, $\mu$ represents a distribution over sequences $\gamma$ living in the configuration space $\Gamma$ which constitutes the set of valid sequences.
-We use $N:\Gamma\to\mathbb{Z}_{\geq 0}$ to denote the counting functional, i.e., $N(\gamma)$ is the number of events in the TPP realization $\gamma$ [^1] This can be thoughts of in terms of the counting process, i.e.
-$N(\gamma)$ corresponds to the value of the associated counting process at the ending time $T$, or the total number of events in $\gamma$ that have occurred in the interval $[0,T]$.
-While it is common to represent TPPs as distributions over random sets of event times, in our approach it will be more convenient to represent TPPs as random measures (Kallenberg et al., 2017).
-We assume all TPP distributions are atomless (Kallenberg et al., 2017, Ch.
-1), i.e., the probability of observing an event at any singleton is zero.
-In addition, we assume all TPPs are simple (Kallenberg et al., 2017, Ch.
-2), i.e., no more than one event can occur simultaneously.
-Under these assumptions, a TPP $\mu$ can be fully characterized (Daley & Vere-Jones, 2003, Prop.
+Informally, $\mu$ represents a distribution over finite sequences $\gamma\in\Gamma$ whose events are in $\mathcal{T}$.
+We use $N:\Gamma\to\mathbb{Z}_{\geq 0}$ to denote the counting functional, i.e., $N(\gamma)$ is the number of events in the sequence $\gamma$ [^1] This can be thought of in terms of the counting process, i.e., $N(\gamma)$ corresponds to the value of the associated counting process at the ending time $T$, or the total number of events in $\gamma$ that have occurred in the interval $[0,T]$.
+While it is common to represent TPPs as distributions over random sets of event times, in our approach it will be more convenient to represent TPPs as random measures (Kallenberg, 2017).
+Under mild assumptions, a TPP $\mu$ can be fully characterized (Daley & Vere-Jones, 2003, Prop.
 5.3.II) by a probability distribution which specifies the number of events and a collection of joint densities corresponding to the event times themselves.
@@
 In other words, for any given $n\in\mathbb{Z}_{\geq 0}$, the probability of observing $n$ events in the interval $\mathcal{T}$ is $\mu(n)$, and $\mu^{n}(t^{1},\dots,t^{n})$ describes the corresponding joint distribution of event times.
-We further restrict each $\mu^{n}$ to be supported only on the ordered sets, so that we may assume $t^{1}<t^{2}<\dots<t^{n}$.
+We further restrict each $\mu^{n}$ to be supported only on the ordered sets, so that $t^{1}<t^{2}<\dots<t^{n}$.
 Let $\mu_{1}$ represent the data distribution and $\mu_{0}$ represent a reference distribution.
@@
 These two time axes are in a sense orthogonal to one another (see Figure 1).
-4.2 Balanced Couplings
-Our first step is to define a useful notion of couplings (Villani et al., 2009), allowing us to pair event sequences drawn from $\mu_{0}$ with those drawn from $\mu_{1}$.
+Balanced Couplings
+Our first step is to define a useful notion of couplings, allowing us to pair event sequences drawn from $\mu_{0}$ with those drawn from $\mu_{1}$.
 A coupling between two TPPs $\mu,\nu\in\mathbb{P}(\Gamma)$ is a joint probability measure $\rho\in\mathbb{P}(\Gamma\times\Gamma)$ over pairs of event sequences $(\gamma_{0},\gamma_{1})$ such that the marginal distributions of $\rho$ are $\mu$ and $\nu$.
@@
 In other words, balanced couplings only pair event sequences with equal numbers of events.
-We will use $\Pi_{b}(\mu,\nu)$ to denote the set of balanced couplings of $\mu,\nu$.
-We can show that the set $\Pi_{b}(\mu,\nu)$ is nonempty if and only if the event count distributions of $\mu$ and $\nu$ are equal, placing a structural constraint on the suitable choices of a reference measure.
+While we will later see how to interpolate between any two given event sequences, this coupling will allow us to decide which sequences to interpolate.
+In particular, a balanced coupling will allow us to pair sequences such that they always have the same number of events, allowing us to avoid the addition or deletion of events during both training and sampling.
+We will use $\Pi_{b}(\mu,\nu)$ to denote the set of balanced couplings.
+The following proposition shows $\Pi_{b}(\mu,\nu)$ is nonempty if and only if the event count distributions of $\mu$ and $\nu$ are equal, placing a structural constraint on the suitable choices of $\mu_{0}$.
 Proposition 1 (Existence of Balanced Couplings).
@@
 We provide a proof in Appendix B.
-To provide some intuition for the proof, in one direction, we may explicitly construct a balanced coupling (which we describe in the next paragraph).
-Conversely, if a balanced coupling exists, the event count distribution of $\mu$ (and hence $\nu$) may be obtained by sampling from the coupling and counting the number of events occurring in either coordinate.
 In practice, we follow a simple strategy for choosing both the reference TPP $\mu_{0}$ and the coupling $\rho$.
 Suppose $q$ is any given density on our state space $\mathcal{T}$, e.g., a uniform distribution.
-We take $\mu_{0}$ to be a mixed binomial process (Kallenberg et al., 2017, Ch.
+We take $\mu_{0}$ to be a mixed binomial process (Kallenberg, 2017, Ch.
 3) whose event count distribution is given by that of the data $\mu_{1}(n)$ and joint event distributions given by independent products of $q$ (up to sorting).
@@
 To draw a sample from our coupling $\rho$, we first sample a data sequence $\gamma_{1}\sim\mu_{1}$, followed by sampling $N(\gamma_{1})$ events independently from $q$ and sorting to produce a draw $\gamma_{0}\sim\mu_{0}$.
-We call this coupling the independent balanced coupling of $\mu$ and $\nu$.
-4.3 Interpolant Construction
+Interpolant Construction
 We now proceed to construct our interpolant $\eta_{s}\in\mathbb{P}(\Gamma)$.
 We will construct this path of TPPs via a local procedure which we then marginalize over a given balanced coupling.
+Here, we extend flow matching techniques (Lipman et al., 2023; Tong et al., 2024) to design our sequence-level interpolants, but we emphasize that this is only possible under a balanced coupling as the number of events is fixed.
 To that end, let $\rho$ be any balanced coupling of the reference measure $\mu_{0}$ and the data measure $\mu_{1}$, and suppose $z:=(\gamma_{0},\gamma_{1})\sim\rho$ is a draw from this coupling.
-As $\rho$ is balanced, we have that
-$\gamma_{0}=\sum_{k=1}^{n}\delta[t_{0}^{k}]\qquad\gamma_{1}=\sum_{k=1}^{n}\delta[t_{1}^{k}]$ (3)
-are both a collection of $n$ events.
+As $\rho$ is balanced, we have $\gamma_{0}=\sum_{k=1}^{n}\delta[t_{0}^{k}]$ and $\gamma_{1}=\sum_{k=1}^{n}\delta[t_{1}^{k}]$ are both a collection of $n$ events.
 As TPPs are fully characterized by their joint distributions over event times, we will henceforth describe our procedure for a fixed (but arbitrary) number of events $n$.
@@
 This defines a path $(\gamma_{s}^{z})_{s=0}^{1}$ in the configuration space $\Gamma$ which evolves the reference sample $\gamma_{0}$ into the data sample $\gamma_{1}$.
-In order to perform the marginalization step, we now lift this deterministic path $(\gamma_{s}^{z})_{s=0}^{1}\subset\Gamma$ to a path of TPP distributions $(\eta_{s}^{z})_{s=0}^{1}\subset\mathbb{P}(\Gamma)$.
+We now lift this deterministic path $(\gamma_{s}^{z})_{s=0}^{1}\subset\Gamma$ to a path of TPP distributions $(\eta_{s}^{z})_{s=0}^{1}\subset\mathbb{P}(\Gamma)$.
 We define the point process distribution $\eta_{s}^{z}\in\mathbb{P}(\Gamma)$ implicitly by adding independent Gaussian noise to each of the events in $\gamma_{s}^{z}$.
@@
 We note that this noising step is typical in flow matching models (Lipman et al., 2023; Tong et al., 2024).
-Finally, for any $s\in[0,1]$, we define the marginal TPP measure $\eta_{s}$ via $\eta_{s}=\int\eta_{s}^{z}\,\mathrm{d}\rho(z)$.
+Finally, for $s\in[0,1]$, we define the marginal TPP measure $\eta_{s}\in\mathbb{P}(\Gamma)$ via $\eta_{s}=\int\eta_{s}^{z}\,\mathrm{d}\rho(z)$.
 Observe that, by construction, the event count distribution $\eta_{s}(n)$ is given by $\mu_{1}(n)$ for all $s\in[0,1]$.
@@
 As $\sigma^{2}\downarrow 0$, it is clear that we recover a genuine interpolant (Tong et al., 2024).
-We now shift our attention to the transport of a single event $t_{s}^{k}$ for a fixed $k$.
-Through the addition of Gaussian noise, we have constructed a path of Gaussian distributions $\mathcal{N}(t_{s}^{k},\sigma^{2})$ whose mean is determined by the location of the $k$th event at the flow time $s$.
-This transport of a Gaussian can be achieved infinitesimally through the constant vector field ${v_{s}^{k}:[0,T]\to\mathbb{R}}$ given by $v_{s}^{k}(t)=t_{1}^{k}-t_{0}^{k}$ (Tong et al., 2024).
-Thus, the evolution in (5) is generated by the vector field $v_{s}^{z}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ given by
+For single event $t_{s}^{k}$, we have constructed a path of distributions $\mathcal{N}(t_{s}^{k},\sigma^{2})$.
+This transport of a Gaussian can be achieved infinitesimally through the vector field $v_{s}^{k}(t)=t_{1}^{k}-t_{0}^{k}$ (Tong et al., 2024).
+Thus, the evolution (4) is generated by the vector field $v_{s}^{z}:\mathcal{T}^{n}\to\mathbb{R}^{n}$,
 $v_{s}^{z}(\gamma)=\begin{bmatrix}v_{s}^{1},\dots,v_{s}^{n}\end{bmatrix}^{\sf T}=\begin{bmatrix}t_{1}^{1}-t_{0}^{1},&\dots,&t_{1}^{n}-t_{0}^{n}\end{bmatrix}^{\sf T}\qquad 0\leq s\leq 1.$ (5)
-Informally, we view $v_{s}^{z}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ as prescribing a constant (but different) velocity to each of the $n$ events.
-For a fixed pair $z=(\gamma_{0},\gamma_{1})$ and a given sample $\gamma_{0}^{\prime}\sim\eta_{0}^{z}$, solving the system of ordinary differential equations $\,\mathrm{d}\gamma_{s}^{\prime}=v_{s}^{z}(\gamma_{s}^{\prime})\,\mathrm{d}s$ with initial condition $\gamma_{0}^{\prime}$ will result in a collection of events which is concentrated around the true event times $\gamma_{1}$.
-Note that here, we view this differential equation as an ODE in $\mathcal{T}^{n}$.
-If we draw many samples $\gamma_{0}\sim\eta_{0}^{z}$ and solve the corresponding ODE, the distribution over events at any intermediate time $s$ will be given by $\eta_{s}^{z}$.
-In other words, the vector field $v_{s}^{z}$ generates the path of distributions $\eta_{s}^{z}$.
+In other words, flowing samples $\hat{\gamma}_{0}^{z}$ along the vector field $v_{s}^{z}$ generates the path of distributions $\eta_{s}^{z}$.
 However, this path is conditioned on $z$, and we would like to find the vector field $v_{s}$ which generates the unconditional path $\eta_{s}$.
-As is standard in flow matching (Lipman et al., 2023; Tong et al., 2024; Albergo & Vanden-Eijnden, 2023), the unconditional vector field $v_{s}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ may be obtained through a weighted marginalization procedure via
+As is standard in flow matching (Lipman et al., 2023; Tong et al., 2024; Albergo & Vanden-Eijnden, 2023), the unconditional vector field $v_{s}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ may be obtained via
 $v_{s}(\gamma)=\int v_{s}^{z}(\gamma)\frac{\,\mathrm{d}\eta_{s}^{z}}{\,\mathrm{d}\eta_{s}}(\gamma)\,\mathrm{d}\rho(z).$ (6)
@@
 We describe our approach for modeling the event count distribution in conditional tasks in the following section.
-4.4 Training, Parametrization, and Sampling
-To train the model, we would like to perform regression on the vector fields $v_{s}$ in Equation (7).
-If we knew this vector field $v_{s}$, we could draw samples from the data TPP by drawing a sample event sequence $\gamma_{0}\sim\mu_{0}$ from the reference TPP, and flowing each event along the vector field $v_{s}$.
-1 Sample $\gamma_{1}\sim\mu_{1}$, $s\sim\mathcal{U}[0,1]$, and $\epsilon\sim\mathcal{N}(0,1)$
+1 Sample $\gamma_{1}\sim\mu_{1}$, $s\sim\mathcal{U}[0,1]$, $\epsilon\sim\mathcal{N}(0,1)$
 2 $e_{\mathcal{H}}=\varnothing$
@@
 4 if forecast then
-5       Sample split time $T_{0}\in[\Delta T,T-\Delta T]$
-6       Construct history $\mathcal{H}\leftarrow\{t\in\gamma_{1}:t\leq T_{0}\}$
-7       Embed history $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H})$
-8       Construct future $\gamma_{1}\leftarrow\{t\in\gamma_{1}:T_{0}<t\leq T_{0}+\Delta T\}$
+5       Sample $T_{0}\in[\Delta T,T-\Delta T]$
+6       $\mathcal{H}\leftarrow\{t\in\gamma_{1}:t\leq T_{0}\}$
+7       $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H})$
+8       $\gamma_{1}\leftarrow\{t\in\gamma_{1}:T_{0}<t\leq T_{0}+\Delta T\}$
 9
-10Set $n\leftarrow N(\gamma_{1})$
-11 Sample $t_{0}^{1},\dots,t_{0}^{n}\sim q$ and sort to construct $\gamma_{0}$
-12 Compute $\gamma_{s}^{z}$ via $t_{s}^{k}=(1-s)t_{0}^{k}+st_{s}^{k}$
+[…]
+11 Sample and sort $t_{0}^{k}\sim q$ for $k=1,\dots,n$
+12 $\gamma_{0}\leftarrow(t_{0}^{1},\dots,t_{0}^{n})$
+13 $t_{s}^{k}\leftarrow(1-s)t_{0}^{k}+st_{1}^{k}$ for $k=1,\dots,n$
+14 $\gamma_{s}^{z}\leftarrow(t_{s}^{1},\dots,t_{s}^{n})$
 Take a gradient step on $\left\lVert\gamma_{1}-\gamma_{0}-v_{\theta}\left(\gamma_{s}^{z}+\epsilon,s,e_{\mathcal{H}}\right)\right\rVert^{2}$
-Algorithm 1 Training Step for EventFlow
-Training
-Foremost, although the marginal vector field in Equation (7) admits an analytical form, it is intractable to compute in practice as the marginal measure $\eta_{s}$ is not available.
-To overcome this, we may instead perform regression on the conditional vector fields $v_{s}^{z}$.
-Here, $v_{\theta}(\gamma_{s},s)$ will represent a neural network with parameters $\theta$ which takes in a sequence $\gamma_{s}$ of $N(\gamma_{s})=n$ event times, along with the flow time $s$.
-That is, we seek to minimize the loss
-$J(\theta)=\mathbb{E}_{s,(\gamma_{0},\gamma_{1}),\hat{\gamma}_{s}^{z}}\left[\left\lVert\gamma_{1}-\gamma_{0}-v_{\theta}(\hat{\gamma}_{s}^{z},s)\right\rVert^{2}\right]$ (8)
-which can be shown to be equal to MSE regression on the unconditional $v_{s}$ up to an additive constant not depending on $\theta$ (Lipman et al., 2023; Tong et al., 2024).
-We note here that, although the regression target $v_{s}^{z}$ is linear, the unconditional vector field $v_{s}$ will in general be nonlinear.
-In practice, this loss is estimated by uniformly sampling a flow time $s\in[0,1]$, a pair $z=(\gamma_{0},\gamma_{1})\sim\rho$ from our balanced independent coupling, and drawing a noisy interpolant $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ according to Equation (5).
-To train the model on a forecasting task, where the goal is to predict a future sequence of events conditioned on a history $\mathcal{H}$, we embed $\mathcal{H}$ into a fixed-dimensional vector representation $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$ via a learned encoder $f_{\theta}$ before providing this to the model $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$ and minimizing Equation (8).
-Note that we jointly train the encoder $f_{\theta}$ and vector field $v_{\theta}$.
-See Algorithm 1.
-Parametrization
-The second challenge is that we must learn a vector field $v_{\theta}(\gamma,s)$ in $n$ dimensions for arbitrary values of $n$.
-In other words, $v_{\theta}$ is a neural network which takes in a flow time $s\in[0,1]$ and a sequence of events $\gamma$, and must produce $N(\gamma)$ scalar outputs.
-We achieve this through an attention-based architecture, which we detail in Appendix C.
-At a high level, the flow time $s$ is transformed via a learnable embedding into a fixed-dimensional vector.
-Similarly, each event in $\gamma$ is transformed into fixed-dimensional vector via a learned embedding (which is shared across the events, but not the flow time).
-The flow-time embedding is then added to each event embedding, and the resulting sequence is passed through a standard transformer architecture (Yang et al., 2022; Vaswani, 2017), resulting in a sequence of $N(\gamma)$ vectors.
-Finally, each of these vectors is projected into one dimension via a linear layer to produce the sequence of $N(\gamma)$ velocities.
-For conditional tasks, we must also compute an encoding $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$ of the history $\mathcal{H}$.
-This is done by a separate transformer encoder, which operates in the same fashion as described in the previous paragraph, but without the use of the flow-time $s$ as an input and without the final linear projection layer.
-This embedding is fed into the intermediate layers of our velocity network via cross-attention.
-Lastly, for forecasting tasks we must also learn a model of the event count distribution $p_{\phi}(n\mid\mathcal{H})$.
-We treat this as a classification problem, where the goal is to predict the number of events $n$ occurring in the forecast window given the history $\mathcal{H}$.
-We again use an attention-based model, analogous to our velocity model, but where we aggregate the final sequence embedding by averaging and passing this through a small MLP.
-The model $p_{\phi}(n\mid\mathcal{H})$ is trained by minimizing the cross-entropy loss.
+Algorithm 1 Training EventFlow
 1 Choose a flow time discretization $10=s_{0}<s_{1}<\dots<s_{K}=$ 1
@@
 4 if forecast then
-5       Embed history $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H}$)
-6       Sample $n\sim p_{\phi}(n\mid\mathcal{H})$
+5       $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H}$)
+6       $n\sim p_{\phi}(n\mid\mathcal{H})$
 7
 8else
-9       Sample $n\sim\mu_{1}(n)$
+9       $n\sim\mu_{1}(n)$
 10
-11Sample $t_{0}^{1},\dots,t_{0}^{n}\sim q$ and sort to construct $\gamma_{0}$
+11Sample and sort $t_{0}^{k}\sim q$ for $k=1,\dots,n$
+12 $\gamma_{0}\leftarrow(t_{0}^{1},\dots,t_{0}^{n})$
 13 for $k=1,2,\dots,K$ do
@@
 15       $\gamma_{s_{k}}\leftarrow\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1},e_{\mathcal{H}})$
-15
-Algorithm 2 Sampling Step for EventFlow
+16
+return $\gamma_{0}$
+Algorithm 2 Sampling EventFlow
+Training
+If we knew $v_{s}$ in Equation (6), we could draw samples from $\mu_{1}$ by drawing a sample event sequence $\gamma_{0}\sim\mu_{0}$ from the reference TPP and flowing each event along $v_{s}$.
+Although the marginal vector field in Equation (6) admits an analytical form, it is intractable to compute in practice as the marginal measure $\eta_{s}$ is not available.
+To overcome this, we instead regress on the conditional vector fields $v_{s}^{z}$.
+Here, $v_{\theta}(\gamma_{s},s)$ will represent a neural network with parameters $\theta$ which takes in a sequence $\gamma_{s}$ along with the flow time $s$.
+We seek to minimize the loss
+$J(\theta)=\mathbb{E}_{s,(\gamma_{0},\gamma_{1}),\hat{\gamma}_{s}^{z}}\left[\left\lVert\gamma_{1}-\gamma_{0}-v_{\theta}(\hat{\gamma}_{s}^{z},s)\right\rVert^{2}\right]$ (7)
+which previous works on flow matching have shown to be equal to MSE regression on the unconditional $v_{s}$ up to an additive constant not depending on $\theta$ (Lipman et al., 2023; Tong et al., 2024).
+We estimate $J(\theta)$ by uniformly sampling a flow time $s\in[0,1]$, a pair $z=(\gamma_{0},\gamma_{1})\sim\rho$ from our balanced coupling and drawing a noisy interpolant $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ according to Equation (4).
+To train the model on a forecasting task, where the goal is to predict a future sequence of events conditioned on a history $\mathcal{H}$, we embed $\mathcal{H}$ into a fixed-dimensional vector representation $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$ via a learned encoder $f_{\theta}$ before providing this to the model $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$ and minimizing Equation (7).
+Note that we jointly train the encoder $f_{\theta}$ and vector field $v_{\theta}$.
+In this case, we must also learn a model of the event count distribution $p_{\phi}(n\mid\mathcal{H})$.
+We train $p_{\phi}(n\mid\mathcal{H})$ by minimizing a cross-entropy loss.
+In practice, we found it important to regularize this loss to encourage smoothness in $n$, which we achieve through an optimal-transport based regularizer.
+As both $v_{\theta}$ and $p_{\phi}$ must be able to operate on variable-length sequences, we use a transformer-based backbone for both models.
+However, our overall method is agnostic to this choice.
+See Algorithm 1 and Appendix C.
 Sampling
-Once the network $v_{\theta}$ is learned, we may draw samples from the model by first drawing a reference sequence $\gamma_{0}\sim\mu_{0}$ and solving the corresponding system of ODEs parametrized by $v_{\theta}$.
-More concretely, we first fix a number of events $n$ to generate.
+Once $v_{\theta}$ is learned, we may sample from the model by drawing a reference sequence $\gamma_{0}\sim\mu_{0}$ and solving the corresponding ODE parametrized by $v_{\theta}$.
+More concretely, we first fix a number of events $n$.
 When seeking to unconditionally generate new sequences from the underlying data TPP $\mu_{1}$, we simply sample $n$ from the empirical event count distribution $\mu_{1}(n)$.
-For conditional tasks, we draw $n\sim q(n\mid\mathcal{H})$ from the learned conditional distribution over event counts.
+For conditional tasks, we draw $n\sim p_{\phi}(n\mid\mathcal{H})$ from the learned conditional distribution over event counts.
 Next, we draw $n$ initial events, corresponding to $s=0$, by sampling and sorting $t_{0}^{1},\dots,t_{0}^{n}\sim q$.
@@
 $\gamma_{s_{k}}=\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1})\qquad k=1,2,\dots,K$ (8)
-where $h_{k}=s_{k}-s_{k-1}$ represents a scalar step size.
-While other choices of numerical solvers are certainly possible, we found that this simple scheme was sufficient as the model sample paths are typically close to linear.
-See Algorithm 2 for the full procedure.
+where $h_{k}=s_{k}-s_{k-1}$ is a step size. While other choices are possible, we found that this simple scheme was sufficient as the model sample paths are typically close to linear. See Algorithm 2.
 5 Experiments
 We study our proposed EventFlow model under two settings.
-The first is a conditional generative modeling task, where we seek to forecast both the number and times of future events given a history.
+The first is a conditional task, where we seek to forecast both the number of future events and their times of occurrence, over a particular horizon, given a history.
 The second is an unconditional task, where we aim to learn a representation of the underlying TPP distribution from empirical observations and generate new sequences from this distribution.
-In a sense, this second task can be viewed as a special case of the first with no observed history.
-Our overall experimental procedure is inspired by that of Lüdke et al.
-(2023).
-Datasets
+This second task can be viewed as a special case of the first with no observed history.
 We evaluate our model across a diverse set of datasets encompassing a wide range of possible point process behaviors.
-First, we use a collection of six synthetic datasets produced by Omi et al.
-(2019).
-We additionally evaluate our model on seven real-world datasets, which are a standard benchmark for modeling unmarked TPPs (Shchur et al., 2020b; Bosser & Taieb, 2023; Lüdke et al., 2023).
-See Appendix A for additional information regarding our datasets.
+We evaluate our model on seven real-world datasets, which are a standard benchmark for modeling unmarked TPPs (Shchur et al., 2020b; Bosser & Taieb, 2023; Lüdke et al., 2023).
+We also evaluate on a collection of six synthetic datasets (Omi et al., 2019).
+See Appendix A.
 Baseline Models
-As a point of comparison, we train and evaluate several baseline models.
-These models were selected as they constitute a set of diverse and highly performant models.
-For an intensity-based method, we compare against the Neural Hawkes Process (NHP) (Mei & Eisner, 2017).
+We selected a set of baselines consisting of a set of diverse and highly performant neural TPP models.
+For an intensity-based method, we compare against the Neural Hawkes Process (NHP) (Mei & Eisner, 2017), a well-known and very widely-used neural TPP model.
 We additionally compare against two intensity-free methods, namely the flow-based IFTPP model (Shchur et al., 2020a) and the diffusion-based model of Lin et al.
@@
 This is particularly challenging for the unconditional setting, where unlike forecasting, we do not have a ground-truth sequence to compare against.
-Our starting point is a metric (Xiao et al., 2017a) on the space of sequences $\Gamma$, allowing us to measure the distance between two sequences $\gamma=\sum_{k=1}^{n}\delta[t_{k}^{\gamma}]$ and $\eta=\sum_{k=1}^{m}\delta[t_{k}^{\eta}]$ with possibly different numbers of events.
-Without loss of generality, we assume $n\leq m$, so the distance is given by
-$d(\gamma,\eta)=\sum_{k=1}^{n}|t_{k}^{\gamma}-t_{k}^{\eta}|+\sum_{k=n+1}^{m}(T-t_{k}^{\eta})$ (10)
-where we recall that we assume the sequences are supported on $\mathcal{T}=[0,T]$.
-This distance can be understood either as an $L^{1}$ distance between the corresponding counting processes of $\gamma,\eta$ or as a generalization of the $1$-Wasserstein distance to measures of unequal mass.
-This metric allows us to compare two sequences of possibly unequal lengths, and will be used in our forecasting experiment.
+Our starting point is a metric (Xiao et al., 2017a) on the space of sequences $\Gamma$, allowing us to measure the distance between two sequences $\gamma=\sum_{k=1}^{n}\delta[t^{k}_{\gamma}]$ and $\eta=\sum_{k=1}^{m}\delta[t^{k}_{\eta}]$ with possibly different numbers of events.
+Without loss of generality, we assume $n\leq m$, where the distance is given by
+$d(\gamma,\eta)=\sum_{k=1}^{n}|t^{k}_{\gamma}-t^{k}_{\eta}|+\sum_{k=n+1}^{m}(T-t^{k}_{\eta})$ (9)
+where we recall that sequences are supported on $\mathcal{T}=[0,T]$.
+This distance can be understood either as an $L^{1}$ distance between the counting processes of $\gamma,\eta$ or as a generalization of the $1$-Wasserstein distance to measures of unequal mass.
 For our unconditional experiment, we require a metric which will capture the distance between the TPP distributions themselves.
-To do so, we follow the approach suggested by Shchur et al.
-(2020b) and Lüdke et al.
-(2023), and use the distance in Equation (10) to calculate an MMD (Gretton et al., 2012).
-Namely, the MMD between two point process distributions $\mu,\nu\in\mathbb{P}(\Gamma)$ is given by
-$\mathsf{MMD}(\mu,\nu)=\mathbb{E}_{\gamma,\gamma^{\prime}\sim\mu}[k(\gamma,\gamma^{\prime})]-2\mathbb{E}_{\gamma\sim\mu,\eta\sim\nu}[k(\gamma,\eta)]+\mathbb{E}_{\eta,\eta^{\prime}\sim\nu}[k(\eta,\eta^{\prime})]$ (11)
-where $k:\Gamma\to\Gamma\to\mathbb{R}_{\geq 0}$ is a specified kernel.
-In accordance with prior work (Shchur et al., 2020b; Lüdke et al., 2023), we use an exponential kernel $k(\gamma,\eta)=\exp\left(-d(\gamma,\eta)/(2\sigma^{2})\right)$ with $\sigma$ chosen to be the median distance between all sequences.
+To do so we use the distance in Equation (9) to calculate an MMD (Gretton et al., 2012; Shchur et al., 2020b).
+We use an exponential kernel $k(\gamma,\eta)=\exp\left(-d(\gamma,\eta)/(2\sigma^{2})\right)$ with $\sigma$ chosen to be the median distance between all sequences (Shchur et al., 2020b; Lüdke et al., 2023).
+Figure 2: Sequence distance (9) between the forecasted and ground-truth event sequences on a held-out test set.
+We report the mean $\pm$ one standard deviation over five random seeds.
+EventFlow (with 25 NFEs) achieves the lowest mean distance (forecasting error) for each of the 7 datasets.
 5.1 Forecasting Event Sequences
-Table 1: Sequence distance (10) between the forecasted and ground-truth event sequences on a held-out test set.
-Lower is better.
-We report the mean $\pm$ one standard deviation over five random seeds.
-The best mean distance on each dataset is indicated in bold, and the second best by an underline.
-PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
-IFTPP $2.6{\scriptstyle\pm 0.38}$ $20.4{\scriptstyle\pm 1.74}$ $43.8{\scriptstyle\pm 5.70}$ $3.4{\scriptstyle\pm 0.17}$ $1.9{\scriptstyle\pm 0.15}$ $1.4{\scriptstyle\pm 0.12}$ $2.6{\scriptstyle\pm 0.21}$
-NHP $2.2{\scriptstyle\pm 0.07}$ $29.9{\scriptstyle\pm 2.17}$ $67.6{\scriptstyle\pm 1.2}$ $3.4{\scriptstyle\pm 0.50}$ $2.6{\scriptstyle\pm 0.53}$ $1.1{\scriptstyle\pm 0.04}$ $2.1{\scriptstyle\pm 0.52}$
-Diffusion $4.5{\scriptstyle\pm 1.26}$ $20.8{\scriptstyle\pm 2.4}$ $64.5{\scriptstyle\pm 19.8}$ $3.8{\scriptstyle\pm 0.51}$ $1.8{\scriptstyle\pm 0.18}$ $1.1{\scriptstyle\pm 0.02}$ $2.6{\scriptstyle\pm 0.12}$
-Add-and-Thin $\underline{2.0}{\scriptstyle\pm 0.03}$ $\underline{13.7}{\scriptstyle\pm 0.46}$ $\underline{21.5}{\scriptstyle\pm 0.51}$ $\underline{2.2}{\scriptstyle\pm 0.06}$ $\underline{1.5}{\scriptstyle\pm 0.06}$ $\underline{0.9}{\scriptstyle\pm 0.02}$ $\underline{1.4}{\scriptstyle\pm 0.04}$
-EventFlow (ours) $\mathbf{0.5}{\scriptstyle\pm 0.01}$ $\mathbf{4.8}{\scriptstyle\pm 0.25}$ $\mathbf{6.8}{\scriptstyle\pm 0.44}$ $\mathbf{1.0}{\scriptstyle\pm 0.10}$ $\mathbf{0.4}{\scriptstyle\pm 0.02}$ $\mathbf{0.3}{\scriptstyle\pm 0.02}$ $\mathbf{0.5}{\scriptstyle\pm 0.04}$
-We first evaluate our model on a multi-step conditional forecasting task, focusing on the real-world datasets.
-To do so, we set a forecast horizon $\Delta T$ for each of our real-world datasets, and generate event sequences in the range $[T_{0},T_{0}+\Delta T]$ for some given $T_{0}$, conditioned on the history of events $\mathcal{H}_{T_{0}}$.
-Up to a shift, this means we are taking $\mathcal{T}=[0,\Delta T]$ as the support for our model TPP.
-The forecast horizon $\Delta T$ is chosen such that the window typically contains multiple events.
-At training time, we uniformly sample $T_{0}\in[\Delta T,T-\Delta T]$ and split a given data sequence $\gamma_{1}$ into a history on $[0,T_{0}]$ and a future $[T_{0},T_{0}+\Delta T]$.
-We encode the history $\mathcal{H}_{T_{0}}$ before training the model to fit the events occurring in the future.
-At testing time, we perform the same splitting procedure, sampling 50 values of $T_{0}$ for each test set sequence.
-We then forecast the sequence in $[T_{0},T_{0}+\Delta T]$ via the model and compute the distance (10) between the ground-truth and generated sequences.
-We report the results of this experiment in Table 1.
-Our proposed EventFlow method obtains the lowest average forecasting error across all datasets, obtaining an error which is 2.2 to 4 times lower than the second-best method, Add-and-Thin.
+We first evaluate our model on a multi-step forecasting task.
+We set a horizon $\Delta T$ for each of our real-world datasets and seek to generate event sequences in the range $[T_{0},T_{0}+\Delta T]$ for some given $T_{0}$, conditioned on the history $\mathcal{H}_{T_{0}}$.
+Up to a shift, this means we are taking $\mathcal{T}=[0,\Delta T]$.
+At training time, we uniformly sample $T_{0}\in[\Delta T,T-\Delta T]$.
+At test time, we sample 50 values of $T_{0}$ for each test set sequence.
+We then forecast the sequence in $[T_{0},T_{0}+\Delta T]$ and compute the distance (9) between the ground-truth and generated sequences.
+We note that the distance in Equation (9) is computed using $T_{0}+\Delta T$ rather than $T$ as the maximum event time, as using $T$ would result in a distance which is sensitive to the location of the forecasting window.
+We further normalize Equation (9) by $\Delta T$.
+This overall approach is similar to that of Lüdke et al.
+(2023).
+For all datasets, the sequences are split into train-validation-test sets, and all metrics are reported on the held-out test set.
+We report the results of these experiments in Figure 2.
+Our proposed EventFlow method obtains the lowest average forecasting error across all datasets, achieving an error which is 20%-53% lower than the nearest baseline.
 Given that the non-autoregressive models (EventFlow, Add-and-Thin) consistently outperform the autoregressive baselines, this is clear evidence that autoregressive models can struggle on multi-step predictions.
 This is especially true on the Reddit-C and Reddit-S datasets, which exhibit long sequence lengths.
+Table 1: Mean sequence distance per dataset for EventFlow as we vary the NFEs used to simulate the ODE.
+NFEs PUBG Red.C Red.S Taxi Twitter Yelp-A Yelp-M
+25 $2.0$ $15.8$ $16.0$ 3.23.2{3.2}3.2 1.41.4{1.4}1.4 1.31.3{1.3}1.3 1.91.9{1.9}1.9
+10 $2.0$ $15.8$ $15.8$ $3.1$ $1.4$ $1.3$ $1.9$
+1 $2.0$ $15.8$ $15.8$ $3.7$ $1.4$ $1.8$ $1.9$
+In Table 1, we additionally vary the number of function evaluations (NFEs) used at test time, i.e., the number of steps $K$ used in the ODE solver in Algorithm 2.
+We find that reducing $K$ results in minimal (or no) loss in performance.
+Even with only a single NFE, EventFlow is able to obtain state-of-the-art performance.
+This is because, although the vector fields (6) in EventFlow are generally non-linear, we find that our construction produces paths which are approximately so.
+In contrast, Add-and-Thin uses 100 NFEs, while the NFEs for the autoregressive models scale linearly with the number of sampled events.
+Table 2: MARE values evaluating the predicted number of events when forecasting. Mean values are reported over five random seeds.
+PUBG Red.C Red.S Taxi Twitter Yelp-A Yelp-M
+IFTPP $1.05$ $1.69$ $0.79$ $0.60$ $0.88$ $0.76$ $0.76$
+NHP $1.02$ $\underline{0.95}$ $1.00$ $0.67$ $2.48$ $0.80$ $1.07$
+Diffusion $1.95$ $1.28$ $1.12$ $0.49$ $0.66$ $0.65$ $0.72$
+Add-and-Thin $\underline{0.43}$ $0.99$ $\underline{0.38}$ $\underline{0.33}$ $\underline{0.60}$ $\mathbf{0.42}$ $\mathbf{0.46}$
+EventFlow (ours) $\mathbf{0.40}$ $\mathbf{0.70}$ $\mathbf{0.16}$ $\mathbf{0.28}$ $\mathbf{0.46}$ $\underline{0.56}$ $\underline{0.50}$
+To evaluate the predictions of event counts, ${p_{\phi}(n\mid\mathcal{H})}$, we report the mean absolute relative error (MARE) between the true and predicted event counts.
+This is reported in Table 2.
+Our model achieves strong performance on this metric as it decouples the event count prediction from the event time prediction, allowing us to train explicitly for each task.
+The baselines, on the other hand, only model $n$ implicitly.
+See Appendix E for details and standard deviations.
 5.2 Unconditional Generation of Event Sequences
-Next, we evaluate our model on an unconditional generation task, where we aim to generate new sequences from the underlying data distribution.
-This task serves as a benchmark to evaluate the methods in terms of how well they are able to fit the underlying TPP.
-Moreover, learning a general-purpose TPP prior could enable downstream tasks, such as data augmentation (Graikos et al., 2022).
-In Tables 2 and 3, we report MMD values (11) for each of the synthetic and real-world datasets.
-We tune the hyperparameters for all models and perform model selection based on the validation set MMD.
-MMDs are calculated by sampling 1,000 sequences from each trained model, and estimating Equation (11) using the generated and test set samples.
-The first row (“data”) is the MMD calculated between samples in the training and validation sets, giving us a sense of the best-case performance.
-Overall, we find that our EventFlow method (mean rank: $\mathbf{1.8}$) exhibits uniformly strong performance, obtaining either the best or second best MMD on 11 of the 13 datasets.
-This is particularly pronounced on the real-world datasets, where we obtain the lowest MMD on 5 of the 7 datasets.
-We see that IFTPP (mean rank: $\mathbf{2.1}$) is a strong baseline, obtaining results which are competitive with our method.
+Finally, we evaluate our model on an unconditional generation task, where we aim to generate new sequences from the underlying data distribution.
+In Table 3 we report MMD values for each of the synthetic and real-world datasets.
+See Appendix E for results with standard deviations.
+EventFlow (mean rank: $\mathbf{1.8}$) exhibits uniformly strong performance, obtaining either the best or second best MMD on 11 of the 13 datasets.
+While IFTPP (mean rank: $\mathbf{2.1}$) performs well on the synthetic datasets, it struggles with the more complex real-world datasets, reflecting its high bias.
 The Add-and-Thin method (mean rank: $\mathbf{2.4}$) is often similarly strong, but struggles on the SC dataset.
@@
 The diffusion baseline (mean rank: $\mathbf{4.8}$) is our weakest baseline, which is perhaps unsurprising as this model can only be trained to maximize the likelihood of a subsequent event and not the overall sequence likelihood.
-Table 2: MMDs (1e-2) between the test set and $1,000$ generated sequences on our synthetic datasets.
-Lower is better.
-We report the mean $\pm$ one standard deviation over five random seeds.
-The lowest MMD distance on each dataset is indicated in bold, and the second lowest is indicated by an underline.
-Hawkes1 Hawkes2 NSP NSR SC SR
-Data $1.3$ $1.3$ $1.8$ $3.0$ $5.7$ $1.1$
-IFTPP $\mathbf{1.5}{\scriptstyle\pm 0.4}$ $\mathbf{1.4}{\scriptstyle\pm 0.5}$ $\mathbf{2.3}{\scriptstyle\pm 0.7}$ $\underline{6.2}{\scriptstyle\pm 2.1}$ $\mathbf{5.8}{\scriptstyle\pm 0.5}$ $\mathbf{1.3}{\scriptstyle\pm 0.3}$
-NHP $\underline{1.9}{\scriptstyle\pm 0.3}$ $5.2{\scriptstyle\pm 1.6}$ $3.6{\scriptstyle\pm 1.3}$ $12.6{\scriptstyle\pm 1.8}$ $25.4{\scriptstyle\pm 11.5}$ $5.0{\scriptstyle\pm 0.7}$
-Diffusion $4.8{\scriptstyle\pm 2.7}$ $5.5{\scriptstyle\pm 3.3}$ $10.8{\scriptstyle\pm 7.5}$ $15.0{\scriptstyle\pm 3.6}$ $9.1{\scriptstyle\pm 1.8}$ $5.1{\scriptstyle\pm 2.8}$
-Add-and-Thin $\underline{1.9}{\scriptstyle\pm 0.5}$ $2.5{\scriptstyle\pm 0.3}$ $\underline{2.6}{\scriptstyle\pm 0.5}$ $7.4{\scriptstyle\pm 1.2}$ $22.5{\scriptstyle\pm 0.5}$ $2.2{\scriptstyle\pm 0.8}$
-EventFlow (ours) $\underline{1.9}{\scriptstyle\pm 0.2}$ $\underline{2.2}{\scriptstyle\pm 0.1}$ $3.8{\scriptstyle\pm 1.2}$ $\mathbf{4.2}{\scriptstyle\pm 0.5}$ $\underline{8.3}{\scriptstyle\pm 0.4}$ $\underline{1.7}{\scriptstyle\pm 0.3}$
-Table 3: MMDs (1e-2) between the test set and $1,000$ generated sequences on our real-world datasets.
-Lower is better.
-We report the mean $\pm$ one standard deviation over five random seeds.
-The lowest MMD distance on each dataset is indicated in bold, and the second lowest is indicated by an underline.
-PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
-Data $1.3$ $0.6$ $0.4$ $3.1$ $2.6$ $3.6$ $3.1$
-IFTPP $5.7{\scriptstyle\pm 1.8}$ $1.3{\scriptstyle\pm 1.2}$ $\underline{1.9}{\scriptstyle\pm 0.6}$ $5.8{\scriptstyle\pm 0.9}$ $\mathbf{2.9}{\scriptstyle\pm 0.6}$ $8.2{\scriptstyle\pm 4.7}$ $\underline{5.1}{\scriptstyle\pm 0.7}$
-NHP $7.2{\scriptstyle\pm 0.2}$ $2.2{\scriptstyle\pm 1.6}$ $22.5{\scriptstyle\pm 0.3}$ $\underline{5.0}{\scriptstyle\pm 0.1}$ $7.3{\scriptstyle\pm 0.7}$ $6.7{\scriptstyle\pm 1.5}$ $6.1{\scriptstyle\pm 2.3}$
-Diffusion $14.3{\scriptstyle\pm 6.5}$ $3.9{\scriptstyle\pm 1.2}$ $6.2{\scriptstyle\pm 3.3}$ $11.7{\scriptstyle\pm 1.8}$ $12.5{\scriptstyle\pm 1.9}$ $10.9{\scriptstyle\pm 3.8}$ $10.5{\scriptstyle\pm 5.2}$
-Add-and-Thin $\underline{2.8}{\scriptstyle\pm 0.5}$ $\underline{1.2}{\scriptstyle\pm 0.27}$ $2.7{\scriptstyle\pm 0.3}$ $5.2{\scriptstyle\pm 0.6}$ $\underline{4.8}{\scriptstyle\pm 0.4}$ $\mathbf{4.5}{\scriptstyle\pm 0.2}$ $\mathbf{3.0}{\scriptstyle\pm 0.5}$
-EventFlow (ours) $\mathbf{1.5}{\scriptstyle\pm 0.2}$ $\mathbf{0.7}{\scriptstyle\pm 0.1}$ $\mathbf{0.7}{\scriptstyle\pm 0.1}$ $\mathbf{3.5}{\scriptstyle\pm 0.1}$ $4.9{\scriptstyle\pm 0.7}$ $\underline{6.6}{\scriptstyle\pm 1.2}$ $\mathbf{3.0}{\scriptstyle\pm 0.5}$
+Table 3: MMDs (1e-2) between the test set and $1,000$ generated unconditional sequences averaged over five random seeds. EventFlow achieves the overall highest mean rank.
+H1 H2 NSP NSR SC SR PUBG Red.C Red.S Taxi Twitter YelpA YelpM
+IFTPP $\mathbf{1.5}$ $\mathbf{1.4}$ $\mathbf{2.3}$ $\underline{6.2}$ $\mathbf{5.8}$ $\mathbf{1.3}$ $5.7$ $1.3$ $1.9$ $5.8$ $\mathbf{2.9}$ $8.2$ $\underline{5.1}$
+NHP $\underline{1.9}$ $5.2$ $3.6$ $12.6$ $25.4$ $5.0$ $7.2$ $2.2$ $22.5$ $\underline{5.0}$ $7.3$ $6.7$ $6.1$
+Diffusion $4.8$ $5.5$ $10.8$ $15.0$ $9.1$ $5.1$ $14.3$ $3.9$ $6.2$ $11.7$ $12.5$ $10.9$ $10.5$
+Add-Thin $\underline{1.9}$ $2.5$ $\underline{2.6}$ $7.4$ $22.5$ $2.2$ $\underline{2.8}$ $\underline{1.2}$ $2.7$ $5.2$ $\underline{4.8}$ $\mathbf{4.5}$ $\mathbf{3.0}$
+EventFlow (25 NFEs) $\underline{1.9}$ $\underline{2.2}$ $3.8$ $\mathbf{4.2}$ $\underline{8.3}$ $\underline{1.7}$ $\mathbf{1.5}$ $\mathbf{0.7}$ $\mathbf{0.7}$ $\mathbf{3.5}$ $4.9$ $\underline{6.6}$ $\mathbf{3.0}$
 6 Conclusion
-In this work, we propose EventFlow, a non-autoregressive and likelihood-free generative model for temporal point processes.
+In this work, we propose EventFlow, a non-autoregressive generative model for temporal point processes.
 We demonstrate that EventFlow is able to achieve state-of-the-art results on a multi-step forecasting task and strong performance on unconditional generation.
-There are several directions in which our work could be extended.
-First, we do not explicitly enforce that the support of our model TPP is $[0,T]$.
-This would necessitate moving beyond the Gaussian setup, which is non-trivial to carry out.
-Second, more sophisticated approaches to learning the event count distribution $p_{\phi}(n\mid\mathcal{H})$ could lead to improved performance.
-Overall, we believe that our work lays a foundation for future flow-based TPPs, and exploring their application in tasks such as imputation, marked TPPs, and querying (Boyd et al., 2023) are exciting directions.
-Reproducibility Statement
-Our experiments were conducted with fixed random seeds and tested to ensure reproducibility. After publication, we will open-source our code to facilitate reproducibility.
+Our work lays a foundation for future flow-based TPPs, and exploring applications in tasks like imputation, marked TPPs, and querying (Boyd et al., 2023) are exciting directions.
+Limitations and Outlook
+In this work, we have focused on modeling the event times, and thus do not experiment with marked TPPs or spatiotemporal TPPs.
+Extending EventFlow to these tasks is a promising directions.
+We also note that we do not explicitly enforce that the support of our model is $[0,T]$, which would necessitate moving beyond the Gaussian setting.
+Moreover, our baseline methods use the neural architectures proposed in the original works, which vary across the methods.
+Carefully studying the effects of these architecture choices is an important direction.
 Acknowledgments
@@
 In Table 4, we report the number of sequences in each dataset, some basic statistics regarding the number of events in each sequence, and their support $[0,T]$ and chosen forecast window $\Delta T$.
-In all datasets, we use 60% of the data for training, 20% for validation, and the remaining 20% for testing.
+In all datasets, we use 60% of the sequences for training, 20% for validation, and the remaining 20% for testing.
 Synthetic Datasets
@@
 We note that PUBG has the largest number of sequences (which can lead to slow training), and the Reddit-C and Reddit-S datasets have long sequences (which can lead to slow training and high memory costs).
+For selecting the forecasting horizon $\Delta T$, we follow the choice made in Lüdke et al.
+(2023) to facilitate comparison and reproducibility.
+We note the initial times $T_{0}\in[\Delta T,T-\Delta T]$ are selected to be at least $\Delta T$ to avoid forecasting sequences with no observations in the history.
 Table 4: Some basic summary statistics of the datasets we consider in this work.
@@
 Appendix B Proofs
+In this section, we provide a proof of Proposition 1, showing a necessary and sufficient condition for the existence of balanced couplings.
 See 1
@@
 Here, we provide additional details regarding the parametrization and training of our EventFlow model.
-In general, our model is based on the transformer architecture (Vaswani, 2017; Yang et al., 2022), due to its general ability to handle variable length inputs and outputs, high flexibility, and ability to incorporate long-range interactions.
+In general, our model is based on the transformer architecture (Vaswani et al., 2017; Yang et al., 2022), due to its general ability to handle variable length inputs and outputs, high flexibility, and ability to incorporate long-range interactions.
+We emphasize that this is an effective architecture choice, but our method is not necessarily tied to this architecture, i.e., other choices like CNNs or RNNs are certainly possible here.
 In all settings, our reference measure $\mu_{0}$ is specified with $q=\mathcal{N}(0,I)$.
+All models (including the baselines) were trained on a cluster of six NVIDIA A6000 GPUs with 24Gb of VRAM.
+No hardware parallelization was used, i.e., each model was only trained on a single GPU.
 Model Parametrization
-For our unconditional model, we first embed the sequence times $\gamma_{s}$, the flow-time $s$, and the sequence position indices using sinusoidal embeddings followed by an additional linear layer.
+Our unconditional model takes in a sequence $\gamma_{s}=(t^{1}_{s},\dots,t^{n}_{s})$ and flow time $s$.
+We first embed the sequence times $t^{k}$, the flow-time $s$, and the sequence position indices.
+These position indices are handled by sinusoidal embeddings followed by an additional linear layer.
 There are three linear layers in total – one for the flow time, one shared across the sequence times, and one for the position indices.
-These embeddings are added together to create a representation of the sequence, and we apply a standard transformer to this sequence to produce a sequence of vectors of length $N(\gamma_{s})$.
+These embeddings are added together to create a representation for each element of the sequence, and we apply a standard transformer to this sequence to produce a sequence of vectors of length $N(\gamma_{s})$.
 Finally, each of these vectors is projected to one dimension via a final linear layer with shared weights to produce the vector field $v_{\theta}(\gamma_{s},s)$.
@@
 Our architecture for predicting the number of future events given a history, i.e.
-$p(n\mid\mathcal{H})$, is again based on the transformer decoder, sharing the same overall architecture as our unconditional model.
+$p_{\phi}(n\mid\mathcal{H})$, is again based on the transformer decoder, sharing the same overall architecture as our unconditional model.
 However, the key difference is that we instead take a mean of the final sequence embeddings before passing this through a small MLP to produce the final logit.
 See Figure 5.
+We note that this requires specifying a maximum value $N_{\max}$, which we set as the maximum sequence length seen in the training data.
+In early experiments, we also parametrized the model not through a logit but rather as a mixture of Poissons or a mixture of Negative Binomials.
+While this no longer has the limitation of assuming a value for $N_{\max}$, we found that this approach yielded worse results.
 Training and Tuning
@@
 All sequences are generated on this normalized scale, prior to re-scaling the sequence back to the original data range before evaluation.
-Our model is trained with the Adam (Kingma & Ba, 2015) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for $30,000$ steps with a cosine scheduler (Loshchilov & Hutter, 2017), which cycled every $10,000$ steps.
+Our models are trained with the Adam (Kingma & Ba, 2015) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for $30,000$ steps with a cosine scheduler (Loshchilov & Hutter, 2017), which cycled every $10,000$ steps.
 Final hyperparameters were selected by best performance on the validation dataset achieved at any point during the training, where models were evaluated 10 times throughout their training.
@@
 We use $6$ transformer layers, $8$ attention heads, and an embedding dimension of $512$ across all settings, except for the Reddit-C and Reddit-S datasets where we use $4$ heads and an embedding dimension of $128$ due to the increased memory cost of these datasets.
+To train the event count model $p_{\phi}(n\mid\mathcal{H})$, we seek to minimize the regularized cross-entropy loss
+$\mathcal{L}(\phi)=\mathbb{E}_{n,\mathcal{H}}\left[-\log p_{\phi}(n\mid\mathcal{H})+\frac{\alpha}{N_{\max}}\sum_{k=1}^{N_{\max}}p_{\phi}(k\mid\mathcal{H})(n-k)^{2}\right]$ (19)
+where $\mathcal{H}$ is a given history and $n$ is the number of events in $\mathcal{T}$ following this history.
+Note that this is a cross-entropy loss where the regularization term can be viewed as a squared optimal transport distance between the distribution $p_{\phi}(n\mid\mathcal{H})$ and a delta distribution $\delta[n]$ at the true $n$.
+This regularization term encourages the values of $p_{\phi}(n\mid\mathcal{H})$ to be smooth in $n$, as it penalizes predictions which are far fronm the true $n$ more heavily than those which are nearby.
+We searched for values of $\alpha$, the weight associated with the OT-loss, over the set $\{0,1/N_{\max},10/N_{\max},100/N_{\max},1000/N_{\max}\}$.
 Figure 3: Overview of our model architecture for unconditional generation.
@@
 As in our other architectures, the events are embedded and passed through a transformer.
-Here, the final sequence embedding output by the transformer is averaged and passed through an additional residual MLP with three layers to produce the logit corresponding to $p(n\mid\mathcal{H})$.
+Here, the final sequence embedding output by the transformer is averaged and passed through an additional residual MLP with three layers to produce the logit corresponding to $p_{\phi}(n\mid\mathcal{H})$.
 Table 5: The best hyperparameter settings found for the vector field $v_{\theta}$ in our EventFlow method on the unconditional generation task.
@@
 Yelp-Miss. $10^{-3}$ $512$ $2048$ $8$ $6$
-Table 7: The best hyperparameter settings found for the event count predictor $p(n\mid\mathcal{H})$ in our EventFlow method on the forecasting task.
-Learning Rate Emb. Dim. MLP Dim. Heads Transformer Layers
-PUBG $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
-Reddit-C $10^{-3}$ $128$ $256$ $4$ $6$
-Reddit-S $10^{-3}$ $128$ $256$ $4$ $6$
-Taxi $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
-Twitter $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
-Yelp-Airport $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
-Yelp-Miss. $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
+Table 7: The best hyperparameter settings found for the event count predictor $p_{\phi}(n\mid\mathcal{H})$ in our EventFlow method on the forecasting task.
+Learning Rate $\alpha\times N_{max}$ Emb. Dim. MLP Dim. Heads Transformer Layers
+PUBG $10^{-3}$ $1.0$ $512$ $2048$ $8$ $6$
+Reddit-C $10^{-3}$ $1000.0$ $128$ $256$ $4$ $6$
+Reddit-S $10^{-3}$ $1000.0$ $128$ $256$ $4$ $6$
+Taxi $10^{-3}$ $1000.0$ $512$ $2048$ $8$ $6$
+Twitter $10^{-3}$ $1000.0$ $512$ $2048$ $8$ $6$
+Yelp-Airport $10^{-3}$ $1000.0$ $512$ $2048$ $8$ $6$
+Yelp-Miss. $10^{-3}$ $100.0$ $512$ $2048$ $8$ $6$
 Appendix D Additional Details on Baselines
@@
 Our best hyperparameters are reported in Table 10 and Table 11.
-Diffusion
-Our diffusion baseline is based on the implementation of Lin et al.
-(2022), and our decoder model architecture is taken directly from the code of Lin et al.
-(2022).
-[^4] URL: https://github.com/EDAPINENUT/GNTPP At a high level, this model is a discrete-time diffusion model (Ho et al., 2020) trained to generate a single inter-arrival time given a history embedding.
-Note that as the likelihood is not available in diffusion models, the CDF in the likelihood in Equation (2) is not tractable.
-Instead, the model is trained by maximizing an ELBO of only the subsequent inter-arrival time.
-In preliminary experiments, we found that the codebase provided by Lin et al.
-(2022) often produced NaN values during sampling, prompting us to make several changes.
-First, we use the RNN encoder from Shchur et al.
-(2020a), i.e.
-the same encoder as the IFTPP baseline, to reduce the memory requirements of the model.
-Second, we do not log-scale the inter-arrival times as suggested by Lin et al.
-(2022), as we found that this often led to overflow and underflow issues at sampling time.
-Third, we do not normalize the data via standardization (i.e., subtracting off the mean inter-arrival time and dividing by the standard deviation), but rather, we scale the inter-arrival times so that they are in the bounded range $[-1,1]$.
-This is aligned with standard diffusion implementations (Ho et al., 2020), and allows us to perform clipping at sampling time to avoid the accumulation of errors.
-With these changes, our diffusion baseline is competitive, and able to obtain stronger results than previous work has reported (Lüdke et al., 2023).
-We use $1000$ diffusion steps and the cosine beta schedule (Nichol & Dhariwal, 2021), and we train the model on the simplified $\epsilon$-prediction loss of Ho et al.
-(2020).
-We train for $1,000$ epochs with early stopping based on the validation set loss.
-To tune this baseline, we performed a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$, weight decays in $\{0,10^{-6},10^{-5},10^{-4}\}$, history embedding dimensions $\{32,64,128\}$, and layer numbers $\{2,4,6\}$.
-Our best hyperparameters can be found in Table 12 and Table 13.
-Add-and-Thin
-We compare to the Add-and-Thin model of Lüdke et al.
-(2023) as a recently proposed non-autoregressive baseline.
-We directly run the code provided by the authors without additional modifications.
-[^5] URL: https://github.com/davecasp/add-thin We do, however, perform a slightly larger hyperparameter sweep than Lüdke et al.
-(2023), in order to ensure a fair comparison between the methods considered.
-We train for $1,000$ epochs with early stopping on the validation loss.
-Tuning is performed via a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$ and number of mixture components in $\{8,16,32,64\}$.
-We choose to tune only these hyperparameters in order to follow the implementation provided by the authors.
-Our best hyperparameters can be found in Table 14 and Table 15.
 Table 10: The best hyperparameter settings found for NHP on the unconditional generation task.
@@
 Yelp-Miss. $10^{-2}$ $64$
+Diffusion
+Our diffusion baseline is based on the implementation of Lin et al.
+(2022), and our decoder model architecture is taken directly from the code of Lin et al.
+(2022).
+[^4] URL: https://github.com/EDAPINENUT/GNTPP At a high level, this model is a discrete-time diffusion model (Ho et al., 2020) trained to generate a single inter-arrival time given a history embedding.
+Note that as the likelihood is not available in diffusion models, the CDF in the likelihood in Equation (2) is not tractable.
+Instead, the model is trained by maximizing an ELBO of only the subsequent inter-arrival time.
+In preliminary experiments, we found that the codebase provided by Lin et al.
+(2022) often produced NaN values during sampling, prompting us to make several changes.
+First, we use the RNN encoder from Shchur et al.
+(2020a), i.e.
+the same encoder as the IFTPP baseline, to reduce the memory requirements of the model.
+Second, we do not log-scale the inter-arrival times as suggested by Lin et al.
+(2022), as we found that this often led to overflow and underflow issues at sampling time.
+Third, we do not normalize the data via standardization (i.e., subtracting off the mean inter-arrival time and dividing by the standard deviation), but rather, we scale the inter-arrival times so that they are in the bounded range $[-1,1]$.
+This is aligned with standard diffusion implementations (Ho et al., 2020), and allows us to perform clipping at sampling time to avoid the accumulation of errors.
+With these changes, our diffusion baseline is competitive, and able to obtain stronger results than previous work has reported (Lüdke et al., 2023).
+We use $1000$ diffusion steps and the cosine beta schedule (Nichol & Dhariwal, 2021), and we train the model on the simplified $\epsilon$-prediction loss of Ho et al.
+(2020).
+We train for $1,000$ epochs with early stopping based on the validation set loss.
+To tune this baseline, we performed a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$, weight decays in $\{0,10^{-6},10^{-5},10^{-4}\}$, history embedding dimensions $\{32,64,128\}$, and layer numbers $\{2,4,6\}$.
+Our best hyperparameters can be found in Table 12 and Table 13.
 Table 12: The best hyperparameter settings found for diffusion on the unconditional generation task.
@@
 Yelp-Miss. $10^{-3}$ $10^{-5}$ $32$ $4$
+Add-and-Thin
+We compare to the Add-and-Thin model of Lüdke et al.
+(2023) as a recently proposed non-autoregressive baseline.
+We directly run the code provided by the authors without additional modifications.
+[^5] URL: https://github.com/davecasp/add-thin We do, however, perform a slightly larger hyperparameter sweep than Lüdke et al.
+(2023), in order to ensure a fair comparison between the methods considered.
+We train for $1,000$ epochs with early stopping on the validation loss.
+Tuning is performed via a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$ and number of mixture components in $\{8,16,32,64\}$.
+We choose to tune only these hyperparameters in order to follow the implementation provided by the authors.
+Our best hyperparameters can be found in Table 14 and Table 15.
 Table 14: The best hyperparameter settings found for Add-and-Thin on the unconditional generation task.
@@
 Yelp-Miss. $10^{-3}$ $16$
+Appendix E Additional Experiments
+This section contains additional empirical evaluations, supplementing Section 5.
+E.1 MARE
+In Table 16, we evaluate the performance of the various models only in terms of the predicted number of events in the forecast. To do so, we measure the mean absolute relative error (MARE) given by
+$\mathrm{MARE}=\mathbb{E}_{n,\mathcal{H},\hat{n}\sim p_{\phi}(n\mid\mathcal{H})}\left|\frac{\hat{n}-n}{n}\right|$ (20)
+where $n$ represents the true number of points in a sequence following a history $\mathcal{H}$ and $\hat{n}\sim p_{\phi}(n\mid\mathcal{H})$ represents the predicted number of points.
+The expectation is estimated empirically on the testing set.
+As our method directly predicts the number of events $n$ by sampling from the learned distribution $p_{\phi}(n\mid\mathcal{H})$, this serves as a direct evaluation of this component of our model.
+We note that the baselines only model $n$ implicitly.
+For example, the autoregressive models sample new events until an event is generated outside the support $\mathcal{T}$.
+As an ablation, we also report the MARE values for our method without regularization, i.e., setting $\alpha=0$.
+We see that the regularization significantly improves the event count predictions.
+Table 16: MARE values evaluating the predicted number of events when forecasting.
+Mean values $\pm$ one standard deviation are reported over five random seeds.
+The lowest MARE on each dataset is indicated and bold, and the second lowest is indicated by an underline.
+PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
+IFTPP $1.05{\scriptstyle\pm 0.14}$ $1.69{\scriptstyle\pm 0.39}$ $0.79{\scriptstyle\pm 0.20}$ $0.60{\scriptstyle\pm 0.11}$ $0.88{\scriptstyle\pm 0.08}$ $0.76{\scriptstyle\pm 0.07}$ $0.76{\scriptstyle\pm 0.05}$
+NHP $1.02{\scriptstyle\pm 0.08}$ $\underline{0.95}{\scriptstyle\pm 0.01}$ $1.00{\scriptstyle\pm 0.0004}$ $0.67{\scriptstyle\pm 0.11}$ $2.48{\scriptstyle\pm 0.40}$ $0.80{\scriptstyle\pm 0.22}$ $1.07{\scriptstyle\pm 0.34}$
+Diffusion $1.95{\scriptstyle\pm 0.48}$ $1.28{\scriptstyle\pm 0.09}$ $1.12{\scriptstyle\pm 0.56}$ $0.49{\scriptstyle\pm 0.07}$ $0.66{\scriptstyle\pm 0.04}$ $0.65{\scriptstyle\pm 0.07}$ $0.72{\scriptstyle\pm 0.07}$
+Add-and-Thin $\underline{0.43}{\scriptstyle\pm 0.01}$ $0.99{\scriptstyle\pm 0.10}$ $\underline{0.38}{\scriptstyle\pm 0.01}$ $\underline{0.33}{\scriptstyle\pm 0.02}$ $\underline{0.60}{\scriptstyle\pm 0.02}$ $\mathbf{0.42}{\scriptstyle\pm 0.01}$ $\mathbf{0.46}{\scriptstyle\pm 0.03}$
+EventFlow (ours) $\mathbf{0.40}{\scriptstyle\pm 0.01}$ $\mathbf{0.70}{\scriptstyle\pm 0.07}$ $\mathbf{0.16}{\scriptstyle\pm 0.01}$ $\mathbf{0.28}{\scriptstyle\pm 0.01}$ $\mathbf{0.46}{\scriptstyle\pm 0.03}$ $\underline{0.56}{\scriptstyle\pm 0.06}$ $\underline{0.50}{\scriptstyle\pm 0.02}$
+EventFlow ($\alpha=0$) $0.69{\scriptstyle\pm 0.17}$ $2.01{\scriptstyle\pm 0.40}$ $0.26{\scriptstyle\pm 0.01}$ $0.47{\scriptstyle\pm 0.03}$ $1.23{\scriptstyle\pm 0.07}$ $0.66{\scriptstyle\pm 0.03}$ $0.80{\scriptstyle\pm 0.05}$
+E.2 MMDs
+In Tables 17 and 18, we report the MMD values appearing in the unconditional experiment (i.e., Table 3 in the main paper) with standard deviations.
+These are omitted from the main paper for the sake of space.
+Table 17: MMDs (1e-2) between the test set and $1,000$ generated sequences on our real-world datasets.
+Lower is better.
+We report the mean $\pm$ one standard deviation over five random seeds.
+The lowest MMD distance on each dataset is indicated in bold, and the second lowest is indicated by an underline.
+PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
+Data $1.3$ $0.6$ $0.4$ $3.1$ $2.6$ $3.6$ $3.1$
+IFTPP $5.7{\scriptstyle\pm 1.8}$ $1.3{\scriptstyle\pm 1.2}$ $\underline{1.9}{\scriptstyle\pm 0.6}$ $5.8{\scriptstyle\pm 0.9}$ $\mathbf{2.9}{\scriptstyle\pm 0.6}$ $8.2{\scriptstyle\pm 4.7}$ $\underline{5.1}{\scriptstyle\pm 0.7}$
+NHP $7.2{\scriptstyle\pm 0.2}$ $2.2{\scriptstyle\pm 1.6}$ $22.5{\scriptstyle\pm 0.3}$ $\underline{5.0}{\scriptstyle\pm 0.1}$ $7.3{\scriptstyle\pm 0.7}$ $6.7{\scriptstyle\pm 1.5}$ $6.1{\scriptstyle\pm 2.3}$
+Diffusion $14.3{\scriptstyle\pm 6.5}$ $3.9{\scriptstyle\pm 1.2}$ $6.2{\scriptstyle\pm 3.3}$ $11.7{\scriptstyle\pm 1.8}$ $12.5{\scriptstyle\pm 1.9}$ $10.9{\scriptstyle\pm 3.8}$ $10.5{\scriptstyle\pm 5.2}$
+Add-and-Thin $\underline{2.8}{\scriptstyle\pm 0.5}$ $\underline{1.2}{\scriptstyle\pm 0.27}$ $2.7{\scriptstyle\pm 0.3}$ $5.2{\scriptstyle\pm 0.6}$ $\underline{4.8}{\scriptstyle\pm 0.4}$ $\mathbf{4.5}{\scriptstyle\pm 0.2}$ $\mathbf{3.0}{\scriptstyle\pm 0.5}$
+EventFlow (ours) $\mathbf{1.5}{\scriptstyle\pm 0.2}$ $\mathbf{0.7}{\scriptstyle\pm 0.1}$ $\mathbf{0.7}{\scriptstyle\pm 0.1}$ $\mathbf{3.5}{\scriptstyle\pm 0.1}$ $4.9{\scriptstyle\pm 0.7}$ $\underline{6.6}{\scriptstyle\pm 1.2}$ $\mathbf{3.0}{\scriptstyle\pm 0.5}$
+Table 18: MMDs (1e-2) between the test set and $1,000$ generated sequences on our synthetic datasets.
+Lower is better.
+We report the mean $\pm$ one standard deviation over five random seeds.
+The lowest MMD distance on each dataset is indicated in bold, and the second lowest is indicated by an underline.
+Hawkes1 Hawkes2 NSP NSR SC SR
+Data $1.3$ $1.3$ $1.8$ $3.0$ $5.7$ $1.1$
+IFTPP $\mathbf{1.5}{\scriptstyle\pm 0.4}$ $\mathbf{1.4}{\scriptstyle\pm 0.5}$ $\mathbf{2.3}{\scriptstyle\pm 0.7}$ $\underline{6.2}{\scriptstyle\pm 2.1}$ $\mathbf{5.8}{\scriptstyle\pm 0.5}$ $\mathbf{1.3}{\scriptstyle\pm 0.3}$
+NHP $\underline{1.9}{\scriptstyle\pm 0.3}$ $5.2{\scriptstyle\pm 1.6}$ $3.6{\scriptstyle\pm 1.3}$ $12.6{\scriptstyle\pm 1.8}$ $25.4{\scriptstyle\pm 11.5}$ $5.0{\scriptstyle\pm 0.7}$
+Diffusion $4.8{\scriptstyle\pm 2.7}$ $5.5{\scriptstyle\pm 3.3}$ $10.8{\scriptstyle\pm 7.5}$ $15.0{\scriptstyle\pm 3.6}$ $9.1{\scriptstyle\pm 1.8}$ $5.1{\scriptstyle\pm 2.8}$
+Add-and-Thin $\underline{1.9}{\scriptstyle\pm 0.5}$ $2.5{\scriptstyle\pm 0.3}$ $\underline{2.6}{\scriptstyle\pm 0.5}$ $7.4{\scriptstyle\pm 1.2}$ $22.5{\scriptstyle\pm 0.5}$ $2.2{\scriptstyle\pm 0.8}$
+EventFlow (ours) $\underline{1.9}{\scriptstyle\pm 0.2}$ $\underline{2.2}{\scriptstyle\pm 0.1}$ $3.8{\scriptstyle\pm 1.2}$ $\mathbf{4.2}{\scriptstyle\pm 0.5}$ $\underline{8.3}{\scriptstyle\pm 0.4}$ $\underline{1.7}{\scriptstyle\pm 0.3}$
+E.3 MSEs
+In Table 19, we evaluate the performance of our model when forecasting only a single subsequent event.
+That is, given a history $\mathcal{H}$, we evaluate the MSE between the first true event time following this history and the first event time generated by each model conditioned on $\mathcal{H}$.
+The results are reported in Table 19.
+Generally, all of the methods show similar results on this metric, despite there being clear differences between methods on the multi-step task.
+We believe this serves to further highlight the necessity of moving beyond single-step prediction tasks.
+Table 19: MSE values evaluating one-step-ahead forecasting performance.
+Mean values $\pm$ one standard deviation are reported over five random seeds.
+The lowest MSE on each dataset is indicated and bold, and the second lowest is indicated by an underline.
+PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
+IFTPP $0.85{\scriptstyle\pm 0.05}$ $\underline{0.32}{\scriptstyle\pm 0.03}$ $0.0047{\scriptstyle\pm 0.0006}$ $0.22{\scriptstyle\pm 0.03}$ $1.74{\scriptstyle\pm 0.10}$ $1.24{\scriptstyle\pm 0.16}$ $1.11{\scriptstyle\pm 0.17}$
+NHP $0.89{\scriptstyle\pm 0.09}$ $0.53{\scriptstyle\pm 0.24}$ $\mathbf{0.0022}{\scriptstyle\pm 0.0007}$ $0.31{\scriptstyle\pm 0.12}$ $2.00{\scriptstyle\pm 0.30}$ $1.30{\scriptstyle\pm 0.26}$ $\underline{1.03}{\scriptstyle\pm 0.35}$
+Diffusion $\mathbf{0.61}{\scriptstyle\pm 0.10}$ $0.33{\scriptstyle\pm 0.04}$ $\underline{0.0037}{\scriptstyle\pm 0.0012}$ $0.23{\scriptstyle\pm 0.14}$ $\mathbf{1.30}{\scriptstyle\pm 0.21}$ $\mathbf{0.86}{\scriptstyle\pm 0.18}$ $\mathbf{0.92}{\scriptstyle\pm 0.20}$
+Add-and-Thin $0.86{\scriptstyle\pm 0.05}$ $\mathbf{0.30}{\scriptstyle\pm 0.04}$ $0.0043{\scriptstyle\pm 0.0007}$ $\underline{0.21}{\scriptstyle\pm 0.03}$ $\underline{1.53}{\scriptstyle\pm 0.14}$ $1.16{\scriptstyle\pm 0.16}$ $1.20{\scriptstyle\pm 0.14}$
+EventFlow (25 NFEs) $\underline{0.68}{\scriptstyle\pm 0.02}$ $1.06{\scriptstyle\pm 0.09}$ $0.028{\scriptstyle\pm 0.0015}$ $\mathbf{0.19}{\scriptstyle\pm 0.01}$ $2.30{\scriptstyle\pm 0.22}$ $\underline{0.90}{\scriptstyle\pm 0.05}$ $1.30{\scriptstyle\pm 0.03}$
+EventFlow (10 NFEs) ${0.60}{\scriptstyle\pm 0.02}$ $1.01{\scriptstyle\pm 0.31}$ $0.0113{\scriptstyle\pm 0.0016}$ ${0.16}{\scriptstyle\pm 0.01}$ $1.93{\scriptstyle\pm 0.07}$ ${0.77}{\scriptstyle\pm 0.05}$ ${1.14}{\scriptstyle\pm 0.07}$
+EventFlow (1 NFE) $0.65{\scriptstyle\pm 0.16}$ $1.30{\scriptstyle\pm 0.61}$ $0.0445{\scriptstyle\pm 0.0104}$ $0.18{\scriptstyle\pm 0.04}$ $1.52{\scriptstyle\pm 0.17}$ $0.95{\scriptstyle\pm 0.17}$ $1.03{\scriptstyle\pm 0.24}$
+E.4 Forecasting TPPs
+In Table 20, we report the sequence distance values appearing in the forecasting experiment (i.e., Figure 2 and Table 1 in the main paper) with standard deviations.
+We additionally report an ablation where we use the true value of $n$, rather than sampling $n\sim p_{\phi}(n\mid\mathcal{H})$.
+This serves to measure how much room for improvement remains from the event count predictor.
+In general, we see that this benchmark is still not saturated, with further gains being possible especially on datasets with long sequences (e.g., the Reddit-C and Reddit-S datasets).
+Table 20: Sequence distance (9) between the forecasted and ground-truth event sequences on a held-out test set.
+Lower is better.
+We report the mean $\pm$ one standard deviation over five random seeds.
+The lowest mean distance on each dataset is indicated in bold, and the second lowest by an underline.
+PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
+IFTPP $4.2{\scriptstyle\pm 0.7}$ $25.6{\scriptstyle\pm 2.3}$ $61.2{\scriptstyle\pm 3.2}$ $5.1{\scriptstyle\pm 0.4}$ $2.9{\scriptstyle\pm 0.2}$ $2.1{\scriptstyle\pm 0.2}$ $3.4{\scriptstyle\pm 0.2}$
+NHP $2.8{\scriptstyle\pm 0.1}$ $31.0{\scriptstyle\pm 1.4}$ $95.7{\scriptstyle\pm 0.7}$ $4.5{\scriptstyle\pm 0.3}$ $3.4{\scriptstyle\pm 0.5}$ $\underline{1.8}{\scriptstyle\pm 0.1}$ $\underline{3.0}{\scriptstyle\pm 0.2}$
+Diffusion $5.4{\scriptstyle\pm 1.2}$ $25.7{\scriptstyle\pm 0.9}$ $80.3{\scriptstyle\pm 11.4}$ $4.6{\scriptstyle\pm 0.7}$ $\underline{2.4}{\scriptstyle\pm 0.2}$ $\underline{1.8}{\scriptstyle\pm 0.1}$ $3.3{\scriptstyle\pm 0.7}$
+Add-and-Thin $\underline{2.5}{\scriptstyle\pm 0.04}$ $\underline{22.2}{\scriptstyle\pm 4.6}$ $\underline{34.3}{\scriptstyle\pm 0.4}$ $\underline{3.7}{\scriptstyle\pm 0.1}$ $3.1{\scriptstyle\pm 0.2}$ $\underline{1.8}{\scriptstyle\pm 0.1}$ $\underline{3.0}{\scriptstyle\pm 0.2}$
+EventFlow (25 NFEs) $\mathbf{2.0}{\scriptstyle\pm 0.03}$ $\mathbf{15.8}{\scriptstyle\pm 2.7}$ $\mathbf{16.0}{\scriptstyle\pm 0.2}$ $\mathbf{3.2}{\scriptstyle\pm 0.1}$ $\mathbf{1.4}{\scriptstyle\pm 0.01}$ $\mathbf{1.3}{\scriptstyle\pm 0.02}$ $\mathbf{1.9}{\scriptstyle\pm 0.1}$
+EventFlow (10 NFEs) $2.0{\scriptstyle\pm 0.03}$ $15.8{\scriptstyle\pm 2.7}$ $15.8{\scriptstyle\pm 0.2}$ $3.1{\scriptstyle\pm 0.1}$ $1.3{\scriptstyle\pm 0.02}$ $1.3{\scriptstyle\pm 0.1}$ $1.9{\scriptstyle\pm 0.1}$
+EventFlow (1 NFE) $2.0{\scriptstyle\pm 0.01}$ $15.8{\scriptstyle\pm 2.7}$ $15.8{\scriptstyle\pm 0.2}$ $3.2{\scriptstyle\pm 0.2}$ $1.4{\scriptstyle\pm 0.03}$ ${1.8\scriptstyle\pm 0.3}$ $1.9{\scriptstyle\pm 0.1}$
+EventFlow (25 NFEs, true $n$) $1.2{\scriptstyle\pm 0.01}$ $5.5{\scriptstyle\pm 0.3}$ $8.8{\scriptstyle\pm 0.2}$ $1.8{\scriptstyle\pm 0.02}$ $0.7{\scriptstyle\pm 0.01}$ $0.7{\scriptstyle\pm 0.02}$ $1.1{\scriptstyle\pm 0.02}$

### v3

@@
 The model builds on the flow matching framework in order to directly learn joint distributions over event times, side-stepping the autoregressive process.
-EventFlow is simple to implement and achieves a 20%-53% lower error than the nearest baseline on standard TPP benchmarks while simultaneously using fewer model calls at sampling time.
+EventFlow is simple to implement and achieves a 20%-53% lower forecast error than the nearest baseline on standard TPP benchmarks while simultaneously using fewer model calls at sampling time.
 1 Introduction
-Many stochastic processes, ranging from the occurrence of earthquakes (Ogata, 1998) to consumer behavior (Xu et al., 2014), are best understood as a sequence of discrete events which occur at random times.
-Any observed event sequence, consisting of one or more event times, may be viewed as a draw from a temporal point process (TPP) (Daley & Vere-Jones, 2003) which characterizes the distribution over such sequences.
+Many stochastic processes, ranging from the occurrence of earthquakes (Ogata, 1998) to consumer behavior (Xu et al., 2014), are best understood as a sequence of discrete events that occur at random times.
+Any observed event sequence, consisting of one or more event times, may be viewed as a draw from a temporal point process (TPP) (Daley and Vere-Jones, 2003) which characterizes the distribution over such sequences.
 Given a collection of observed event sequences, faithfully modeling the underlying TPP is critical in both understanding and forecasting the phenomenon of interest.
-While multiple different parametric TPP models have been proposed (Hawkes, 1971; Isham & Westcott, 1979), their limited flexibility limits their application when modeling complex real-world sequences.
-This has motivated the use of neural networks (Du et al., 2016; Mei & Eisner, 2017) in modeling TPPs.
-To date, most neural network based TPP models are autoregressive in nature (Shchur et al., 2020a; Zhang et al., 2020), where a model is trained to predict the next event time given an observed history of events.
+While multiple different parametric TPP models have been proposed (Hawkes, 1971; Isham and Westcott, 1979), their limited flexibility limits their application when modeling complex real-world sequences.
+This has motivated the use of neural networks (Du et al., 2016; Mei and Eisner, 2017) in modeling TPPs.
+To date, most neural TPP models are autoregressive in nature (Shchur et al., 2020a; Zhang et al., 2020; Shchur et al., 2021), where a model is trained to predict only a single subsequent event time given an observed history of events.
 However, in many tasks, we are interested not only in the next event, but in the entire sequence of events which is to follow.
-Despite their popularity, their performance in many-step forecasting tasks can be unsatisfactory due to compounding errors arising from the autoregressive sampling procedure (Xue et al., 2022; Lüdke et al., 2023).
+While autoregressive neural TPP models can be applied in this setting, their performance in many-step forecasting tasks can be unsatisfactory due to compounding errors arising from the autoregressive sampling procedure (Xue et al., 2022; Lüdke et al., 2023).
 Moreover, existing models are typically trained via a maximum likelihood procedure (see Section 3) which involves computing the CDF implied by the learned model.
-When using a neural model, computing this CDF necessitates techniques such as Monte Carlo estimation to properly compute the loss (Mei & Eisner, 2017).
-In addition, sampling from intensity-based models (Du et al., 2016; Mei & Eisner, 2017; Yang et al., 2022) is nontrivial, requiring an expensive and difficult to implement approach based on the thinning algorithm (Lewis & Shedler, 1979; Ogata, 1981; Xue et al., 2024).
-Motivated by these limitations, we propose EventFlow, a generative model which directly learns the joint event time distributions, thus allowing us to avoid autoregressive sampling altogether.
-Our proposed model extends the flow matching framework (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) to the setting of TPPs, where we learn a continuous flow from a reference TPP to our data TPP.
-At an intuitive level, samples from our model are generated by drawing a collection of event times from a reference distribution and flowing these events along a learned vector field.
-The number of events is fixed throughout this process, decoupling the event counts and their times, so that the distribution over event counts can be learned or otherwise specified.
-See Figure 1 for an illustration.
-More specifically,
-  • We propose EventFlow, a novel generative model for temporal point processes.
-Our model is suitable for both unconditional generation tasks (i.e., generating draws from the underlying data TPP) and conditional generation tasks (e.g., forecasting future events given a history), and is able to forecast multiple events simultaneously.
-  • We provide new perspective on modeling TPPs which sidesteps common pitfalls.
-In particular, the key idea of EventFlow is to decompose the generative process into a learned event count distribution and a flow-based generative model for the joint distribution of event times.
-Our model is non-autoregressive, easy to sample from, and straightforward to implement.
-  • On standard benchmarks, we demonstrate that EventFlow obtains state-of-the-art performance on a multi-step forecasting task and matches or exceeds the performance of state-of-the-art models for unconditional generation.
-In particular, we achieve a 20%-53% reduction in forecasting error from the nearest baseline.
-Compared with existing methods, which can require hundreds of model evaluations at sampling time, we demonstrate that EventFlow can achieve competitive performance with only a single forward pass.
+When using a neural model, computing this CDF often requires techniques such as Monte Carlo estimation to compute the loss (Mei and Eisner, 2017).
+In addition, sampling from intensity-based models (Du et al., 2016; Mei and Eisner, 2017; Yang et al., 2022) is nontrivial, requiring an expensive and difficult to implement approach based on the thinning algorithm (Lewis and Shedler, 1979; Ogata, 1981; Xue et al., 2024).
 Figure 1: All illustration of forecasting with our EventFlow method.
@@
 We first encode the observed history $\mathcal{H}$ into an embedding $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$.
-At $s=0$, we independently draw $n$ events in the forecasting window $[T_{0},T_{0}+\Delta T]$ from a fixed reference distribution, constituting a sample $\gamma_{0}$ from a mixed-binomial TPP.
+At $s=0$, we independently draw $n$ events in the forecasting window $T_{0},T_{0}+\Delta T]$ from a fixed reference distribution, constituting a sample $\gamma_{0}$ from a mixed-binomial TPP.
 Each event can be thought of as a particle, which is assigned a velocity by a neural network $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$.
-Each particle flows along its corresponding velocity field until reaching its terminal point at $s=1$, whereby we obtain a forecasted sequence $\gamma_{1}$.
+Each particle flows along its velocity field until reaching its terminal point at $s=1$, whereby we obtain a forecasted sequence $\gamma_{1}$.
+Motivated by these limitations, we propose EventFlow, a generative model which directly learns the full joint distribution over future event times, allowing us to avoid autoregressive sampling altogether.
+Our proposed model extends the flow matching framework (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023) to the setting of TPPs.
+To generate a sample, we first draw a set of random event times from a simple reference distribution and then transport them through a learned vector field to obtain a sequence of predicted times.
+The number of events is held fixed during this transformation, decoupling the modeling of event counts from that of event times.
+See Figure 1 for an illustration.
+This yields a flexible formulation that can be used for both unconditional generation (sampling from the underlying TPP) and conditional generation (forecasting future events given a history).
+More broadly, EventFlow provides an alternative perspective on TPP modeling: rather than specifying an intensity function and relying on sequential simulation, we decompose the generative process into a distribution over event counts and a flow-based model over event times.
+This results in a non-autoregressive model that is straightforward to train and efficient to sample from.
+Empirically, we demonstrate that EventFlow obtains state-of-the-art performance on multi-step forecasting benchmarks and is competitive with leading approaches for unconditional generation.
+In particular, our model reduces forecast error by 20%–53% relative to the strongest baselines.
+Further, compared with existing methods, which can require hundreds of model evaluations at sampling time, we demonstrate that EventFlow can achieve competitive performance with only a single forward pass of the underlying neural network.
 2 Related Work
 Temporal Point Processes
-The statistical modeling of temporal point processes (TPPs) is a classical subject with a long history (Daley & Vere-Jones, 2003; Hawkes, 1971; Isham & Westcott, 1979).
-The contemporary modeling paradigm, based on neural networks (Du et al., 2016), typically operates by learning a history encoder and an event decoder.
-The history encoder seeks to learn a fixed-dimensional vector representation of the history of a sequence up to some given time, and the decoder seeks to model a distribution over the subsequent event(s).
-Numerous models have been proposed for both components.
+The statistical modeling of temporal point processes (TPPs) has a long history (Daley and Vere-Jones, 2003; Hawkes, 1971; Isham and Westcott, 1979).
+The contemporary modeling paradigm, based on neural networks (Du et al., 2016; Shchur et al., 2021), typically operates by learning a history encoder and an event decoder.
+The history encoder seeks to learn a fixed-dimensional vector representation of the sequence of events which have been observed or already predicted, and the decoder seeks to model a distribution over the subsequent event time(s).
 Popular choices for the history encoder include RNN-based models (Du et al., 2016; Shchur et al., 2020a; Mei et al., 2019) or attention-based models (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022).
 While attention-based encoders can provide longer-range contexts, this benefit typically comes at the cost of additional memory overhead.
-Similarly, a wide range of forms for the event decoder have also been proposed.
-The most common approach is to parametrize a conditional intensity function via a neural network (Mei & Eisner, 2017; Zuo et al., 2020; Du et al., 2016).
-An alternative approach is to use generative models as decoders.
+For the event decoder, a common approach is to parametrize a conditional intensity function using a neural network that takes as input a learned representation of the event history (Du et al., 2016; Mei and Eisner, 2017; Zuo et al., 2020).
+An alternative and more recent approach is to use generative models as decoders.
 These models often do not assume a parametric form for the decoder, enhancing their flexibility.
@@
 Similarly, Shchur et al.
-(2020a) learn the distribution over the next inter-arrival time via a normalizing flow.
+(2020a) learn a distribution over the next event time via a normalizing flow.
 Lin et al.
 (2022) benchmark several choices of generative models, including diffusion, GANs, and VAEs.
-Despite the flexibility of these models, these approaches are all autoregressive in nature, making them ill-suited for multi-step forecasting tasks.
-In contrast, Lüdke et al.
-(2023) propose a diffusion-style model which is able to avoid autoregressive sampling via an iterative refinement procedure.
-Our work can be viewed as a novel approach to building flexible decoders for TPPs, extending flow matching to the setting of continuous-time event sequences.
-In contrast previous generative approaches, our model is trained in a likelihood-free manner and is non-autoregressive.
-The work of Lüdke et al.
-(2023) is most closely related to ours, but we emphasize that Lüdke et al.
-(2023) require an involved training and sampling procedure.
-In contrast, our method is straightforward to both implement and sample from, while simultaneously outperforming existing approaches.
+Regardless of the specific choice of decoder, these approaches are all autoregressive, making them ill-suited for multi-step forecasting tasks.
+Notable exceptions are the works of Lüdke et al.
+(2023, 2026), which propose flow and diffusion based models that avoids autoregressive sampling through an iterative refinement procedure.
+Our EventFlow model can be viewed as a flexible, non-autoregressive decoder for TPPs, obtained by extending flow matching to continuous-time event sequences.
+Among existing approaches, Lüdke et al.
+(2023, 2026) are most closely related in spirit.
+However, their method relies on a relatively involved training and sampling pipeline based on iterative refinement.
+In contrast, EventFlow admits a particularly simple regression-based formulation which is straightforward to implement.
+While previous works have studied marked TPPs (Du et al., 2016; Mei and Eisner, 2017; Draxler et al., 2025; Chang et al., 2025), EventFlow is focused on modeling the event times themselves.
 Flow Matching
-The recently introduced flow matching framework (or stochastic interpolants) (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) describes a class of generative models which are closely related to both normalizing flows (Papamakarios et al., 2021) and diffusion models (Ho et al., 2020; Song et al., 2021).
-Intuitively, these models learn a path of probability distributions which interpolates between a fixed reference distribution and the data distribution.
-These models are a popular alternative to diffusion, providing greater flexibility in model design, with recent applications in image generation (Ma et al., 2024; Dao et al., 2023), DNA and protein design (Stark et al., 2024; Campbell et al., 2024), and point cloud generation (Buhmann et al., 2023; Wu et al., 2023).
-To the best of our knowledge, our work is the first to explore flow matching for TPPs.
+The flow matching framework (or stochastic interpolants) (Lipman et al., 2023, 2024; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023) describes a class of generative models which are closely related to normalizing flows (Papamakarios et al., 2021).
+Flow matching has been explored for applications including image generation (Dao et al., 2023; Ma et al., 2024), DNA and protein design (Stark et al., 2024; Campbell et al., 2024), and 3D modeling (Buhmann et al., 2023; Wu et al., 2023; Xiang et al., 2025), but our work is the first to apply flow matching for TPPs.
 3 Autoregressive TPP Models
-We first provide a brief review of autoregressive point process models and discuss their shortcomings.
-Informally, one may think of an event sequence as a set $\{t_{k}\}_{k=1}^{n}$ of increasing event times.
-We will use $\mathcal{H}_{t}$ to represent the history of a sample up to (and including) time $t$, i.e., $\mathcal{H}_{t}=\{t_{k}:t_{k}\leq t\}$.
-Similarly, we use $\mathcal{H}_{t^{-}}=\{t_{k}:t_{k}<t\}$ to represent the history of a sample prior to time $t$.
+We first provide a brief, informal review of autoregressive point process models and discuss their shortcomings.
+Informally, an event sequence is a set $\{t^{k}\}_{k=1}^{n}$ of increasing event times.
+We will use $\mathcal{H}_{t}$ to represent the history of a sample up to (and including) time $t$, i.e., $\mathcal{H}_{t}=\{t^{k}:t^{k}\leq t\}$.
+Similarly, $\mathcal{H}_{t^{-}}=\{t^{k}:t^{k}<t\}$ represents the history prior to time $t$.
 In the autoregressive setting, the time of a single future event $t$ is modeled conditioned on the observed history of a sequence.
-This is typically achieved by either directly modeling a distribution over $t$ (Shchur et al., 2020a), or equivalently by modeling a conditional intensity function (Du et al., 2016).
-In the first approach, a conditional probability density of the form $p_{n}(t\mid\mathcal{H}_{t_{n}})$ is learned, allowing us to specify a joint distribution over event times $p(t_{1},\dots,t_{n})$ autoregressively via ${p_{n}(t_{1},\dots,t_{n})=p(t_{1})\prod_{k=2}^{n}p(t_{k}\mid\mathcal{H}_{t_{k-1}})}$.
-Alternatively, we may define the conditional intensity $\lambda^{*}(t):=\lambda(t\mid\mathcal{H}_{t^{-}})=p(t\mid\mathcal{H}_{t_{n}})/(1-F(t\mid\mathcal{H}_{t_{n}}))$, where $F(t\mid\mathcal{H}_{t_{n}})=\int_{t_{n}}^{t}p(s\mid\mathcal{H}_{t_{n}})\,\mathrm{d}s$ is the CDF associated with $p(t\mid\mathcal{H}_{t_{n}})$.
-Informally, the conditional intensity $\lambda^{*}(t)$ can be thought of (Rasmussen, 2011) as the instantaneous rate of occurrence of events at time $t$ given the previous $n$ events and given that no events have occurred since $t_{n}$.
-By integrating $\lambda^{*}(t)$, one can show that
-$F(t\mid\mathcal{H}_{t_{n}})=1-\exp\left(-\int_{t_{n}}^{t}\lambda^{*}(s)\,\mathrm{d}s\right)\qquad p(t\mid\mathcal{H}_{t_{n}})=\lambda^{*}(t)\exp\left(-\int_{t_{n}}^{t}\lambda^{*}(s)\,\mathrm{d}s\right)$ (1)
+This is often achieved by either directly modeling a distribution over $t$ (Shchur et al., 2020a), or equivalently by modeling a conditional intensity function (Du et al., 2016).
+More precisely, one models conditional densities of the form $p(t^{k}\mid\mathcal{H}_{t^{k-1}})$, which describe the distribution of the next event time given the history up to the previous event.
+This induces a joint distribution over event times $p(t^{1},\dots,t^{n})$ autoregressively via $p(t^{1},\dots,t^{n})=p(t^{1})\prod_{k=2}^{n}p(t^{k}\mid\mathcal{H}_{t^{k-1}})$}.
+Alternatively, we may define the conditional intensity $\lambda^{*}(t):=\lambda(t\mid\mathcal{H}_{t^{-}})=p(t\mid\mathcal{H}_{t^{n}})/(1-F(t\mid\mathcal{H}_{t^{n}}))$, where $F(t\mid\mathcal{H}_{t^{n}})=\int_{t^{n}}^{t}p(s\mid\mathcal{H}_{t^{n}})\,\mathrm{d}s$ is the CDF associated with $p(t\mid\mathcal{H}_{t^{n}})$.
+Informally, the conditional intensity $\lambda^{*}(t)$ can be thought of (Rasmussen, 2011) as the instantaneous rate of occurrence of events at time $t$ given the history up to $t^{n}$ and that no events have occurred in $(t^{n},t)$.
+By integrating $\lambda^{*}(t)$, one can verify
+$F(t\mid\mathcal{H}_{t^{n}})=1-\exp\left(-\int_{t_{n}}^{t}\lambda^{*}(s)\,\mathrm{d}s\right)$ (1)
+$p(t\mid\mathcal{H}_{t^{n}})=\lambda^{*}(t)\exp\left(-\int_{t_{n}}^{t}\lambda^{*}(s)\,\mathrm{d}s\right)$ (2)
 and thus one may recover the conditional distribution from the conditional intensity under mild additional assumptions (Rasmussen, 2011, Prop 2.2).
 The Likelihood Function
-Suppose we observe an event sequence $\{t_{k}\}_{k=1}^{n}$ on the interval $[0,T]$.
-The likelihood of this sequence can be seen loosely as the probability of seeing precisely $n$ events at these times.
+Suppose we observe an event sequence $\{t^{k}\}_{k=1}^{n}$ on the interval $[0,T]$.
+The likelihood of this sequence can be seen loosely as the probability of observing events at these times and no others within the observation window.
 The likelihood may be expressed in terms of either the density or intensity via
-$L\left(\{t_{k}\}_{k=1}^{n}\right)=p_{n}(t_{1},\dots,t_{n})\left(1-F(T\mid\mathcal{H}_{t_{n}})\right)=\left(\prod_{k=1}^{n}\lambda^{*}(t_{k})\right)\exp\left(-\Lambda^{*}(T)\right)$ (2)
-where the CDF term is included to indicate that no events beyond $t_{n}$ have occurred and $\Lambda^{*}(T)=\int_{0}^{T}\lambda^{*}(s)\,\mathrm{d}s$ is the total intensity.
-Autoregressive models are typically trained by maximizing this likelihood (Du et al., 2016; Mei & Eisner, 2017; Shchur et al., 2020a).
-We emphasize that this likelihood is not simply the joint event-time density $p_{n}(t_{1},\dots,t_{n})$, as the likelihood also measures the fact that no events occur between $t_{n}$ and $T$.
+$L\left(\{t^{k}\}_{k=1}^{n}\right)$ $=p(t^{1},\dots,t^{n})\left(1-F(T\mid\mathcal{H}_{t^{n}})\right)$ (3)
+$=\left(\prod_{k=1}^{n}\lambda^{*}(t^{k})\right)\exp\left(-\Lambda^{*}(T)\right)$ (4)
+where $\Lambda^{*}(T)=\int_{0}^{T}\lambda^{*}(s)\,\mathrm{d}s$ is the total intensity.
+The survival term $1-F(T\mid\mathcal{H}_{t^{n}})$ accounts for the absence of events in $(t^{n},T]$.
+Autoregressive models are typically trained by maximizing this likelihood (Du et al., 2016; Mei and Eisner, 2017; Shchur et al., 2020a).
+Critically, we emphasize that $p(t^{1},\dots,t^{n})$ captures a joint density over $n$ event times, but does not account for what occurs after the final event $t^{n}$.
+In contrast, the likelihood $L\left(\{t^{k}\}_{k=1}^{n}\right)$ corresponds to observations in a finite window $[0,T]$, and therefore includes an additional CDF term to account for the fact that no events were observed in $(t^{n},T]$.
 It is worth noting that evaluating $L(\{t_{k}\})$ can be non-trivial in practice.
-For models that parametrize $\lambda^{*}(t)$ via a neural network (Du et al., 2016; Mei & Eisner, 2017), computing the total intensity $\Lambda^{*}(T)$ is often done via a Monte Carlo integral, requiring many forward passes of the model to evaluate $\lambda^{*}(t)$ at different values of $t$.
+For models that parametrize $\lambda^{*}(t)$ via a neural network (Du et al., 2016; Mei and Eisner, 2017), computing the total intensity $\Lambda^{*}(T)$ is often done via a Monte Carlo integral, requiring many forward passes of the model to evaluate $\lambda^{*}(t)$ at different values of $t$.
 Models which directly parametrize the density $p(t\mid\mathcal{H}_{t})$ suffer from the same drawback when computing the corresponding CDF in Equation (3).
@@
 Moreover, the predictive performance of autoregressive models is often unsatisfactory on multi-step generation tasks due to the accumulation of errors over many steps (Lin et al., 2021; Lüdke et al., 2023).
-This difficulty is particularly pronounced for intensity-based models (Du et al., 2016; Mei & Eisner, 2017; Zhang et al., 2020), where naively computing the implied distribution in Equation (1) is prohibitively expensive.
-Instead, sampling from intensity-based models is typically achieved via the thinning algorithm (Ogata, 1981; Lewis & Shedler, 1979).
+This difficulty is particularly pronounced for intensity-based models (Du et al., 2016; Mei and Eisner, 2017; Zhang et al., 2020), where naively computing the implied distribution in Equation (2) is prohibitively expensive.
+Instead, sampling from intensity-based models is typically achieved via the thinning algorithm (Ogata, 1981; Lewis and Shedler, 1979).
 However, this algorithm has several hyperparameters to tune, is challenging to parallelize, and can be difficult for practitioners to implement (Xue et al., 2024).
@@
 4 EventFlow
-Motivated by the limitations of autoregressive models, we propose EventFlow, which has a number of distinct advantages.
+Motivated by the limitations of autoregressive models, we propose EventFlow, which has a number of distinct advantages over prior work.
 First, EventFlow directly models the joint distribution over event times, thereby avoiding autoregression entirely.
@@
 In fact, EventFlow can achieve competitive performance using only a single forward pass at sampling time.
-We build upon the flow matching (or stochastic interpolant) framework (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) to develop our model.
-We begin below by focusing on the unconditional setting, followed by an extension for conditional generation.
+We build upon the flow matching (or stochastic interpolant) framework (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023) to develop our model.
+We begin below by focusing on the unconditional setting (i.e., generating an event sequence without being given an observed history), followed by an extension for conditional generation.
 Preliminaries
@@
 A temporal point process (TPP) on $\mathcal{T}$ is a probability distribution $\mu$ over the configuration space $\Gamma$.
-Informally, $\mu$ represents a distribution over finite sequences $\gamma\in\Gamma$ whose events are in $\mathcal{T}$.
-We use $N:\Gamma\to\mathbb{Z}_{\geq 0}$ to denote the counting functional, i.e., $N(\gamma)$ is the number of events in the sequence $\gamma$ [^1] This can be thought of in terms of the counting process, i.e., $N(\gamma)$ corresponds to the value of the associated counting process at the ending time $T$, or the total number of events in $\gamma$ that have occurred in the interval $[0,T]$.
+We use $\mathbb{P}(\Gamma)$ to denote the set of all such distributions.
+Informally, a TPP $\mu$ is a distribution over finite sequences $\gamma\in\Gamma$ whose events are in $\mathcal{T}$.
+We use $N:\Gamma\to\mathbb{Z}_{\geq 0}$ to denote the counting functional, i.e., $N(\gamma)$ is the number of events in the sequence $\gamma$.
+[^1] This can be thought of in terms of the counting process, i.e., $N(\gamma)$ corresponds to the value of the associated counting process at the ending time $T$, or the total number of events in $\gamma$ that have occurred in the interval $[0,T]$.
 While it is common to represent TPPs as distributions over random sets of event times, in our approach it will be more convenient to represent TPPs as random measures (Kallenberg, 2017).
-Under mild assumptions, a TPP $\mu$ can be fully characterized (Daley & Vere-Jones, 2003, Prop.
+Under mild assumptions, a TPP $\mu$ can be fully characterized (Daley and Vere-Jones, 2003, Prop.
 5.3.II) by a probability distribution which specifies the number of events and a collection of joint densities corresponding to the event times themselves.
@@
 We further restrict each $\mu^{n}$ to be supported only on the ordered sets, so that $t^{1}<t^{2}<\dots<t^{n}$.
-Let $\mu_{1}$ represent the data distribution and $\mu_{0}$ represent a reference distribution.
-That is, both $\mu_{0},\mu_{1}\in\mathbb{P}(\Gamma)$ are TPP distributions.
-To construct our model, we will define a path of TPPs $\eta_{s}\in\mathbb{P}(\Gamma)$ which approximately interpolates from our reference distribution to our data distribution.
+Let $\mu_{1}$ represent the data distribution and $\mu_{0}$ represent a reference distribution, i.e., both $\mu_{0},\mu_{1}\in\mathbb{P}(\Gamma)$ are distributions over event sequences.
+To construct our model, we will define a path of distributions $(\eta_{s})_{s\in[0.1]}\subset\mathbb{P}(\Gamma)$ which approximately interpolates from our reference TPP to our data TPP.
+Intuitively, this path provides us with a way to transform samples: we can draw a sequence of events from $\mu_{0}$ and gradually move them along this path to obtain a sequence approximately distributed according to $\mu_{1}$.
 Throughout, we use $s\in[0,1]$ to denote a flow time and $t\in[0,T]$ to denote an event time.
@@
 Balanced Couplings
-Our first step is to define a useful notion of couplings, allowing us to pair event sequences drawn from $\mu_{0}$ with those drawn from $\mu_{1}$.
-A coupling between two TPPs $\mu,\nu\in\mathbb{P}(\Gamma)$ is a joint probability measure $\rho\in\mathbb{P}(\Gamma\times\Gamma)$ over pairs of event sequences $(\gamma_{0},\gamma_{1})$ such that the marginal distributions of $\rho$ are $\mu$ and $\nu$.
+Towards constructing such a path, our first step is to define a notion of coupling, which allows us to pair event sequences drawn from the reference distribution $\mu_{0}$ with those drawn from the data distribution $\mu_{1}$.
+Formally, a coupling between two TPPs $\mu_{0},\mu_{1}\in\mathbb{P}(\Gamma)$ is a joint probability distribution $\rho\in\mathbb{P}(\Gamma\times\Gamma)$ over pairs of event sequences $(\gamma_{0},\gamma_{1})$ such that the marginal distributions of $\rho$ are $\mu_{0}$ and $\mu_{1}$.
+Sampling $(\gamma_{0},\gamma_{1})\sim\rho$ from a coupling gives us paired sequences of events.
+While couplings are broadly used in transport-based generative modeling (Tong et al., 2024; Villani et al., 2009), we introduce a new class of couplings which is particularly well-suited to the TPP setting.
 We say that the coupling $\rho$ is balanced if draws $(\gamma_{0},\gamma_{1})\sim\rho$ are such that $N(\gamma_{0})=N(\gamma_{1})$ almost surely.
-In other words, balanced couplings only pair event sequences with equal numbers of events.
-While we will later see how to interpolate between any two given event sequences, this coupling will allow us to decide which sequences to interpolate.
-In particular, a balanced coupling will allow us to pair sequences such that they always have the same number of events, allowing us to avoid the addition or deletion of events during both training and sampling.
-We will use $\Pi_{b}(\mu,\nu)$ to denote the set of balanced couplings.
-The following proposition shows $\Pi_{b}(\mu,\nu)$ is nonempty if and only if the event count distributions of $\mu$ and $\nu$ are equal, placing a structural constraint on the suitable choices of $\mu_{0}$.
+That is, balanced couplings only pair event sequences with equal numbers of events.
+This ensures that when we later construct interpolations between paired sequences, we do not need to add or remove events, simplifying both training and sampling.
+We will use $\Pi_{b}(\mu_{0},\mu_{1})$ to denote the set of balanced couplings.
+The following proposition characterizes when balanced couplings exist.
+In particular, $\Pi_{b}(\mu_{0},\mu_{1})$ is nonempty if and only if the event count distributions of $\mu_{0}$ and $\mu_{1}$ are identical, imposing a structural constraint on suitable choices of $\mu_{0}$: the reference TPP $\mu_{0}$ must have the same event count distribution as the data TPP $\mu_{1}$.
 Proposition 1 (Existence of Balanced Couplings).
-Let $\mu,\nu\in\mathbb{P}(\Gamma)$ be two TPPs. The set of balanced couplings $\Pi_{b}(\mu,\nu)$ is nonempty if and only if $\mu(n)=\nu(n)$ have the same distribution over event counts.
+Let $\mu_{0},\mu_{1}\in\mathbb{P}(\Gamma)$ be two TPPs.
+The set of balanced couplings $\Pi_{b}(\mu_{0},\mu_{1})$ is nonempty if and only if have the same distribution over event counts, i.e., $\mu_{0}(n)=\mu_{1}(n)$ for every $n\in\mathbb{Z}_{\geq 0}$.
 We provide a proof in Appendix B.
 In practice, we follow a simple strategy for choosing both the reference TPP $\mu_{0}$ and the coupling $\rho$.
-Suppose $q$ is any given density on our state space $\mathcal{T}$, e.g., a uniform distribution.
+Let $q$ be a density on $\mathcal{T}$ (e.g., uniform).
 We take $\mu_{0}$ to be a mixed binomial process (Kallenberg, 2017, Ch.
-3) whose event count distribution is given by that of the data $\mu_{1}(n)$ and joint event distributions given by independent products of $q$ (up to sorting).
-That is, to sample from $\mu_{0}$, one can simply sample $n\sim\mu_{1}(n)$ from the empirical event count distribution followed by sampling and sorting $n$ i.i.d.
-points $t^{k}\sim q$.
-To draw a sample from our coupling $\rho$, we first sample a data sequence $\gamma_{1}\sim\mu_{1}$, followed by sampling $N(\gamma_{1})$ events independently from $q$ and sorting to produce a draw $\gamma_{0}\sim\mu_{0}$.
+3) whose event count distribution is given by that of the data $\mu_{1}(n)$, and joint event time distributions given by independent products of $q$ (up to sorting).
+Sampling $\gamma_{0}\sim\mu_{0}$ amounts to sampling $n\sim\mu_{1}(n)$ from the data event count distribution followed by sampling and sorting $n$ i.i.d.
+times $t^{k}\sim q$.
+Under this choice, a sample from a balanced coupling $\rho\in\Pi_{b}(\mu_{0},\mu_{1})$ can be produced by first sampling a data sequence $\gamma_{1}\sim\mu_{1}$, followed by sampling $N(\gamma_{1})$ events independently from $q$ and sorting to produce a paired draw $\gamma_{0}\sim\mu_{0}$.
 Interpolant Construction
-We now proceed to construct our interpolant $\eta_{s}\in\mathbb{P}(\Gamma)$.
-We will construct this path of TPPs via a local procedure which we then marginalize over a given balanced coupling.
-Here, we extend flow matching techniques (Lipman et al., 2023; Tong et al., 2024) to design our sequence-level interpolants, but we emphasize that this is only possible under a balanced coupling as the number of events is fixed.
-To that end, let $\rho$ be any balanced coupling of the reference measure $\mu_{0}$ and the data measure $\mu_{1}$, and suppose $z:=(\gamma_{0},\gamma_{1})\sim\rho$ is a draw from this coupling.
+We now construct our path $(\eta_{s})_{s\in[0,1]}\subset\mathbb{P}(\Gamma)$ using a two-stage procedure.
+First, for a given pair of sequences $(\gamma_{0},\gamma_{1})$ drawn from a balanced coupling, we define a sequence-level interpolation $(\gamma_{s})_{s\in[0,1]}$ that smoothly transforms $\gamma_{0}$ into $\gamma_{1}$.
+Second, the distribution $\eta_{s}$ is obtained by averaging these interpolated sequences over the coupling.
+We extend flow matching techniques (Lipman et al., 2023; Tong et al., 2024) to define the interpolants, and emphasize that fixing the number of events via a balanced coupling is essential for this construction.
+To that end, let $\rho$ be any balanced coupling of the reference distribution $\mu_{0}$ and the data distribution $\mu_{1}$, and suppose $z:=(\gamma_{0},\gamma_{1})\sim\rho$ is a draw from this coupling.
 As $\rho$ is balanced, we have $\gamma_{0}=\sum_{k=1}^{n}\delta[t_{0}^{k}]$ and $\gamma_{1}=\sum_{k=1}^{n}\delta[t_{1}^{k}]$ are both a collection of $n$ events.
-As TPPs are fully characterized by their joint distributions over event times, we will henceforth describe our procedure for a fixed (but arbitrary) number of events $n$.
-First, we the define measure $\gamma_{s}^{z}\in\Gamma$ via
+We will henceforth describe our procedure for a fixed (but arbitrary) number of events $n$, and we will later describe how to model the number of events itself.
+First, we define the interpolant sequence $\gamma_{s}^{z}\in\Gamma$ via
 $\gamma_{s}^{z}=\sum_{k=1}^{n}\delta[t_{s}^{k}]\qquad t_{s}^{k}=(1-s)t_{0}^{k}+st_{1}^{k}\qquad 0\leq s\leq 1$ (5)
 where we use the superscript $z$ to denote the dependence on the pair $z=(\gamma_{0},\gamma_{1})$.
-In other words, $\gamma_{s}^{z}$ linearly interpolates each corresponding event in $\gamma_{0}$ and $\gamma_{1}$.
-This defines a path $(\gamma_{s}^{z})_{s=0}^{1}$ in the configuration space $\Gamma$ which evolves the reference sample $\gamma_{0}$ into the data sample $\gamma_{1}$.
-We now lift this deterministic path $(\gamma_{s}^{z})_{s=0}^{1}\subset\Gamma$ to a path of TPP distributions $(\eta_{s}^{z})_{s=0}^{1}\subset\mathbb{P}(\Gamma)$.
-We define the point process distribution $\eta_{s}^{z}\in\mathbb{P}(\Gamma)$ implicitly by adding independent Gaussian noise to each of the events in $\gamma_{s}^{z}$.
+In other words, $\gamma_{s}^{z}$ linearly interpolates each corresponding event in $\gamma_{0}$ and $\gamma_{1}$, defining a path $(\gamma_{s}^{z})_{s\in[0,1]}\subset\Gamma$ which evolves the reference sample $\gamma_{0}$ into the data sample $\gamma_{1}$.
+We now lift this deterministic path $(\gamma_{s}^{z})_{s\in[0,1]}\subset\Gamma$ to a path of TPP distributions $(\eta_{s}^{z})_{s\in[0,1]}\subset\mathbb{P}(\Gamma)$.
+We define the point process distribution $\eta_{s}^{z}\in\mathbb{P}(\Gamma)$ implicitly by adding independent Gaussian noise to each event in $\gamma_{s}^{z}$.
 That is, a draw $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ may be simulated via
 $\hat{\gamma}_{s}^{z}=\sum_{k=1}^{n}\delta\left[t_{s}^{k}+\epsilon^{k}\right]\qquad\epsilon^{k}\sim\mathcal{N}(0,\sigma^{2}).$ (6)
-In principle this means that the support of $\eta_{s}^{z}$ is larger than $\mathcal{T}$, but in practice we choose $\sigma^{2}$ sufficiently small such that this is not a concern.
-The addition of noise $\epsilon_{k}$ is instrumental in obtaining a well-specified model, but in practice the noise variance $\sigma^{2}$ is not a critical hyperparameter.
+In principle, using Gaussian noise means that the support of event times in $\eta_{s}^{z}$ is larger than $\mathcal{T}$, but in practice we choose $\sigma^{2}$ sufficiently small such that this is not a concern.
+The addition of noise $\epsilon^{k}$ is instrumental in obtaining a well-specified model, but in practice we found the noise variance $\sigma^{2}$ to not be a critical hyperparameter.
 We note that this noising step is typical in flow matching models (Lipman et al., 2023; Tong et al., 2024).
-Finally, for $s\in[0,1]$, we define the marginal TPP measure $\eta_{s}\in\mathbb{P}(\Gamma)$ via $\eta_{s}=\int\eta_{s}^{z}\,\mathrm{d}\rho(z)$.
-Observe that, by construction, the event count distribution $\eta_{s}(n)$ is given by $\mu_{1}(n)$ for all $s\in[0,1]$.
+Finally, for $s\in[0,1]$, we define the marginal TPP measure $\eta_{s}\in\mathbb{P}(\Gamma)$ via the mixture distribution
+$\eta_{s}=\int\eta_{s}^{z}\,\mathrm{d}\rho(z)$ (7)
+which, intuitively, corresponds to the two-stage procedure of drawing a sample $z\sim\rho$ and subsequently following the conditional flow defined in Equation (6).
+By construction, the event count distribution $\eta_{s}(n)$ is given by $\mu_{1}(n)$ for all $s\in[0,1]$.
 This path of TPP distributions $\eta_{s}$ approximately interpolates from the reference TPP $\mu_{0}$ at $s=0$ to the data TPP $\mu_{1}$ at $s=1$, in the sense that at the endpoints, the joint event time distributions $\eta_{0}^{n}(t^{1},\dots,t^{n})$ and $\eta_{1}^{n}(t^{1},\dots,t^{n})$ are given by a convolution of $\mu_{0}^{n}(t^{1},\dots,t^{n})$ and $\mu_{1}^{n}(t^{1},\dots,t^{n})$ with the Gaussian $\mathcal{N}(0,\sigma^{2}I_{n})$.
-As $\sigma^{2}\downarrow 0$, it is clear that we recover a genuine interpolant (Tong et al., 2024).
-For single event $t_{s}^{k}$, we have constructed a path of distributions $\mathcal{N}(t_{s}^{k},\sigma^{2})$.
-This transport of a Gaussian can be achieved infinitesimally through the vector field $v_{s}^{k}(t)=t_{1}^{k}-t_{0}^{k}$ (Tong et al., 2024).
-Thus, the evolution (4) is generated by the vector field $v_{s}^{z}:\mathcal{T}^{n}\to\mathbb{R}^{n}$,
-$v_{s}^{z}(\gamma)=\begin{bmatrix}v_{s}^{1},\dots,v_{s}^{n}\end{bmatrix}^{\sf T}=\begin{bmatrix}t_{1}^{1}-t_{0}^{1},&\dots,&t_{1}^{n}-t_{0}^{n}\end{bmatrix}^{\sf T}\qquad 0\leq s\leq 1.$ (5)
-In other words, flowing samples $\hat{\gamma}_{0}^{z}$ along the vector field $v_{s}^{z}$ generates the path of distributions $\eta_{s}^{z}$.
-However, this path is conditioned on $z$, and we would like to find the vector field $v_{s}$ which generates the unconditional path $\eta_{s}$.
-As is standard in flow matching (Lipman et al., 2023; Tong et al., 2024; Albergo & Vanden-Eijnden, 2023), the unconditional vector field $v_{s}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ may be obtained via
-$v_{s}(\gamma)=\int v_{s}^{z}(\gamma)\frac{\,\mathrm{d}\eta_{s}^{z}}{\,\mathrm{d}\eta_{s}}(\gamma)\,\mathrm{d}\rho(z).$ (6)
-We have thus far described a procedure for interpolating between a given reference distribution $\mu_{0}^{n}$ and the data distribution $\mu_{1}^{n}$ for a given, fixed number of events $n$.
-As $n$ was arbitrary, we have successfully constructed a family of interpolants which will enable us to sample from the joint event distribution for any $n$.
-However, to fully characterize the TPP distribution, we need to also specify the event count distribution.
-For unconditional generation tasks, this is straightforward – we simply follow the empirical event count distribution see in the training data.
-We describe our approach for modeling the event count distribution in conditional tasks in the following section.
-1 Sample $\gamma_{1}\sim\mu_{1}$, $s\sim\mathcal{U}[0,1]$, $\epsilon\sim\mathcal{N}(0,1)$
-2 $e_{\mathcal{H}}=\varnothing$
-3  /* Null history */
-4 if forecast then
-5       Sample $T_{0}\in[\Delta T,T-\Delta T]$
-6       $\mathcal{H}\leftarrow\{t\in\gamma_{1}:t\leq T_{0}\}$
-7       $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H})$
-8       $\gamma_{1}\leftarrow\{t\in\gamma_{1}:T_{0}<t\leq T_{0}+\Delta T\}$
-9
-[…]
-11 Sample and sort $t_{0}^{k}\sim q$ for $k=1,\dots,n$
-12 $\gamma_{0}\leftarrow(t_{0}^{1},\dots,t_{0}^{n})$
-13 $t_{s}^{k}\leftarrow(1-s)t_{0}^{k}+st_{1}^{k}$ for $k=1,\dots,n$
-14 $\gamma_{s}^{z}\leftarrow(t_{s}^{1},\dots,t_{s}^{n})$
+As $\sigma^{2}\downarrow 0$, it is clear that we recover a genuine interpolant.
+Observe that in Equation (6), for each individual event $t_{s}^{k}$, we have constructed a path of Gaussian distributions $\mathcal{N}(t_{s}^{k},\sigma^{2})$ centered around said event.
+This path of Gaussians can be simulated via the vector field $v_{s}^{k}:=t_{1}^{k}-t_{0}^{k}$ (Tong et al., 2024).
+That is, if we draw an initial condition $\tau_{0}^{k}\sim\mathcal{N}(t_{0}^{k},\sigma^{2})$ and solve the differential equation $\mathrm{d}\tau_{s}^{k}=v_{s}^{k}\,\mathrm{d}s$ for $s\in[0,1]$ starting from $\tau_{0}^{k}$, we obtain a draw $\tau_{1}^{k}\sim\mathcal{N}(t_{1}^{k},\sigma^{2})$.
+Loosely, $v_{s}^{k}$ is a velocity associated with the event $t_{s}^{k}$.
+Since each event in Equation (6) evolves independently, the path of distributions ${\eta}_{s}^{z}$ is generated by the constant vector field $v_{s}^{z}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ defined as
+$v_{s}^{z}(\gamma):=\begin{bmatrix}v_{s}^{1},\dots,v_{s}^{n}\end{bmatrix}^{\sf T}=\begin{bmatrix}t_{1}^{1}-t_{0}^{1},&\dots,&t_{1}^{n}-t_{0}^{n}\end{bmatrix}^{\sf T}.$ (8)
+In other words, flowing initial samples $\hat{\gamma}_{0}^{z}\sim\eta_{0}^{z}$ along the vector field $v_{s}^{z}(\gamma)$ generates the path of distributions $\eta_{s}^{z}$, approximately interpolating from the noise sequence $\gamma_{0}$ to the data sequence $\gamma_{1}$.
+However, this path is conditioned on $z=(\gamma_{0},\gamma_{1})$.
+Since we seek to generate $\gamma_{1}$, this is intractable, and we would instead like to find the vector field $v_{s}$ that generates the unconditional path $\eta_{s}$ defined in Equation (6).
+To this end, we aggregate the conditional vector fields $v_{s}^{z}(\gamma)$ across the coupling $z\sim\rho$.
+In particular, the unconditional vector field $v_{s}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ is given by
+$v_{s}(\gamma)=\mathbb{E}_{z\sim\rho}[v_{s}^{z}(\gamma)\mid\gamma_{s}^{z}=\gamma]=\int v_{s}^{z}(\gamma)\frac{\,\mathrm{d}\eta_{s}^{z}}{\,\mathrm{d}\eta_{s}}(\gamma)\,\mathrm{d}\rho(z)$ (9)
+i.e., the average of the conditional vector fields over all pairs $z$ that could have produced the sequence $\gamma$ at time $s$.
+This corresponds to an extension of the standard flow matching construction (Lipman et al., 2023; Tong et al., 2024; Albergo and Vanden-Eijnden, 2023) to the setting of TPPs.
+Sample $\gamma_{1}\sim\mu_{1}$, $s\sim\mathcal{U}[0,1]$, $\epsilon\sim\mathcal{N}(0,1)$
+$e_{\mathcal{H}}=\varnothing$
+/* Null history */
+if forecast then
+   Sample $T_{0}\in[\Delta T,T-\Delta T]$
+ $\mathcal{H}\leftarrow\{t\in\gamma_{1}:t\leq T_{0}\}$
+ $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H})$
+ $\gamma_{1}\leftarrow\{t\in\gamma_{1}:T_{0}<t\leq T_{0}+\Delta T\}$
+end if
+$n\leftarrow N(\gamma_{1})$
+Sample and sort $t_{0}^{k}\sim q$ for $k=1,\dots,n$
+$\gamma_{0}\leftarrow(t_{0}^{1},\dots,t_{0}^{n})$
+$t_{s}^{k}\leftarrow(1-s)t_{0}^{k}+st_{1}^{k}$ for $k=1,\dots,n$
+$\gamma_{s}^{z}\leftarrow(t_{s}^{1},\dots,t_{s}^{n})$
 Take a gradient step on $\left\lVert\gamma_{1}-\gamma_{0}-v_{\theta}\left(\gamma_{s}^{z}+\epsilon,s,e_{\mathcal{H}}\right)\right\rVert^{2}$
 Algorithm 1 Training EventFlow
-1 Choose a flow time discretization $10=s_{0}<s_{1}<\dots<s_{K}=$ 1
-2 $e_{\mathcal{H}}=\varnothing$
-3  /* Null history */
-4 if forecast then
-5       $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H}$)
-6       $n\sim p_{\phi}(n\mid\mathcal{H})$
-7
-8else
-9       $n\sim\mu_{1}(n)$
-10
-11Sample and sort $t_{0}^{k}\sim q$ for $k=1,\dots,n$
-12 $\gamma_{0}\leftarrow(t_{0}^{1},\dots,t_{0}^{n})$
-13 for $k=1,2,\dots,K$ do
-14       $h_{k}\leftarrow s_{k}-s_{k-1}$
-15       $\gamma_{s_{k}}\leftarrow\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1},e_{\mathcal{H}})$
-16
+Training
+If we knew $v_{s}$ in Equation (9), we could draw a sample from $\gamma_{1}\sim\mu_{1}$ by drawing a sample event sequence $\gamma_{0}\sim\mu_{0}$ from the reference TPP and flowing each event along $v_{s}$.
+More precisely, $\gamma_{0}$ will consist of $N(\gamma_{0})$ events and $v_{s}(\gamma)$ will be a vector field with $N(\gamma_{0})$ components, so that we may solve the differential equation $\mathrm{d}\gamma_{s}=v_{s}(\gamma)\,\mathrm{d}s$ on $s\in[0,1]$.
+Although the marginal vector field in Equation (9) admits an analytical form, it is intractable to compute in practice as the density ratio $\mathrm{d}\eta_{s}^{z}/\,\mathrm{d}\eta_{s}$ is unknown.
+To overcome this, we instead regress on the conditional vector fields $v_{s}^{z}$.
+Here, $v_{\theta}(\gamma_{s},s)$ will represent a neural network with parameters $\theta$ which takes in a sequence $\gamma_{s}$ along with the flow time $s$.
+We aim to minimize
+$J(\theta)=\mathbb{E}_{s,(\gamma_{0},\gamma_{1}),\hat{\gamma}_{s}^{z}}\left[\left\lVert\gamma_{1}-\gamma_{0}-v_{\theta}(\hat{\gamma}_{s}^{z},s)\right\rVert^{2}\right]$ (10)
+which is equal to the MSE loss on the unconditional $v_{s}$ up to an additive constant not depending on $\theta$ (Lipman et al., 2023; Tong et al., 2024).
+We estimate $J(\theta)$ by uniformly sampling a flow time $s\in[0,1]$, a pair $z=(\gamma_{0},\gamma_{1})\sim\rho$ from our balanced coupling and drawing a noisy interpolant $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ according to Equation (6).
+To train the model for forecasting, where the goal is to predict a future sequence of events conditioned on a history $\mathcal{H}$, we embed $\mathcal{H}$ into a fixed-dimensional vector representation $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$ via a learned encoder $f_{\theta}$ before providing this to the model $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$ and minimizing Equation (10).
+Note that we jointly train the encoder $f_{\theta}$ and vector field $v_{\theta}$.
+See Algorithm 1 for training pseudocode.
+Event Count Distributions
+We have thus far described a procedure for interpolating between a given reference distribution $\mu_{0}^{n}$ and the data distribution $\mu_{1}^{n}$ for a given, fixed number of events $n$.
+As $n$ was arbitrary, we have successfully constructed a family of interpolants which will enable us to sample from the joint event time distributions $\mu_{1}^{n}(t^{1},\dots,t^{n})$.
+However, recall that fully characterizing the TPP distribution requires us to also specify the event count distribution.
+For unconditional generation tasks, this is straightforward: we simply follow the empirical event count distribution see in the training data.
+When forecasting, though, we must also learn a model of the event count distribution $p_{\phi}(n\mid\mathcal{H})$, i.e., how many events are likely to occur after observing a history $\mathcal{H}$.
+Here, $\phi$ represents the parameters of this model.
+We train $p_{\phi}(n\mid\mathcal{H})$ by minimizing a cross-entropy loss.
+In practice, we found it important to regularize this loss to encourage smoothness in $n$, which we achieve through an optimal-transport based regularizer.
+See Appendix C for details.
+As both $v_{\theta}$ and $p_{\phi}$ must be able to operate on variable-length sequences, we use a transformer-based backbone for both models. However, our overall method is agnostic to this choice.
+Choose a flow time discretization $0=s_{0}<s_{1}<\dots<s_{K}=1$
+$e_{\mathcal{H}}=\varnothing$
+/* Null history */
+if forecast then
+ $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H})$
+ $n\sim p_{\phi}(n\mid\mathcal{H})$
+else
+ $n\sim\mu_{1}(n)$
+end if
+Sample and sort $t_{0}^{k}\sim q$ for $k=1,\dots,n$
+$\gamma_{0}\leftarrow(t_{0}^{1},\dots,t_{0}^{n})$
+for $k=1,2,\dots,K$ do
+ $h_{k}\leftarrow s_{k}-s_{k-1}$
+ $\gamma_{s_{k}}\leftarrow\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1},e_{\mathcal{H}})$
+end for
 return $\gamma_{0}$
 Algorithm 2 Sampling EventFlow
-Training
-If we knew $v_{s}$ in Equation (6), we could draw samples from $\mu_{1}$ by drawing a sample event sequence $\gamma_{0}\sim\mu_{0}$ from the reference TPP and flowing each event along $v_{s}$.
-Although the marginal vector field in Equation (6) admits an analytical form, it is intractable to compute in practice as the marginal measure $\eta_{s}$ is not available.
-To overcome this, we instead regress on the conditional vector fields $v_{s}^{z}$.
-Here, $v_{\theta}(\gamma_{s},s)$ will represent a neural network with parameters $\theta$ which takes in a sequence $\gamma_{s}$ along with the flow time $s$.
-We seek to minimize the loss
-$J(\theta)=\mathbb{E}_{s,(\gamma_{0},\gamma_{1}),\hat{\gamma}_{s}^{z}}\left[\left\lVert\gamma_{1}-\gamma_{0}-v_{\theta}(\hat{\gamma}_{s}^{z},s)\right\rVert^{2}\right]$ (7)
-which previous works on flow matching have shown to be equal to MSE regression on the unconditional $v_{s}$ up to an additive constant not depending on $\theta$ (Lipman et al., 2023; Tong et al., 2024).
-We estimate $J(\theta)$ by uniformly sampling a flow time $s\in[0,1]$, a pair $z=(\gamma_{0},\gamma_{1})\sim\rho$ from our balanced coupling and drawing a noisy interpolant $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ according to Equation (4).
-To train the model on a forecasting task, where the goal is to predict a future sequence of events conditioned on a history $\mathcal{H}$, we embed $\mathcal{H}$ into a fixed-dimensional vector representation $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$ via a learned encoder $f_{\theta}$ before providing this to the model $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$ and minimizing Equation (7).
-Note that we jointly train the encoder $f_{\theta}$ and vector field $v_{\theta}$.
-In this case, we must also learn a model of the event count distribution $p_{\phi}(n\mid\mathcal{H})$.
-We train $p_{\phi}(n\mid\mathcal{H})$ by minimizing a cross-entropy loss.
-In practice, we found it important to regularize this loss to encourage smoothness in $n$, which we achieve through an optimal-transport based regularizer.
-As both $v_{\theta}$ and $p_{\phi}$ must be able to operate on variable-length sequences, we use a transformer-based backbone for both models.
-However, our overall method is agnostic to this choice.
-See Algorithm 1 and Appendix C.
 Sampling
 Once $v_{\theta}$ is learned, we may sample from the model by drawing a reference sequence $\gamma_{0}\sim\mu_{0}$ and solving the corresponding ODE parametrized by $v_{\theta}$.
-More concretely, we first fix a number of events $n$.
-When seeking to unconditionally generate new sequences from the underlying data TPP $\mu_{1}$, we simply sample $n$ from the empirical event count distribution $\mu_{1}(n)$.
+We fix a number of events $n$, which can be sampled from the empirical event count distribution $\mu_{1}(n)$ for unconditional generation.
 For conditional tasks, we draw $n\sim p_{\phi}(n\mid\mathcal{H})$ from the learned conditional distribution over event counts.
 Next, we draw $n$ initial events, corresponding to $s=0$, by sampling and sorting $t_{0}^{1},\dots,t_{0}^{n}\sim q$.
-In practice, we use $q=\mathcal{N}(0,I_{n})$ as we normalize our sequences into the range $[-1,1]$ during training and sampling (followed by renormalization to the data scale).
+In practice, we use $q=\mathcal{N}(0,I_{n})$ as we normalize our sequences into the range [1,1][-1,1] during training and sampling (followed by renormalization to the data scale).
 Since we have fixed $n$, we may view this initial draw as a vector $\gamma_{0}=[t_{0}^{1},\dots,t_{0}^{n}]\in\mathcal{T}^{n}$.
@@
 In our experiments, we use the forward Euler scheme, i.e., we specify a discretization $\{0=s_{0}<s_{1}<\dots<s_{K}=1\}$ of the flow time (in practice, uniform) and recursively compute
-$\gamma_{s_{k}}=\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1})\qquad k=1,2,\dots,K$ (8)
-where $h_{k}=s_{k}-s_{k-1}$ is a step size. While other choices are possible, we found that this simple scheme was sufficient as the model sample paths are typically close to linear. See Algorithm 2.
+$\gamma_{s_{k}}=\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1}),\quad k=1,\dots,K$ (11)
+where $h_{k}=s_{k}-s_{k-1}$ is a step size.
+While other choices are possible, we found that this simple scheme was sufficient as the model sample paths are, qualitatively, typically close to linear.
+See Algorithm 2.
 5 Experiments
-We study our proposed EventFlow model under two settings.
+Figure 2: Sequence distance (12) between the forecasted and ground-truth event sequences on a held-out test set.
+We report the mean $\pm$ one standard deviation over five random seeds.
+EventFlow (with 25 NFEs) achieves the lowest mean distance (forecasting error) for each of the 7 datasets.
+We study our proposed EventFlow model under two settings [^2] Code for our experiments is available at this URL: https://github.com/GavinKerrigan/eventflow.
 The first is a conditional task, where we seek to forecast both the number of future events and their times of occurrence, over a particular horizon, given a history.
@@
 We evaluate our model across a diverse set of datasets encompassing a wide range of possible point process behaviors.
-We evaluate our model on seven real-world datasets, which are a standard benchmark for modeling unmarked TPPs (Shchur et al., 2020b; Bosser & Taieb, 2023; Lüdke et al., 2023).
+We evaluate our model on seven real-world datasets, which are a standard benchmark for modeling unmarked TPPs (Shchur et al., 2020b; Bosser and Taieb, 2023; Lüdke et al., 2023).
 We also evaluate on a collection of six synthetic datasets (Omi et al., 2019).
-See Appendix A.
+See Appendix A for additional information regarding our datasets.
 Baseline Models
 We selected a set of baselines consisting of a set of diverse and highly performant neural TPP models.
-For an intensity-based method, we compare against the Neural Hawkes Process (NHP) (Mei & Eisner, 2017), a well-known and very widely-used neural TPP model.
+For an intensity-based method, we compare against the Neural Hawkes Process (NHP) (Mei and Eisner, 2017), a well-known and very widely-used neural TPP model.
 We additionally compare against two intensity-free methods, namely the flow-based IFTPP model (Shchur et al., 2020a) and the diffusion-based model of Lin et al.
 (2022).
-Lastly, our strongest baseline is the recently proposed Add-and-Thin model of Lüdke et al.
+Lastly, our strongest baseline is the Add-and-Thin model of Lüdke et al.
 (2023), which can be loosely viewed as a non-autoregressive diffusion model.
 These models use an RNN-based history encoder, with the exception of Add-and-Thin which uses a CNN-based encoder.
-See Appendix D for additional details.
+See Appendix D for additional details on our baselines.
 Metrics
 Evaluating generative TPP models is challenging, as one must take into account both the variable locations and numbers of events.
-This is particularly challenging for the unconditional setting, where unlike forecasting, we do not have a ground-truth sequence to compare against.
+This is particularly challenging for the unconditional setting, where unlike forecasting, we do not have a ground-truth sequence.
 Our starting point is a metric (Xiao et al., 2017a) on the space of sequences $\Gamma$, allowing us to measure the distance between two sequences $\gamma=\sum_{k=1}^{n}\delta[t^{k}_{\gamma}]$ and $\eta=\sum_{k=1}^{m}\delta[t^{k}_{\eta}]$ with possibly different numbers of events.
-Without loss of generality, we assume $n\leq m$, where the distance is given by
+Without loss of generality, we assume $n\leq m$, in which case the distance is given by
 $d(\gamma,\eta)=\sum_{k=1}^{n}|t^{k}_{\gamma}-t^{k}_{\eta}|+\sum_{k=n+1}^{m}(T-t^{k}_{\eta})$ (12)
@@
 This distance can be understood either as an $L^{1}$ distance between the counting processes of $\gamma,\eta$ or as a generalization of the $1$-Wasserstein distance to measures of unequal mass.
-For our unconditional experiment, we require a metric which will capture the distance between the TPP distributions themselves.
+For our unconditional experiment, we require a metric that will capture the distance between the TPP distributions themselves.
 To do so we use the distance in Equation (12) to calculate an MMD (Gretton et al., 2012; Shchur et al., 2020b).
-We use an exponential kernel $k(\gamma,\eta)=\exp\left(-d(\gamma,\eta)/(2\sigma^{2})\right)$ with $\sigma$ chosen to be the median distance between all sequences (Shchur et al., 2020b; Lüdke et al., 2023).
-Figure 2: Sequence distance (9) between the forecasted and ground-truth event sequences on a held-out test set.
-We report the mean $\pm$ one standard deviation over five random seeds.
-EventFlow (with 25 NFEs) achieves the lowest mean distance (forecasting error) for each of the 7 datasets.
+We use an exponential kernel $k(\gamma,\eta)=\exp\left(-d(\gamma,\eta)/(2\sigma^{2})\right)$ with $\sigma$ chosen to be the median distance between all sequences in a dataset (Shchur et al., 2020b; Lüdke et al., 2023).
 5.1 Forecasting Event Sequences
 We first evaluate our model on a multi-step forecasting task.
-We set a horizon $\Delta T$ for each of our real-world datasets and seek to generate event sequences in the range $[T_{0},T_{0}+\Delta T]$ for some given $T_{0}$, conditioned on the history $\mathcal{H}_{T_{0}}$.
+We set a horizon $\Delta T$ for each of our real-world datasets and seek to generate event sequences in the range $T_{0},T_{0}+\Delta T]$ for some given $T_{0}$, conditioned on the history $\mathcal{H}_{T_{0}}$.
 Up to a shift, this means we are taking $\mathcal{T}=[0,\Delta T]$.
@@
 At test time, we sample 50 values of $T_{0}$ for each test set sequence.
-We then forecast the sequence in $[T_{0},T_{0}+\Delta T]$ and compute the distance (9) between the ground-truth and generated sequences.
+We then generate one forecast for the sequence in $T_{0},T_{0}+\Delta T]$ and compute the distance (12) between the ground-truth and generated sequences.
 We note that the distance in Equation (12) is computed using $T_{0}+\Delta T$ rather than $T$ as the maximum event time, as using $T$ would result in a distance which is sensitive to the location of the forecasting window.
@@
 Our proposed EventFlow method obtains the lowest average forecasting error across all datasets, achieving an error which is 20%-53% lower than the nearest baseline.
-Given that the non-autoregressive models (EventFlow, Add-and-Thin) consistently outperform the autoregressive baselines, this is clear evidence that autoregressive models can struggle on multi-step predictions.
-This is especially true on the Reddit-C and Reddit-S datasets, which exhibit long sequence lengths.
-Table 1: Mean sequence distance per dataset for EventFlow as we vary the NFEs used to simulate the ODE.
-NFEs PUBG Red.C Red.S Taxi Twitter Yelp-A Yelp-M
-25 $2.0$ $15.8$ $16.0$ 3.23.2{3.2}3.2 1.41.4{1.4}1.4 1.31.3{1.3}1.3 1.91.9{1.9}1.9
+Given that the non-autoregressive models (EventFlow, Add-and-Thin) typically outperform the autoregressive baselines, we see clear evidence that autoregressive models can struggle on multi-step tasks.
+This is especially true on Reddit-C and Reddit-S which exhibit long sequence lengths.
+Table 1: Mean sequence distance for EventFlow as we vary the NFEs used to simulate the ODE.
+NFEs PUBG Red.C Red.S Taxi Tw. Yelp-A Yelp-M
+25 $2.0$ $15.8$ $16.0$ 3.2{3.2} 1.4{1.4} 1.3{1.3} 1.9{1.9}
 10 $2.0$ $15.8$ $15.8$ $3.1$ $1.4$ $1.3$ $1.9$
 1 $2.0$ $15.8$ $15.8$ $3.7$ $1.4$ $1.8$ $1.9$
+Table 2: MARE values evaluating the predicted number of events when forecasting. Mean values are reported over five random seeds.
+PUBG Red.C Red.S Taxi Tw. Yelp-A Yelp-M
+IFTPP $1.05$ $1.69$ $0.79$ $0.60$ $0.88$ $0.76$ $0.76$
+NHP $1.02$ $\underline{0.95}$ $1.00$ $0.67$ $2.48$ $0.80$ $1.07$
+Diffusion $1.95$ $1.28$ $1.12$ $0.49$ $0.66$ $0.65$ $0.72$
+Add-Thin $\underline{0.43}$ $0.99$ $\underline{0.38}$ $\underline{0.33}$ $\underline{0.60}$ $\mathbf{0.42}$ $\mathbf{0.46}$
+EventFlow $\mathbf{0.40}$ $\mathbf{0.70}$ $\mathbf{0.16}$ $\mathbf{0.28}$ $\mathbf{0.46}$ $\underline{0.56}$ $\underline{0.50}$
 In Table 1, we additionally vary the number of function evaluations (NFEs) used at test time, i.e., the number of steps $K$ used in the ODE solver in Algorithm 2.
@@
 Even with only a single NFE, EventFlow is able to obtain state-of-the-art performance.
-This is because, although the vector fields (6) in EventFlow are generally non-linear, we find that our construction produces paths which are approximately so.
-In contrast, Add-and-Thin uses 100 NFEs, while the NFEs for the autoregressive models scale linearly with the number of sampled events.
-Table 2: MARE values evaluating the predicted number of events when forecasting. Mean values are reported over five random seeds.
-PUBG Red.C Red.S Taxi Twitter Yelp-A Yelp-M
-IFTPP $1.05$ $1.69$ $0.79$ $0.60$ $0.88$ $0.76$ $0.76$
-NHP $1.02$ $\underline{0.95}$ $1.00$ $0.67$ $2.48$ $0.80$ $1.07$
-Diffusion $1.95$ $1.28$ $1.12$ $0.49$ $0.66$ $0.65$ $0.72$
-Add-and-Thin $\underline{0.43}$ $0.99$ $\underline{0.38}$ $\underline{0.33}$ $\underline{0.60}$ $\mathbf{0.42}$ $\mathbf{0.46}$
-EventFlow (ours) $\mathbf{0.40}$ $\mathbf{0.70}$ $\mathbf{0.16}$ $\mathbf{0.28}$ $\mathbf{0.46}$ $\underline{0.56}$ $\underline{0.50}$
-To evaluate the predictions of event counts, ${p_{\phi}(n\mid\mathcal{H})}$, we report the mean absolute relative error (MARE) between the true and predicted event counts.
-This is reported in Table 2.
-Our model achieves strong performance on this metric as it decouples the event count prediction from the event time prediction, allowing us to train explicitly for each task.
-The baselines, on the other hand, only model $n$ implicitly.
-See Appendix E for details and standard deviations.
-5.2 Unconditional Generation of Event Sequences
-Finally, we evaluate our model on an unconditional generation task, where we aim to generate new sequences from the underlying data distribution.
-In Table 3 we report MMD values for each of the synthetic and real-world datasets.
-See Appendix E for results with standard deviations.
-EventFlow (mean rank: $\mathbf{1.8}$) exhibits uniformly strong performance, obtaining either the best or second best MMD on 11 of the 13 datasets.
-While IFTPP (mean rank: $\mathbf{2.1}$) performs well on the synthetic datasets, it struggles with the more complex real-world datasets, reflecting its high bias.
-The Add-and-Thin method (mean rank: $\mathbf{2.4}$) is often similarly strong, but struggles on the SC dataset.
-While the NHP (mean rank: $\mathbf{3.7}$) can obtain good fits, this appears to be dataset dependent, with weak results on the NSR, SC, and Reddit-S datasets.
-The diffusion baseline (mean rank: $\mathbf{4.8}$) is our weakest baseline, which is perhaps unsurprising as this model can only be trained to maximize the likelihood of a subsequent event and not the overall sequence likelihood.
+This is due to the fact that, although the vector fields (Equation (9)) in EventFlow are generally non-linear, we find qualitatively that our construction produces paths which are approximately so.
+In contrast, Add-and-Thin uses 100 NFEs, while the NFEs for the autoregressive models scale linearly with the number of generated events.
 Table 3: MMDs (1e-2) between the test set and $1,000$ generated unconditional sequences averaged over five random seeds. EventFlow achieves the overall highest mean rank.
@@
 EventFlow (25 NFEs) $\underline{1.9}$ $\underline{2.2}$ $3.8$ $\mathbf{4.2}$ $\underline{8.3}$ $\underline{1.7}$ $\mathbf{1.5}$ $\mathbf{0.7}$ $\mathbf{0.7}$ $\mathbf{3.5}$ $4.9$ $\underline{6.6}$ $\mathbf{3.0}$
+To evaluate the event count predictions, $p_{\phi}(n\mid\mathcal{H})$}, we report the mean absolute relative error (MARE) between the true and predicted counts in Table 2.
+Our model achieves strong performance on this metric as it decouples the event count prediction from the event time prediction, training explicitly for each task.
+The baselines, on the other hand, only model $n$ implicitly.
+See Appendix E for details and standard deviations.
+5.2 Unconditional Generation of Event Sequences
+Finally, we evaluate our model on an unconditional generation task, where we aim to generate new sequences from the underlying data distribution without conditioning on an observed history.
+In Table 3 we report MMD values for each of the synthetic and real-world datasets.
+MMDs are calculated by sampling 1,000 sequences from each trained model.
+The MMD values in Table 3 correspond to the mean test set MMD ($\pm$ one standard deviation) across five random seeds.
+See Appendix E for results with standard deviations.
+EventFlow (mean rank: $\mathbf{1.8}$) exhibits uniformly strong performance, obtaining either the best or second best MMD on 11 of the 13 datasets.
+This is particularly pronounced on the real-world datasets, where we obtain the lowest MMD on 5 of the 7 datasets.
+Notably, IFTPP (mean rank: $\mathbf{2.1}$) is one of the strongest methods on the synthetic data, but is relatively weak on the real-world data.
+We believe that this can be understood in terms of model flexibility.
+IFTPP uses a simple parametric form (a mixture of log-normals) to parametrize the subsequent event time distribution (see Section 3).
+This provides a strong inductive bias and allows it to fit simple synthetic distributions very effectively, but limits its capability to model more realistic data sets.
+EventFlow, in contrast, makes no parametric assumptions, offering greater flexibility.
+As a result, it may not outperform simple parametric models on very simple synthetic datasets, but this flexibility is precisely what enables strong performance on complex real-world data, which we consider the more meaningful benchmark.
+The Add-and-Thin method (mean rank: $\mathbf{2.4}$) is often similarly strong, but struggles on the SC dataset.
+While the NHP (mean rank: $\mathbf{3.7}$) can obtain good fits, this appears to be dataset dependent, with weak results on the NSR, SC, and Reddit-S datasets.
+The diffusion baseline (mean rank: $\mathbf{4.8}$) is our weakest baseline, which is perhaps unsurprising as this model can only be trained to maximize the likelihood of a subsequent event and not the overall sequence likelihood.
 6 Conclusion
-In this work, we propose EventFlow, a non-autoregressive generative model for temporal point processes.
-We demonstrate that EventFlow is able to achieve state-of-the-art results on a multi-step forecasting task and strong performance on unconditional generation.
-Our work lays a foundation for future flow-based TPPs, and exploring applications in tasks like imputation, marked TPPs, and querying (Boyd et al., 2023) are exciting directions.
+We introduced EventFlow, a non-autoregressive generative model for temporal point processes.
+Our approach extends flow matching to the TPP setting by constructing a continuous transformation between a simple reference distribution and the data distribution over event sequences.
+Concretely, EventFlow generates event sequences by first drawing a sequence of event times from a reference process, and then transporting these events along a learned vector field.
+Beyond modeling event times, our framework also captures the distribution over the number of events, which we learn via a regularized cross-entropy objective.
+Empirically, we demonstrate that EventFlow achieves state-of-the-art performance on multi-step forecasting tasks and competitive results on unconditional generation across a range of standard benchmarks.
+These results highlight the promise of non-autoregressive, flow-based approaches as a flexible alternative to classical intensity-based models for TPPs.
 Limitations and Outlook
-In this work, we have focused on modeling the event times, and thus do not experiment with marked TPPs or spatiotemporal TPPs.
-Extending EventFlow to these tasks is a promising directions.
-We also note that we do not explicitly enforce that the support of our model is $[0,T]$, which would necessitate moving beyond the Gaussian setting.
-Moreover, our baseline methods use the neural architectures proposed in the original works, which vary across the methods.
-Carefully studying the effects of these architecture choices is an important direction.
-Acknowledgments
-This research was supported by the National Institutes of Health under award R01-LM013344 and by the Hasso Plattner Institute (HPI) Research Center in Machine Learning and Data Science at the University of California, Irvine.
+Our work focuses on modeling event times, and we do not consider marked or spatiotemporal point processes.
+Extending EventFlow to incorporate marks or spatial structure is a natural and promising direction for future work, though it may require new coupling and interpolation strategies.
+Additionally, our current construction does not explicitly enforce support constraints on $[0,T]$, as it relies on Gaussian noising in Equation (6).
+Addressing this limitation, e.g., by designing constrained flows, could potentially lead to further gains in performance.
+Finally, our experimental comparisons adopt the neural architectures used in prior work, which differ across baseline methods.
+A more controlled study of architectural choices, and their interaction with flow-based objectives, would help isolate the gains attributable to the modeling framework itself.
+More broadly, we believe our work opens the door to a new class of generative models for point processes based on flow-based ideas, and we hope it encourages further exploration of non-autoregressive approaches.
+Acknowledgements
+We thank the reviewers for their feedback on improving the paper.
+This work was supported by National Science Foundation under awards NSF 2505006 and NSF 2425932, by the National Institutes of Health under awards R01-LM013344 and R01CA297869, by the UK EPSRC grant EP/Y037200/1, by the Hasso Plattner Institute (HPI) Research Center in Machine Learning and Data Science at UCI, and by funding support from Google and from SAP.
+Supplementary Materials
 Appendix A Datasets
@@
 Our synthetic datasets are adopted from those proposed by Omi et al.
-(2019).
+[2019].
 Each of these datasets consists of $1,000$ sequences supported on $\mathcal{T}=[0,100]$.
@@
 We refer to Section 4 of Omi et al.
-(2019) for details.
+[2019] for details.
 Real-World Datasets
 We use the set of real-world datasets proposed in Shchur et al.
-(2020b), which constitute a set of standard benchmark datasets for unmarked TPPs.
+[2020b], which constitute a set of standard benchmark datasets for unmarked TPPs.
 We refer to Appendix D of Shchur et al.
-(2020b) for additional details.
+[2020b] for additional details.
 With the exception of PUBG, these datasets are supported on $\mathcal{T}=[0,24]$, i.e.
@@
 For selecting the forecasting horizon $\Delta T$, we follow the choice made in Lüdke et al.
-(2023) to facilitate comparison and reproducibility.
+[2023] to facilitate comparison and reproducibility.
 We note the initial times $T_{0}\in[\Delta T,T-\Delta T]$ are selected to be at least $\Delta T$ to avoid forecasting sequences with no observations in the history.
@@
 Sequences Mean length Std length Range length Support $\Delta T$
-Hawkes1 1000 95.4 45.8 [14, 300] [0, 100] {-}-
-Hawkes2 1000 97.2 49.1 [18, 355] [0, 100] {-}-
-Nonstationary Poisson 1000 100.3 9.8 [71, 134] [0, 100] {-}-
-Nonstationary Renewal 1000 98 2.9 [86, 100] [0, 100] {-}-
-Stationary Renewal 1000 109.2 38.1 [1, 219] [0, 100] {-}-
-Self-Correcting 1000 100.3 0.74 [98, 102] [0, 100] {-}-
-PUBG 3001 76.5 8.8 [26, 97] [0, 38] $5$
-Reddit-C 1356 295.7 317.9 [1, 2137] [0, 24] $4$
-Reddit-S 1094 1129 359.5 [363, 2658] [0, 24] $4$
-Taxi 182 98.4 20 [12, 140] [0, 24] $4$
-Twitter 2019 14.9 14 [1, 169] [0, 24] $4$
-Yelp-Airport 319 30.5 7.5 [9, 55] [0, 24] $4$
-Yelp-Miss. 319 55.2 15.9 [3, 107] [0, 24] $4$
+Hawkes1 1000 95.4 45.8 [14, 300] [0, 100] {-}
+Hawkes2 1000 97.2 49.1 [18, 355] [0, 100] {-}
+Nonstationary Poisson 1000 100.3 9.8 [71, 134] [0, 100] {-}
+Nonstationary Renewal 1000 98 2.9 [86, 100] [0, 100] {-}
+Stationary Renewal 1000 109.2 38.1 [1, 219] [0, 100] {-}
+Self-Correcting 1000 100.3 0.74 [98, 102] [0, 100] {-}
+PUBG 3001 76.5 8.8 [26, 97] [0, 38] 55
+Reddit-C 1356 295.7 317.9 [1, 2137] [0, 24] 44
+Reddit-S 1094 1129 359.5 [363, 2658] [0, 24] 44
+Taxi 182 98.4 20 [12, 140] [0, 24] 44
+Twitter 2019 14.9 14 [1, 169] [0, 24] 44
+Yelp-Airport 319 30.5 7.5 [9, 55] [0, 24] 44
+Yelp-Miss. 319 55.2 15.9 [3, 107] [0, 24] 44
 Appendix B Proofs
@@
 Proof.
-Let $A_{1},A_{2}\subseteq\Gamma$ be Borel measurable (Daley & Vere-Jones, 2003, Prop.
-5.3) subsets of the configuration space $\Gamma$, i.e.
+Let $A_{1},A_{2}\subseteq\Gamma$ be Borel measurable [Daley and Vere-Jones, 2003, Prop.
+5.3] subsets of the configuration space $\Gamma$, i.e.
 each of $A_{1},A_{2}$ is a measurable collection of event sequences.
@@
 Here, we provide additional details regarding the parametrization and training of our EventFlow model.
-In general, our model is based on the transformer architecture (Vaswani et al., 2017; Yang et al., 2022), due to its general ability to handle variable length inputs and outputs, high flexibility, and ability to incorporate long-range interactions.
-We emphasize that this is an effective architecture choice, but our method is not necessarily tied to this architecture, i.e., other choices like CNNs or RNNs are certainly possible here.
+Our model is based on the transformer architecture [Vaswani et al., 2017, Yang et al., 2022], due to its general ability to handle variable length inputs and outputs, high flexibility, and ability to incorporate long-range interactions.
+We emphasize that this is an effective architecture choice, but our method is not necessarily tied to this architecture.
 In all settings, our reference measure $\mu_{0}$ is specified with $q=\mathcal{N}(0,I)$.
@@
 These position indices are handled by sinusoidal embeddings followed by an additional linear layer.
-There are three linear layers in total – one for the flow time, one shared across the sequence times, and one for the position indices.
+There are three linear layers in total: one for the flow time, one shared across the sequence times, and one for the position indices.
 These embeddings are added together to create a representation for each element of the sequence, and we apply a standard transformer to this sequence to produce a sequence of vectors of length $N(\gamma_{s})$.
@@
 We first embed the history sequence times $\mathcal{H}$ and the sequence position indices in a manner analogous to the above.
-In addition, the model was provided the start of the prediction window $T_{0}$ by concatenating it as the final event in $\mathcal{H}$.
+The model was provided the start of the prediction window $T_{0}$ by concatenating it as the final event in $\mathcal{H}$.
 This yielded better results than encoding the start of the prediction window separately.
@@
 $p_{\phi}(n\mid\mathcal{H})$, is again based on the transformer decoder, sharing the same overall architecture as our unconditional model.
-However, the key difference is that we instead take a mean of the final sequence embeddings before passing this through a small MLP to produce the final logit.
+The key difference is that we instead take a mean of the final sequence embeddings before passing this through a small MLP to produce the final logit.
 See Figure 5.
@@
 Training and Tuning
-We normalize all sequences to the range $[-1,1]$, using the overall min/max event time seen in the training data.
+We normalize all sequences to the range [1,1][-1,1], using the overall min/max event time seen in the training data.
 All sequences are generated on this normalized scale, prior to re-scaling the sequence back to the original data range before evaluation.
-Our models are trained with the Adam (Kingma & Ba, 2015) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for $30,000$ steps with a cosine scheduler (Loshchilov & Hutter, 2017), which cycled every $10,000$ steps.
+Our models are trained with the Adam [Kingma and Ba, 2015] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for $30,000$ steps with a cosine scheduler [Loshchilov and Hutter, 2017], which cycled every $10,000$ steps.
 Final hyperparameters were selected by best performance on the validation dataset achieved at any point during the training, where models were evaluated 10 times throughout their training.
 To tune our model, we performed a grid search over learning rates in $\{5\times 10^{-3},10^{-3},5\times 10^{-4}\}$ and dropout probabilities in $\{0,0.1,0.2\}$.
-Overall, we found that learning rates of $10^{-2}$ or larger often caused the model to diverge, and a dropout of $0.1$ yielded the best results across all settings.
+Overall, we found that learning rates of $10^{-2}$ or larger often caused the model to diverge.
 We use $6$ transformer layers, $8$ attention heads, and an embedding dimension of $512$ across all settings, except for the Reddit-C and Reddit-S datasets where we use $4$ heads and an embedding dimension of $128$ due to the increased memory cost of these datasets.
@@
 Table 5: The best hyperparameter settings found for the vector field $v_{\theta}$ in our EventFlow method on the unconditional generation task.
-Learning Rate Emb. Dim. MLP Dim Heads Transformer Layers
-Hawkes1 $10^{-3}$ $512$ $2048$ $8$ $6$
-Hawkes2 $10^{-3}$ $512$ $2048$ $8$ $6$
-Nonstationary Poisson $10^{-3}$ $512$ $2048$ $8$ $6$
-Nonstationary Renewal $10^{-3}$ $512$ $2048$ $8$ $6$
-Stationary Renewal $10^{-3}$ $512$ $2048$ $8$ $6$
-Self-Correcting $10^{-3}$ $512$ $2048$ $8$ $6$
-PUBG $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
-Reddit-C $10^{-3}$ $128$ $256$ $4$ $6$
-Reddit-S $5\times 10^{-3}$ $128$ $256$ $4$ $6$
-Taxi $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
-Twitter $10^{-3}$ $512$ $2048$ $8$ $6$
-Yelp-Airport $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
-Yelp-Miss. $10^{-3}$ $512$ $2048$ $8$ $6$
+Learning Rate Emb. Dim. MLP Dim Heads Transformer Layers Dropout
+Hawkes1 $10^{-3}$ $512$ $2048$ 88 66 0.1
+Hawkes2 $10^{-3}$ $512$ $2048$ 88 66 0.1
+Nonstationary Poisson $10^{-3}$ $512$ $2048$ 88 66 0.1
+Nonstationary Renewal $10^{-3}$ $512$ $2048$ 88 66 0.1
+Stationary Renewal $10^{-3}$ $512$ $2048$ 88 66 0.1
+Self-Correcting $10^{-3}$ $512$ $2048$ 88 66 0.1
+PUBG $5\times 10^{-4}$ $512$ $2048$ 88 66 0.1
+Reddit-C $10^{-3}$ $128$ $256$ 44 66 0.1
+Reddit-S $5\times 10^{-3}$ $128$ $256$ 44 66 0.1
+Taxi $5\times 10^{-4}$ $512$ $2048$ 88 66 0.1
+Twitter $10^{-3}$ $512$ $2048$ 88 66 0.1
+Yelp-Airport $5\times 10^{-4}$ $512$ $2048$ 88 66 0.1
+Yelp-Miss. $10^{-3}$ $512$ $2048$ 88 66 0.1
 Table 6: The best hyperparameter settings found for the vector field $v_{\theta}$ in our EventFlow method on the forecasting task.
-Learning Rate Emb. Dim. MLP Dim. Heads Transformer Layers
-PUBG $10^{-3}$ $512$ $2048$ $8$ $6$
-Reddit-C $10^{-3}$ $128$ $256$ $4$ $6$
-Reddit-S $10^{-3}$ $128$ $256$ $4$ $6$
-Taxi $10^{-3}$ $512$ $2048$ $8$ $6$
-Twitter $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
-Yelp-Airport $10^{-3}$ $512$ $2048$ $8$ $6$
-Yelp-Miss. $10^{-3}$ $512$ $2048$ $8$ $6$
+Learning Rate Emb. Dim. MLP Dim. Heads Transformer Layers Dropout
+PUBG $5\times 10^{-3}$ $512$ $2048$ 88 66 $0.1$
+Reddit-C $10^{-3}$ $128$ $256$ 44 66 $0.2$
+Reddit-S $10^{-3}$ $128$ $256$ 44 66 $0.2$
+Taxi $5\times 10^{-4}$ $512$ $2048$ 88 66 $0.1$
+Twitter $5\times 10^{-4}$ $512$ $2048$ 88 66 $0.1$
+Yelp-Airport $10^{-3}$ $512$ $2048$ 88 66 $0.2$
+Yelp-Miss. $5\times 10^{-4}$ $512$ $2048$ 88 66 $0.1$
 Table 7: The best hyperparameter settings found for the event count predictor $p_{\phi}(n\mid\mathcal{H})$ in our EventFlow method on the forecasting task.
-Learning Rate $\alpha\times N_{max}$ Emb. Dim. MLP Dim. Heads Transformer Layers
-PUBG $10^{-3}$ $1.0$ $512$ $2048$ $8$ $6$
-Reddit-C $10^{-3}$ $1000.0$ $128$ $256$ $4$ $6$
-Reddit-S $10^{-3}$ $1000.0$ $128$ $256$ $4$ $6$
-Taxi $10^{-3}$ $1000.0$ $512$ $2048$ $8$ $6$
-Twitter $10^{-3}$ $1000.0$ $512$ $2048$ $8$ $6$
-Yelp-Airport $10^{-3}$ $1000.0$ $512$ $2048$ $8$ $6$
-Yelp-Miss. $10^{-3}$ $100.0$ $512$ $2048$ $8$ $6$
+Learning Rate $\alpha/N_{max}$ Emb. Dim. MLP Dim. Heads Transformer Layers Dropout
+PUBG $^{-3} 1.01.0$ $512$ $2048$ 88 66 $0.2$
+Reddit-C $10^{-3}$ $1000.0$ $128$ $256$ 44 66 $0.2$
+Reddit-S $10^{-3}$ $1000.0$ $128$ $256$ 44 66 $0.0$
+Taxi $10^{-3}$ $1000.0$ $512$ $2048$ 88 66 $0.2$
+Twitter $10^{-3}$ $1000.0$ $512$ $2048$ 88 66 $0.2$
+Yelp-Airport $10^{-3}$ $1000.0$ $512$ $2048$ 88 66 $0.2$
+Yelp-Miss. $10^{-3}$ $100.0$ $512$ $2048$ 88 66 $0.2$
 Appendix D Additional Details on Baselines
@@
 All methods are trained at a batch size of $64$ for $1,000$ epochs, using early stopping on the validation set loss.
-In early experiments, we also evaluated AttNHP (Zuo et al., 2020), a variant of the NHP which uses an attention-based encoder, but found it to be prohibitively expensive in terms of memory cost (requiring more than 24 GB of VRAM) and, as a result, do not include it in our results.
+In early experiments, we also evaluated AttNHP [Zuo et al., 2020], a variant of the NHP which uses an attention-based encoder, but found it to be prohibitively expensive in terms of memory cost (requiring more than 24 GB of VRAM) and, as a result, do not include it in our results.
 IFTPP
 Our first baseline is the intensity-free TPP model of Shchur et al.
-(2020a).
+[2020a].
 This model uses an RNN encoder and a mixture of log-normal distributions to parametrize the decoder.
@@
 Learning Rate Weight Decay Embedding Dimension Mixture Components
-Hawkes1 $10^{-3}$ $10^{-4}$ $32$ $8$
-Hawkes2 $10^{-2}$ 00 $32$ $8$
-Nonstationary Poisson $10^{-3}$ $10^{-6}$ $128$ $8$
+Hawkes1 $10^{-3}$ $10^{-4}$ $32$ 88
+Hawkes2 $10^{-2}$ 0 $32$ 88
+Nonstationary Poisson $10^{-3}$ $10^{-6}$ $128$ 88
 Nonstationary Renewal $10^{-2}$ $10^{-6}$ $64$ $16$
-Stationary Renewal $10^{-3}$ $10^{-4}$ $32$ $8$
+Stationary Renewal $10^{-3}$ $10^{-4}$ $32$ 88
 Self-Correcting $10^{-3}$ $10^{-6}$ $32$ $64$
-PUBG $10^{-2}$ 00 $128$ $32$
+PUBG $10^{-2}$ 0 $128$ $32$
 Reddit-C $10^{-3}$ $10^{-4}$ $64$ $16$
@@
 Taxi $10^{-2}$ $10^{-5}$ $128$ $64$
-Twitter $10^{-3}$ $10^{-4}$ $64$ $6$
+Twitter $10^{-3}$ $10^{-4}$ $64$ 66
 Yelp-Airport $10^{-2}$ $10^{-6}$ $64$ $64$
-Yelp-Miss. $10^{-3}$ $10^{-4}$ $32$ $8$
+Yelp-Miss. $10^{-3}$ $10^{-4}$ $32$ 88
 Table 9: The best hyperparameter settings found for IFTPP on the forecasting task.
@@
 PUBG $10^{-4}$ $10^{-6}$ $32$ $32$
-Reddit-C $10^{-2}$ 00 $64$ $8$
-Reddit-S $10^{-2}$ 00 $64$ $16$
-Taxi $10^{-3}$ $10^{-6}$ $128$ $8$
-Twitter $10^{-2}$ $10^{-5}$ $32$ $8$
+Reddit-C $10^{-2}$ 0 $64$ 88
+Reddit-S $10^{-2}$ 0 $64$ $16$
+Taxi $10^{-3}$ $10^{-6}$ $128$ 88
+Twitter $10^{-2}$ $10^{-5}$ $32$ 88
 Yelp-Airport $10^{-2}$ $10^{-6}$ $128$ $32$
-Yelp-Miss. $10^{-2}$ $10^{-6}$ $32$ $8$
+Yelp-Miss. $10^{-2}$ $10^{-6}$ $32$ 88
 NHP
-We additionally compare against the Neural Hawkes Process of Mei & Eisner (2017).
+We additionally compare against the Neural Hawkes Process of Mei and Eisner [2017].
 This model uses an LSTM encoder and a parametric form, whose weights are modeled by a neural network, to model the conditional intensity function.
-In practice, we use the implementation proved by the EasyTPP benchmark (Xue et al., 2024), as this version implements the necessary thinning algorithm for sampling.
+In practice, we use the implementation proved by the EasyTPP benchmark [Xue et al., 2024], as this version implements the necessary thinning algorithm for sampling.
 [^4] URL: https://github.com/ant-research/EasyTemporalPointProcess We perform a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$ and embedding dimensions in $\{32,64,128\}$.
@@
 Our diffusion baseline is based on the implementation of Lin et al.
-(2022), and our decoder model architecture is taken directly from the code of Lin et al.
-(2022).
-[^4] URL: https://github.com/EDAPINENUT/GNTPP At a high level, this model is a discrete-time diffusion model (Ho et al., 2020) trained to generate a single inter-arrival time given a history embedding.
+[2022], and our decoder model architecture is taken directly from the code of Lin et al.
+[2022].
+[^5] URL: https://github.com/EDAPINENUT/GNTPP At a high level, this model is a discrete-time diffusion model [Ho et al., 2020] trained to generate a single inter-arrival time given a history embedding.
 Note that as the likelihood is not available in diffusion models, the CDF in the likelihood in Equation (3) is not tractable.
@@
 In preliminary experiments, we found that the codebase provided by Lin et al.
-(2022) often produced NaN values during sampling, prompting us to make several changes.
+[2022] often produced NaN values during sampling, prompting us to make several changes.
 First, we use the RNN encoder from Shchur et al.
-(2020a), i.e.
+[2020a], i.e.
 the same encoder as the IFTPP baseline, to reduce the memory requirements of the model.
 Second, we do not log-scale the inter-arrival times as suggested by Lin et al.
-(2022), as we found that this often led to overflow and underflow issues at sampling time.
-Third, we do not normalize the data via standardization (i.e., subtracting off the mean inter-arrival time and dividing by the standard deviation), but rather, we scale the inter-arrival times so that they are in the bounded range $[-1,1]$.
-This is aligned with standard diffusion implementations (Ho et al., 2020), and allows us to perform clipping at sampling time to avoid the accumulation of errors.
-With these changes, our diffusion baseline is competitive, and able to obtain stronger results than previous work has reported (Lüdke et al., 2023).
-We use $1000$ diffusion steps and the cosine beta schedule (Nichol & Dhariwal, 2021), and we train the model on the simplified $\epsilon$-prediction loss of Ho et al.
-(2020).
+[2022], as we found that this often led to overflow and underflow issues at sampling time.
+Third, we do not normalize the data via standardization (i.e., subtracting off the mean inter-arrival time and dividing by the standard deviation), but rather, we scale the inter-arrival times so that they are in the bounded range [1,1][-1,1].
+This is aligned with standard diffusion implementations [Ho et al., 2020], and allows us to perform clipping at sampling time to avoid the accumulation of errors.
+With these changes, our diffusion baseline is competitive, and able to obtain stronger results than previous work has reported [Lüdke et al., 2023].
+We use $1000$ diffusion steps and the cosine beta schedule [Nichol and Dhariwal, 2021], and we train the model on the simplified $\epsilon$-prediction loss of Ho et al.
+[2020].
 We train for $1,000$ epochs with early stopping based on the validation set loss.
@@
 Learning Rate Weight Decay Embedding Dimension Layers
-Hawkes1 $10^{-3}$ $10^{-6}$ $64$ $2$
-Hawkes2 $10^{-2}$ $10^{-5}$ $64$ $4$
-Nonstationary Poisson $10^{-3}$ $10^{-5}$ $128$ $2$
-Nonstationary Renewal $10^{-3}$ $10^{-4}$ $64$ $2$
-Stationary Renewal $10^{-2}$ 00 $32$ $6$
-Self-Correcting $10^{-3}$ 00 $32$ $6$
-PUBG $10^{-3}$ 00 $64$ $2$
-Reddit-C $10^{-3}$ $10^{-6}$ $128$ $4$
-Reddit-S $10^{-3}$ 00 $64$ $4$
-Taxi $10^{-2}$ 00 $128$ $4$
-Twitter $10^{-3}$ $10^{-4}$ $64$ $6$
-Yelp-Airport $10^{-2}$ 00 $32$ $2$
-Yelp-Miss. $10^{-2}$ $10^{-5}$ $128$ $2$
+Hawkes1 $10^{-3}$ $10^{-6}$ $64$ 22
+Hawkes2 $10^{-2}$ $10^{-5}$ $64$ 44
+Nonstationary Poisson $10^{-3}$ $10^{-5}$ $128$ 22
+Nonstationary Renewal $10^{-3}$ $10^{-4}$ $64$ 22
+Stationary Renewal $10^{-2}$ 0 $32$ 66
+Self-Correcting $10^{-3}$ 0 $32$ 66
+PUBG $10^{-3}$ 0 $64$ 22
+Reddit-C $10^{-3}$ $10^{-6}$ $128$ 44
+Reddit-S $10^{-3}$ 0 $64$ 44
+Taxi $10^{-2}$ 0 $128$ 44
+Twitter $10^{-3}$ $10^{-4}$ $64$ 66
+Yelp-Airport $10^{-2}$ 0 $32$ 22
+Yelp-Miss. $10^{-2}$ $10^{-5}$ $128$ 22
 Table 13: The best hyperparameter settings found for diffusion on the forecasting task.
 Learning Rate Weight Decay Embedding Dimension Layers
-PUBG $10^{-4}$ $10^{-5}$ $32$ $6$
-Reddit-C $10^{-2}$ $10^{-6}$ $64$ $6$
-Reddit-S $10^{-3}$ 00 $64$ $4$
-Taxi $10^{-3}$ $10^{-6}$ $32$ $2$
-Twitter $10^{-4}$ $10^{-5}$ $64$ $6$
-Yelp-Airport $10^{-4}$ $10^{-5}$ $64$ $6$
-Yelp-Miss. $10^{-3}$ $10^{-5}$ $32$ $4$
+PUBG $10^{-4}$ $10^{-5}$ $32$ 66
+Reddit-C $10^{-2}$ $10^{-6}$ $64$ 66
+Reddit-S $10^{-3}$ 0 $64$ 44
+Taxi $10^{-3}$ $10^{-6}$ $32$ 22
+Twitter $10^{-4}$ $10^{-5}$ $64$ 66
+Yelp-Airport $10^{-4}$ $10^{-5}$ $64$ 66
+Yelp-Miss. $10^{-3}$ $10^{-5}$ $32$ 44
 Add-and-Thin
 We compare to the Add-and-Thin model of Lüdke et al.
-(2023) as a recently proposed non-autoregressive baseline.
+[2023] as a recently proposed non-autoregressive baseline.
 We directly run the code provided by the authors without additional modifications.
 [^6] URL: https://github.com/davecasp/add-thin We do, however, perform a slightly larger hyperparameter sweep than Lüdke et al.
-(2023), in order to ensure a fair comparison between the methods considered.
+[2023], in order to ensure a fair comparison between the methods considered.
 We train for $1,000$ epochs with early stopping on the validation loss.
@@
 Nonstationary Poisson $10^{-2}$ $16$
-Nonstationary Renewal $10^{-2}$ $8$
-Stationary Renewal $10^{-2}$ $8$
-Self-Correcting $10^{-4}$ $8$
-PUBG $10^{-3}$ $8$
+Nonstationary Renewal $10^{-2}$ 88
+Stationary Renewal $10^{-2}$ 88
+Self-Correcting $10^{-4}$ 88
+PUBG $10^{-3}$ 88
 Reddit-C $10^{-2}$ $32$
 Reddit-S $10^{-2}$ $16$
-Taxi $10^{-2}$ $8$
+Taxi $10^{-2}$ 88
 Twitter $10^{-4}$ $32$
-Yelp-Airport $10^{-4}$ $8$
+Yelp-Airport $10^{-4}$ 88
 Yelp-Miss. $10^{-2}$ $64$
@@
 Reddit-S $10^{-2}$ $64$
-Taxi $10^{-2}$ $8$
-Twitter $10^{-3}$ $8$
+Taxi $10^{-2}$ 88
+Twitter $10^{-3}$ 88
 Yelp-Airport $10^{-2}$ $32$
@@
 NHP $1.02{\scriptstyle\pm 0.08}$ $\underline{0.95}{\scriptstyle\pm 0.01}$ $1.00{\scriptstyle\pm 0.0004}$ $0.67{\scriptstyle\pm 0.11}$ $2.48{\scriptstyle\pm 0.40}$ $0.80{\scriptstyle\pm 0.22}$ $1.07{\scriptstyle\pm 0.34}$
-Diffusion $1.95{\scriptstyle\pm 0.48}$ $1.28{\scriptstyle\pm 0.09}$ $1.12{\scriptstyle\pm 0.56}$ $0.49{\scriptstyle\pm 0.07}$ $0.66{\scriptstyle\pm 0.04}$ $0.65{\scriptstyle\pm 0.07}$ $0.72{\scriptstyle\pm 0.07}$
+Diffusion $1.95{\scriptstyle\pm 0.48}$ $28{\scriptstyle\pm 0.09} 1.1$$12{\scriptstyle\pm 0.56}$ $0.49{\scriptstyle\pm 0.07}$ $0.66{\scriptstyle\pm 0.04}$ $0.65{\scriptstyle\pm 0.07}$ $0.72{\scriptstyle\pm 0.07}$
 Add-and-Thin $\underline{0.43}{\scriptstyle\pm 0.01}$ $0.99{\scriptstyle\pm 0.10}$ $\underline{0.38}{\scriptstyle\pm 0.01}$ $\underline{0.33}{\scriptstyle\pm 0.02}$ $\underline{0.60}{\scriptstyle\pm 0.02}$ $\mathbf{0.42}{\scriptstyle\pm 0.01}$ $\mathbf{0.46}{\scriptstyle\pm 0.03}$
@@
 IFTPP $5.7{\scriptstyle\pm 1.8}$ $1.3{\scriptstyle\pm 1.2}$ $\underline{1.9}{\scriptstyle\pm 0.6}$ $5.8{\scriptstyle\pm 0.9}$ $\mathbf{2.9}{\scriptstyle\pm 0.6}$ $8.2{\scriptstyle\pm 4.7}$ $\underline{5.1}{\scriptstyle\pm 0.7}$
-NHP $7.2{\scriptstyle\pm 0.2}$ $2.2{\scriptstyle\pm 1.6}$ $22.5{\scriptstyle\pm 0.3}$ $\underline{5.0}{\scriptstyle\pm 0.1}$ $7.3{\scriptstyle\pm 0.7}$ $6.7{\scriptstyle\pm 1.5}$ $6.1{\scriptstyle\pm 2.3}$
+NHP $7.2{\scriptstyle\pm 0.2}$ $2{\scriptstyle\pm 1.6} 22.$$2.5{\scriptstyle\pm 0.3}$ $\underline{5.0}{\scriptstyle\pm 0.1}$ $7.3{\scriptstyle\pm 0.7}$ $6.7{\scriptstyle\pm 1.5}$ $6.1{\scriptstyle\pm 2.3}$
 Diffusion $14.3{\scriptstyle\pm 6.5}$ $3.9{\scriptstyle\pm 1.2}$ $6.2{\scriptstyle\pm 3.3}$ $11.7{\scriptstyle\pm 1.8}$ $12.5{\scriptstyle\pm 1.9}$ $10.9{\scriptstyle\pm 3.8}$ $10.5{\scriptstyle\pm 5.2}$
@@
 Diffusion $4.8{\scriptstyle\pm 2.7}$ $5.5{\scriptstyle\pm 3.3}$ $10.8{\scriptstyle\pm 7.5}$ $15.0{\scriptstyle\pm 3.6}$ $9.1{\scriptstyle\pm 1.8}$ $5.1{\scriptstyle\pm 2.8}$
-Add-and-Thin $\underline{1.9}{\scriptstyle\pm 0.5}$ $2.5{\scriptstyle\pm 0.3}$ $\underline{2.6}{\scriptstyle\pm 0.5}$ $7.4{\scriptstyle\pm 1.2}$ $22.5{\scriptstyle\pm 0.5}$ $2.2{\scriptstyle\pm 0.8}$
+Add-and-Thin $\underline{1.9}{\scriptstyle\pm 0.5}$ $2.5{\scriptstyle\pm 0.3}$ $\underline{2.6}{\scriptstyle\pm 0.5}$ $7.4{\scriptstyle\pm 1.2}$ $2.5{\scriptstyle\pm 0.5} 2.2$$2{\scriptstyle\pm 0.8}$
 EventFlow (ours) $\underline{1.9}{\scriptstyle\pm 0.2}$ $\underline{2.2}{\scriptstyle\pm 0.1}$ $3.8{\scriptstyle\pm 1.2}$ $\mathbf{4.2}{\scriptstyle\pm 0.5}$ $\underline{8.3}{\scriptstyle\pm 0.4}$ $\underline{1.7}{\scriptstyle\pm 0.3}$
@@
 PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
-IFTPP $0.85{\scriptstyle\pm 0.05}$ $\underline{0.32}{\scriptstyle\pm 0.03}$ $0.0047{\scriptstyle\pm 0.0006}$ $0.22{\scriptstyle\pm 0.03}$ $1.74{\scriptstyle\pm 0.10}$ $1.24{\scriptstyle\pm 0.16}$ $1.11{\scriptstyle\pm 0.17}$
-NHP $0.89{\scriptstyle\pm 0.09}$ $0.53{\scriptstyle\pm 0.24}$ $\mathbf{0.0022}{\scriptstyle\pm 0.0007}$ $0.31{\scriptstyle\pm 0.12}$ $2.00{\scriptstyle\pm 0.30}$ $1.30{\scriptstyle\pm 0.26}$ $\underline{1.03}{\scriptstyle\pm 0.35}$
-Diffusion $\mathbf{0.61}{\scriptstyle\pm 0.10}$ $0.33{\scriptstyle\pm 0.04}$ $\underline{0.0037}{\scriptstyle\pm 0.0012}$ $0.23{\scriptstyle\pm 0.14}$ $\mathbf{1.30}{\scriptstyle\pm 0.21}$ $\mathbf{0.86}{\scriptstyle\pm 0.18}$ $\mathbf{0.92}{\scriptstyle\pm 0.20}$
+IFTPP $0.85{\scriptstyle\pm 0.05}$ $\underline{0.32}{\scriptstyle\pm 0.03}$ $0.0047{\scriptstyle\pm 0.0006}$ $0.22{\scriptstyle\pm 0.03}$ $1.74{\scriptstyle\pm 0.10}$ $24{\scriptstyle\pm 0.16} 1.1$$11{\scriptstyle\pm 0.17}$
+[…]
+[…]
 Add-and-Thin $0.86{\scriptstyle\pm 0.05}$ $\mathbf{0.30}{\scriptstyle\pm 0.04}$ $0.0043{\scriptstyle\pm 0.0007}$ $\underline{0.21}{\scriptstyle\pm 0.03}$ $\underline{1.53}{\scriptstyle\pm 0.14}$ $1.16{\scriptstyle\pm 0.16}$ $1.20{\scriptstyle\pm 0.14}$
 EventFlow (25 NFEs) $\underline{0.68}{\scriptstyle\pm 0.02}$ $1.06{\scriptstyle\pm 0.09}$ $0.028{\scriptstyle\pm 0.0015}$ $\mathbf{0.19}{\scriptstyle\pm 0.01}$ $2.30{\scriptstyle\pm 0.22}$ $\underline{0.90}{\scriptstyle\pm 0.05}$ $1.30{\scriptstyle\pm 0.03}$
-EventFlow (10 NFEs) ${0.60}{\scriptstyle\pm 0.02}$ $1.01{\scriptstyle\pm 0.31}$ $0.0113{\scriptstyle\pm 0.0016}$ ${0.16}{\scriptstyle\pm 0.01}$ $1.93{\scriptstyle\pm 0.07}$ ${0.77}{\scriptstyle\pm 0.05}$ ${1.14}{\scriptstyle\pm 0.07}$
+[…]
 EventFlow (1 NFE) $0.65{\scriptstyle\pm 0.16}$ $1.30{\scriptstyle\pm 0.61}$ $0.0445{\scriptstyle\pm 0.0104}$ $0.18{\scriptstyle\pm 0.04}$ $1.52{\scriptstyle\pm 0.17}$ $0.95{\scriptstyle\pm 0.17}$ $1.03{\scriptstyle\pm 0.24}$
@@
 EventFlow (25 NFEs, true $n$) $1.2{\scriptstyle\pm 0.01}$ $5.5{\scriptstyle\pm 0.3}$ $8.8{\scriptstyle\pm 0.2}$ $1.8{\scriptstyle\pm 0.02}$ $0.7{\scriptstyle\pm 0.01}$ $0.7{\scriptstyle\pm 0.02}$ $1.1{\scriptstyle\pm 0.02}$
+E.5 Runtimes
+In Table 21, we report the wall-clock time required to generate 1 0001\,000 sequences for each method.
+As expected, NHP and IFTPP achieve the fastest generation times, though this comes at the cost of substantially lower forecast quality (see Figure 2).
+Both Diffusion and Add-and-Thin are relatively slower due to their iterative refinement procedures.
+EventFlow (with 1 NFE) attains state-of-the-art forecasting accuracy while maintaining generation times comparable to IFTPP on datasets with moderate sequence lengths (Yelp-A, Yelp-M, Taxi).
+On datasets with longer sequences (e.g., Reddit-S), the transformer-based architecture of EventFlow naturally incurs higher computational cost.
+Although EventFlow with 25 NFEs is slower, Table 1 shows that a single NFE already suffices to reach SOTA performance.
+We note that these runtime comparisons are meant as approximate indicators, since implementation and architectural details can substantially influence wall-clock times.
+Overall, EventFlow achieves competitive generation efficiency despite not being explicitly optimized for speed.
+Table 21: Wall-clock runtimes (seconds) required to generate 1 0001\,000 sequences from each model, at the largest batch size which fit in memory.
+PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
+IFTPP $16.15$ $20.77$ $69.14$ $22.62$ $1.85$ $6.55$ $9.60$
+NHP $0.99$ $6.13$ $11.94$ $1.19$ $0.83$ $0.74$ $0.89$
+Diffusion $43.77$ $99.82$ $262.26$ $78.77$ $43.41$ $29.82$ $32.85$
+Add-and-Thin $149.09$ $261.53$ $292.51$ $42.60$ $16.71$ $34.73$ $34.87$
+EventFlow (1 NFE) $35$ $739$ $885$ 88 $42$ 66 $10$
+EventFlow (25 NFEs) $769$ $17889$ $19624$ $187$ $960$ $142$ $233$